A Complete Bibliography of Publications in Biometrics: 1960–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

20 September 2024 Version 2.01

Title word cross-reference

2^{n} [189, 756]. 2 × 2 [798, 381, 735, 263, 460, 257]. 2 × 3 [890]. 2 × m [194]. 2 × N [455]. 3 × 4 [2]. 41 ≤ A ≤ 50, B ≤ A [798]. 5 × 2^{2} [74, 10]. χ^{2} [816]. F [797, 683, 629]. k [593]. λ [954]. Q [573, 569]. r [875]. S [231]. t^{2} [815]. U [755]. y = α + δx + βρx [156].

-Direct [316, 427, 553, 648]. -log [881]. -Ratio [797, 629]. -Stage [231]. -Way [875].

1940 [958]. 1949 [575]. 1958 [354]. 1959-61 [989]. 1965 [800, 764]. 1965-30 [706]. 1966 [706]. 1967 [760]. 1st [800].

2 [805, 307]. 27th [800]. 2nd [800].

3 [306]. 34th [488]. 3rd [136].

57 [448].

650 [28].

704 [198].

Abbreviated [261]. Abilities [691]. absorbent [938]. Abstract [829, 932]. Abstracts [15, 34, 53, 78, 97, 117, 137, 202, 221, 241, 269, 287, 308, 326, 361, 386, 413, 449, 491, 547, 579, 612, 634, 662, 684, 718, 744, 768, 807, 836, 867, 931, 968, 998, 1026, 1053]. Academie [448]. Accept [336]. Acceptance [669]. Accident [727]. Accommodate [208]. Accountancy [897]. Achieving [340]. Action [765, 956, 401, 556]. Active [891]. Activities [380, 270]. Activity [891]. Adaptation [781]. Addition [424]. Additional [208, 652]. Additive [691, 45]. Additivity [138, 101]. Adequacy [504, 1018]. Adjoint [850, 927]. Adjusted [923]. Adjustment [842, 274]. Adolphe [893]. Adult [840]. Advanced [990]. Advisory [800]. Aerosols [529]. Affected [617]. after [545]. Against [587]. Age [3, 122]. Age-Distribution [122]. Ageing [993]. Agents [465]. Agricultural [486]. Agriculture [523, 305]. Agronomie [95]. Akin [6]. Alcool [576]. Alcoolisation [576]. Alcoolisme [576]. Algebra [861, 657, 659]. Algorithm [324, 757, 944, 756, 952]. All-or-None [845, 978, 260]. Alleles [592, 24, 7, 593]. Allelic [929]. Allocation [48, 321, 317, 128]. Allometry [672, 303]. Almost [774]. Alternative [72, 163, 468, 239, 21]. Alternatives [883]. among [292, 398, 639, 914, 586, 941, 824]. Anaesthetic [27]. Analogues [180]. analyse [624]. Analyses [492, 339, 168, 322, 1012, 282]. Analysis [849, 138, 685, 930, 554, 1035, 994, 391, 759, 844, 438, 945, 63, 470, 506, 40, 349, 891, 676, 218, 501, 560, 62, 787, 300, 681, 20, 502, 101, 421, 709, 995, 322, 643, 571, 393, 638, 877, 231, 259, 61, 497, 1032, 462, 773, 1004, 979, 25, 91, 828, 912, 911, 565, 749, 774, 910, 881, 513, 700, 503, 566, 640, 641, 673, 2, 689, 846, 782, 89, 500, 595, 348, 372, 195, 1034, 472, 65, 564, 209, 169, 346, 991, 432, 731, 852, 90, 481, 46, 642, 471, 507, 814, 813, 276, 624, 294, 1056]. Analysis [1057, 865, 740, 603]. Analytical [677]. Analyzing [754]. Angles [438, 851]. Angular [1033]. Animal [682, 59, 151, 608, 44]. Announcements [17, 36, 56, 80, 99, 120, 162, 184, 204, 223, 244, 272, 290, 311, 330, 363, 388, 418, 451, 494, 525, 549, 582, 614, 636, 664, 687, 720, 746, 770, 809, 838, 869, 904, 934, 970, 1000, 1028, 1055, 141]. Annual [714, 715, 535, 764, 706, 306]. Anova [545, 796, 1018, 695, 920]. Antagonistic [319]. Anti [465]. Anti-Cancer [465]. Antibiotic [402]. Antibiotics [673]. Antigenicity [426, 669]. Anwendung [899]. Aplicatii [1022]. Application [816, 938, 70, 154, 256, 589, 675, 31, 440, 1031, 577]. Applications [164, 165, 328, 1046, 67, 86, 676, 337, 651, 233, 212, 382, 553, 659, 250, 710, 710]. Applied [689, 846, 740, 996, 410]. Appreciation [601]. Appreciations [225]. Approach [52, 709, 675, 340, 876, 341, 431]. Approaches [690]. Approximate [345, 542, 503]. Approximation [589, 504, 1018].

Approximations [597, 734, 211]. Aquatic [909]. Arbitrarily [693]. Architect [372]. Areas [627]. Arising [751, 428]. Arithmetic [476]. Armitage [200]. Army [489, 96]. Arrangement [750]. Ascent [85]. Aspects [152, 813]. Assay [468, 728, 622, 588, 27, 42]. Assays [670, 1017, 782, 316, 427, 553, 648, 1059, 87]. Assessment [341]. Assisted [487]. Associate [2]. Associated [397, 286, 1041, 338]. Association [573]. Assumption [755, 427]. Assumption-free [755]. Asymmetrical [787, 564]. Asymptotic [843, 1018, 178, 14, 111, 856, 90, 1042]. asymptotically [318]. Attachment [463, 190]. Attribute [913]. Attributes [286]. Augmented [129]. Augmenting [208]. Australia [270, 380]. Authorship [606]. Automatic [306]. Available [521, 456]. Average [348]. Avram [574]. Axis [885]. Aylmer [365, 270, 367, 372].

B [705, 134, 966, 706, 800, 798, 609, 712, 484, 33, 957, 861, 381, 897, 610]. Back [389]. Background [168]. Backgrounds [972]. Bacteria [463, 1008, 673, 532, 190]. Bacterial [589, 789, 701]. Bacteriophage [559]. Bacterium [296]. Bailey [352]. Baker [444]. Balance [423, 759, 126, 610]. Balanced [849, 228, 783, 1011, 89, 758, 558, 626, 128, 786, 276]. Bancroft [265]. Bands [245, 187]. Bartlett [199, 198]. Based [581, 131, 444, 465, 300, 680, 499, 822, 786, 318]. Bases [862]. Basic [411, 897, 959, 442]. Basis [913, 32, 133]. Baule [951]. Bayesian [604, 707, 1031]. Bayo [989]. be [421]. Beament [180]. Bechhofer [1020]. Becker [479, 857]. Beetle [283]. Beetles [461, 840]. Behavior [108]. Behavioral [1051]. Behaviour [697]. Behrens [345, 242, 377]. Benjamin [33]. Bennett [798, 381]. Berckel [609]. Berlin [965]. Berlyand [304]. Best [171, 318]. Between [13, 501, 560, 913, 397, 617, 168, 1019, 674, 585, 113, 322, 434, 543, 319, 248]. between-Plot [585]. Beyer [762]. Bias [842, 919]. BIB [1017]. Bibliography [317, 354, 477, 575, 713, 958]. Bicking [477]. Binomial [797, 256, 589, 629, 949, 93, 848, 170, 60]. Bio [1059, 670, 1017, 487]. Bio-Assays [670, 1017, 1059]. Bio-Medical [487]. Bioassay [414, 371, 1037, 702, 552, 761]. Biological [468, 27, 972, 956, 42, 529, 457, 340, 181, 897, 530, 659, 658]. Biologie [1022, 95, 710]. Biologist [373]. Biologists [657]. Biologov [360]. Biology [180, 445, 180, 961, 220]. Biomathematics [861]. Biometeorology [805]. Biometria [267]. Biometric [164, 387, 67, 375, 107, 1025, 16, 35, 79, 98, 119, 140, 161, 183, 203, 222, 243, 271, 289, 310, 362, 417, 450, 548, 580, 613, 635, 663, 686, 719, 745, 769, 808, 837, 933, 969, 999, 1054, 1027, 329, 903, 55, 524, 493, 868]. Biometrics [470, 506, 748, 469, 963, 972]. Biometrie [962]. Biometrische [898]. Biometry [201, 811, 369, 892, 167, 359]. Biostatistics [574]. Birth [351, 328, 1032, 282, 233, 781]. Birth-Rank [282]. Bivariate [799, 885]. Bladder [587]. Bliss [894]. Block [849, 1059, 63, 228, 677, 670, 20, 783, 1011, 89, 275, 758, 558, 626, 652].

Blocking [678]. Blocks [651, 422, 626]. Blood [191, 777, 378]. Body [640]. Boehm [1048]. Book [518, 990, 858, 865, 574, 482, 486, 602, 660, 705, 894, 994, 268, 199, 52, 359, 448, 1020, 115, 444, 134, 656, 180, 604, 707, 895, 860, 411, 517, 607, 966, 116, 354, 357, 412, 446, 358, 477, 575, 519, 522, 478, 572, 479, 523, 488, 489, 520, 515, 573, 706, 713, 805, 802, 703, 711, 762, 800, 764, 803, 709, 716, 765, 766, 798, 804, 799, 829, 834, 958, 956, 961, 1049, 992, 993, 995, 989, 965, 158, 960, 352, 742, 360, 898, 900, 899, 896, 307, 741, 265, 658, 609, 443, 483, 159, 135, 306, 384, 655, 356, 712, 96, 355, 160, 32, 606, 94, 480, 484]. Book [901, 240, 33, 445, 1021, 830, 833, 863, 957, 1048, 1051, 1050, 861, 136, 181, 381, 385, 485, 75, 266, 1022, 897, 305, 577, 657, 708, 831, 832, 1023, 383, 267, 95, 743, 859, 76, 77, 576, 862, 200, 1024, 382, 487, 608, 740, 763, 864, 959, 964, 133, 304, 710, 220, 605, 996, 353, 659, 603, 410, 962, 991, 654, 610, 857, 1025, 516, 704, 442, 963, 801, 481, 546, 997]. Books [447, 490, 578, 611, 661, 717, 767, 806, 835, 866, 902, 967, 1052]. Both [730]. Bottenberg [410]. Boundary [428]. Bradley [31]. Branching [409, 214, 781]. Breast [945]. Breeding [691, 1010]. Brown [478]. Brownlee [602]. Bücher [965]. Buckley [711]. Budapest [965]. Bulgarie [448]. Bulgren [382]. Bulk [477]. Bulman [763]. Bulmer [1021]. Business [765].

C [441, 858, 482, 894, 994, 607, 446, 477, 478, 572, 703, 711, 764, 803, 798, 804, 658, 443, 483, 135, 160, 901, 833, 861, 385, 383, 76, 487, 710, 610]. Cacenes [487]. calcul [710]. Calculated [756]. Calculation [28, 890]. Calculus [861]. Calot [480]. Campbell [703]. Can [421]. Canada [711]. Cancer [945, 465, 556]. Cannibalism [840]. Canonical [912]. Capecchi [833]. Caractères [841, 974]. Carcinogenesis [595]. Carcinogenic [956]. Carcinogens [556]. Caries [675]. Carlo [173, 171, 1018]. Carriers [512]. Carry [620]. Carry-Over [620]. Case [885, 845, 978]. Cases [405, 834]. Catch [40]. Catchability [623]. Categorical [1056, 1004, 403]. Causal [697]. Causes [989]. Cell [777]. Cells [191]. Cellular [779]. Censored [497, 298, 318]. Census [127]. Censuses [136, 26]. Cent [798]. Certain [416, 645, 218, 504, 939, 456, 212, 595, 315, 472]. Chain [825]. Chains [675, 938, 822]. Challenging [519]. Chamber [529]. Change [1060, 423, 343, 511, 983, 731, 803]. Change-Over [1060, 423, 511, 983, 731]. Changeover [759]. Characteristic [577, 336]. Characteristics [621]. Characterizing [855]. Characters [61, 248]. Charles [446]. Charts [133]. Chassan [264]. Check [756]. Checks [757]. Chemical [956, 559]. Chemotherapy [671]. Chi [288, 843, 105, 45, 457, 630, 1042]. Chi-Square [288, 105, 45, 457]. Chiang [572]. Chiasma [368]. Chick [588]. Children [157]. Choice [830]. Chronic [69]. Cigarette [556]. Clair [608]. Clark [266]. Class [751, 595]. Classes [147, 432]. Classical [802]. Classification [794, 468, 644, 875, 83, 681, 322, 231, 950, 253, 286, 498, 219, 150]. Classifications [643, 987, 232, 952]. Classified [941]. Classifying [656]. Clines [640]. Clinical [1060, 946, 511, 533, 822, 192]. cliniques [385]. closed [192]. Cluster [502, 773]. Clustering [876]. Clusters [405]. Cochran [515, 569]. Coefficient [942, 573, 775, 910, 542, 509, 537, 917]. Coefficients [1058, 339, 699, 11, 871, 598, 65, 599, 21, 188, 492]. Coincident [777]. Coleman [703]. Coll. [640]. Collection [1048]. Column [1045, 584]. Combination [671, 309, 237, 790, 460, 194]. Combinations [174]. Combined [277, 259, 212, 89]. Combining [585, 11, 691, 735, 258, 533]. Come [405]. Comment [702, 761]. Commerce [897]. Committee [956]. Common [849, 171, 259, 89]. Commonly [253]. Communautés [501, 560]. Communities [501, 560]. Comparative [341, 669]. Comparing [280, 922, 435]. Comparison [818, 333, 886, 350, 1036, 8, 773, 692, 640, 536, 690, 819, 1006, 319]. Comparisons [1019, 113, 395, 20, 274, 434, 104, 114, 145, 567, 652, 752, 654]. Compartment [431]. Competing [165, 173, 86, 977]. Competition [3, 928, 566, 723]. Competitive [401]. Complementary [21]. Complete [416, 423, 759, 315]. Complex [1038, 709, 590]. Component [232, 624, 814]. Component-estimation [814]. Components [794, 930, 1043, 696, 644, 875, 397, 505, 812, 48, 944, 937, 41, 586, 510, 176, 219, 150, 907, 1040, 852, 1009, 248]. composantes [624]. Composed [248]. Composite [499, 581]. Compounds [23]. Computation [923]. Computer [981, 1047, 198, 487, 642, 382]. Computer-Assisted [487]. Computers [616, 355]. Computing [72, 163, 171, 851, 699, 570, 198, 150, 794]. Concentrates [38]. Concentration [532]. Concept [561]. Concepts [399, 442, 21]. Concerned [937]. Concerning [265, 982, 853, 153]. Concomitant [555, 667]. conditional [22]. Conditions [122]. Confectionery [341]. Conference [489, 800]. Confidence [245, 475, 302, 986, 647, 187, 299, 68, 92, 774, 955, 542, 854]. Confounded [787, 564]. Confounding [923, 176]. Consideration [165, 86, 148]. Considerations [672, 735]. Consistency [505]. Consisting [3]. Constant [880]. Constantino [653]. Constructing [73]. Construction [1044, 496, 787, 887, 690, 564]. Consulting [908, 988]. Contagious [521, 802, 727, 456]. Containing [514, 254]. Content [436]. Contests [764]. Context [821, 888]. Contingency [844, 878, 980, 947, 175, 798, 45, 890, 590, 381, 460, 210, 432, 194]. Continuous [819]. Contributed [248]. Contribution [937]. Contributions [414, 371, 375, 682, 378, 607]. Control [113, 395, 38, 151, 133, 652]. Controls [730, 978]. Correct [681]. Corrected [112, 598]. Correction [74, 54, 138, 139, 118, 245, 327, 288, 328, 581, 492, 521, 633, 685, 794, 795, 796, 1059, 1061, 1062]. Corrections [163, 164, 165, 201, 182, 309, 414, 415, 416, 797, 927, 928, 929, 930, 1042, 1043, 1044, 1045, 1046, 1056, 1057, 1058, 1060]. Correlated [753, 229]. Correlation [416, 1058, 942, 11, 910, 598, 854, 315, 599, 543, 430, 210, 943, 919]. Correlations [302, 922, 674, 130, 24, 123]. Corstius [609]. Count [9]. Counter [777]. Counts [532]. Courbes [668]. Cours [480]. Covariance [201, 1035, 397, 886, 168, 218, 814, 912, 2, 348, 1034, 167, 65, 248, 1005].

Covariances [796, 920, 695]. Covariate [685, 565]. Cows [731]. Cox [75]. Creel [26, 127]. Crew [704]. Criteria [327, 1037, 293]. Criterion [286]. Critical [597, 1008, 71]. Cronbach [858]. Crop [347, 896, 627, 396, 628]. Crops [750]. Cross [182, 1061, 872, 349, 279, 398, 638, 916, 103, 195]. Cross-Fertilization [872]. Cross-Overs [349]. Crosses [278, 472]. Crossing [733, 472]. Crump [314]. Culling [753, 146, 563]. Cultures [779]. Cumulative [847]. Curvature [168]. Curve [139, 40, 1014, 879, 88, 156, 514, 823]. Curves [245, 1057, 280, 619, 187, 407, 531, 979, 828, 435, 725, 668, 590, 336, 819]. Cyclic [783, 1011, 758].

D [518, 115, 604, 446, 703, 899, 265, 606, 181, 381, 385, 75, 305]. D. [640]. Dam [788]. Dams [1039]. Dans [698, 974, 384]. Darling [706]. Dasgupta [897]. Data [521, 1056, 554, 913, 568, 83, 825, 13, 619, 589, 709, 995, 625, 691, 462, 1004, 28, 540, 734, 727, 977, 282, 456, 754, 921, 689, 846, 315, 1007, 403, 1040, 93, 82, 781, 991]. Date [416]. Daughter [788]. Daughters [1010]. David [1051, 654]. Davis [573]. Death [351, 328, 989, 1032, 233, 781]. Deaths [834]. Decay [529, 819]. Decision [171, 996]. Decisions [858, 307]. Decomposition [944, 240]. Decrement [749]. Defares [220]. Defense [587]. Deficiency [630]. Degree [853]. Degrees [65, 188]. Demographie [710]. Denominator [543]. Densities [820]. Dental [675]. Depend [1032]. Dependent [350, 497]. deren [899]. Derivation [397]. Deriving [261]. Descriptive [480]. Design [74, 1060, 994, 728, 154, 683, 489, 347, 459, 511, 914, 295, 39, 854, 845, 2, 348, 10, 374, 815, 188, 252]. Designs [581, 1045, 1059, 620, 423, 759, 63, 228, 670, 787, 946, 392, 20, 129, 317, 343, 424, 783, 1011, 259, 651, 261, 110, 189, 1017, 207, 89, 275, 584, 758, 558, 784, 626, 785, 564, 737, 499, 128, 822, 786, 192]. Detection [64, 876]. Determination [647]. Determine [891]. Determinism [841]. Déterminisme [841]. Deterministic [319]. Deutsch [895]. Deutscher [965]. Development [489]. Developments [724]. Deviates [799]. Deviations [370]. Diagnosing [9]. Diagnosis [818, 425]. Diagnostic [109]. Diallel [182, 1061, 279, 278, 638, 914, 916, 103, 915, 733, 941, 472]. Diary [731]. Dictionary [517, 992]. Difference [292, 939]. Differences [697]. Different [333, 228, 639]. Differential [791, 818]. Diffusion [504]. Dilution [64, 729, 701, 532, 316, 427, 553, 648, 87]. Dimensional [351, 947, 233, 275, 328]. Dimensionen [1024]. Dimensions [640]. Diog [575]. Diploid [509, 1009]. Direct [73, 316, 427, 553, 648]. Direction [85]. Directional [791, 783, 1011]. Directions [1041]. Disc [77]. Discontinuity [1012]. Discrete [645, 802, 734, 147, 973]. Discriminant [591, 699, 732, 681, 1036, 570, 911, 774, 910, 226, 1005]. Discrimination [889, 618]. discriminatoire [624]. Discriminators [912]. discriminatory [624]. Discussion [587]. Disease [535, 151, 876, 209]. Diseased [415, 297, 320]. Diseases [834, 396]. Dispersions [570]. Disproportionate [322, 462, 82]. Disputed [606]. Distance [913, 886, 724, 639, 107].

Distances [591]. Distinct [458]. Distinguish [4]. Distributed [476]. Distribution [696, 324, 942, 6, 429, 61, 775, 68, 944, 885, 144, 727, 984, 726, 240, 694, 541, 949, 212, 172, 297, 855, 316, 648, 567, 528, 122, 298, 824, 22, 47, 170, 211, 778, 60, 134]. Distribution-Free [316, 648]. Distributions [164, 521, 333, 67, 473, 645, 504, 586, 147, 955, 542, 456, 841, 466, 819, 530, 802]. dlya [360]. Do [405]. Doig [354, 958]. Dominance [853]. Donald [134]. Dosage [531, 590, 782]. Dose [1057, 884, 979]. Double [327, 398, 343, 293, 255, 195, 626, 426, 186]. doubly [318]. Douglas [358]. Dr. [242]. Draper [740]. Dressing [132]. Drosophila [640]. Drug [1003, 295, 743]. Drugs [416, 468, 401, 315, 690]. Due [777, 257]. Duncan [597, 71, 112]. During [1008]. Dutch [609]. dyad [888]. Dyadenkontextredundanz [888]. Dynamic [209].

Early [596]. East [628]. Ecological [501, 560, 6]. Écologiques [501, 560]. Ecology [199, 52]. Econometric [901, 516]. Econometrics [607, 357]. Economic [765]. Economics [897]. Ecosystem [909]. Edgerton [992]. Edition [136]. Editorial [217, 437, 230, 739]. Educational [703]. Edward [485]. Effect [277, 678, 1005]. Effective [730, 912]. Effectiveness [193, 587, 842]. Effects [1043, 620, 423, 759, 812, 679, 639, 340, 229, 498, 673, 756, 509, 126, 76]. Efficiencies [131]. Efficiency [534, 408, 1017, 44, 543, 257, 318]. Ehrenfeld [705]. Eight [914, 461]. einfaches [888]. Ekambaram [133]. Electronic [777, 355, 487]. Elementary [128, 518, 966]. Elemente [900]. Elements [33]. Elimination [129, 618, 276]. Embodiments [605]. Empirical [534, 942, 886, 556, 123]. Employing [20]. Endocarditis [589]. Energy [961]. Engineering [602]. Engineers [660, 960]. English [517]. Entre [501, 560]. Enumeration [940]. Enumerations [751, 571]. Environment [399]. Environmental [277, 459, 639, 733, 130]. Epidemic [924, 877]. Epidemics [512]. Epidemiologic [433]. Epidemiological [714]. Epidemiology [199]. Epistatic [1009]. Equal [623]. Equality [703]. Equation [436, 303, 239, 951]. Equations [49, 939, 864]. Equilibria [1038, 593]. Equilibrium [929, 826, 825, 748, 7]. Equipment [96]. Equivalence [699, 700]. Eradication [766]. Error [849, 554, 1035, 847, 48, 112, 880, 215, 304, 114]. Errors [617, 218, 701, 458, 229, 782, 338, 652]. Erzeugung [888]. Espèces [501, 560]. Especially [247]. Essais [385]. Essay [707]. Essays [607, 863, 1048]. Essence [359]. Essentials [470, 506, 471, 507]. Estimate [1016, 816, 215, 621]. Estimated [1058, 302, 599]. Estimates [492, 644, 339, 568, 131, 562, 505, 1014, 874, 335, 436, 249, 737, 123, 318]. Estimating [794, 1043, 736, 22, 825, 812, 392, 407, 464, 820, 459, 124, 701, 179, 191, 694, 229, 476, 150, 907, 781]. Estimation [118, 795, 928, 193, 49, 397, 85, 707, 527, 557, 780, 146, 43, 187, 555, 879, 1008, 132, 232, 649, 937, 41, 66, 1039, 889, 885, 984, 59, 977, 39, 668, 792, 723, 856, 733, 627, 281, 212, 172, 788, 596, 532, 219, 26, 1006, 316, 396, 628, 648, 545, 210, 234, 1040, 254, 30, 919, 93, 481, 298, 848, 917, 667, 814, 170, 186, 245, 668, 895, 356]. Estimator [985, 87]. Estimators [875, 108, 976, 408, 880, 255, 755, 258, 106, 284]. Étude [501, 560]. Evaluation [650, 733, 425, 213, 743]. Events [256, 589, 943]. Evidence [505]. Evolution [1025]. Exact [302, 890, 567]. Examination [459]. Example [560]. Examples [54, 457]. Exemple [560]. Exercises [831]. Expectation [406, 370]. Expectations [695, 261, 949, 920, 1009, 817, 796]. Expected [791, 508, 987, 335, 457, 910, 510, 509]. Experiences [373]. Experiment [795, 438, 527, 891, 840, 756, 472, 563, 652]. Experimental [482, 180, 995, 39, 854, 445, 689, 846, 737, 128, 815]. Experimentation [399]. Experimenting [422]. Experiments [671, 849, 1046, 1058, 1061, 394, 508, 85, 676, 227, 23, 168, 677, 218, 62, 350, 48, 321, 489, 729, 8, 730, 25, 91, 916, 1039, 880, 59, 92, 915, 566, 207, 595, 599, 346, 250, 440, 374, 642, 260, 994, 160, 75, 655]. Exploration [446]. Exponential [557, 780, 847, 555, 1014, 255, 985, 855, 819, 262, 298, 667, 781]. Exponentials [174]. Expository [49, 469]. exposure [621]. Expression [1008]. Expressions [517]. Extension [196, 178, 111, 47]. Extinction [789]. Extreme [789, 694]. Extreme-Value [694]. Extremes [94].

F [860, 519, 523, 703, 989, 307, 32, 606, 1050, 136, 743, 704]. F.R.S. [270]. Face [513]. Factor [470, 506, 471, 507, 683, 594, 481]. Factorial [208, 74, 787, 392, 756, 564, 10, 642, 188]. Factorials [581, 63, 499]. Factors [208, 406, 1032]. failure [621]. Faktorenanalyse [1024]. Falconer [115]. Family [28, 29, 338]. FAO [896]. Fascicule [710]. Federalist [606]. Feed [543]. Feedback [50]. Feeding [69]. Female [993]. Fertilization [872]. Fertilizer [132]. Fertilizers [896]. Fiducial [377]. Field [52, 399, 678, 266]. Fifty [519]. Finite [791, 825, 498, 509, 938]. Finite-State [825]. Finney [381, 305]. First [409, 423, 759]. Fisher [270, 414, 365, 225, 371, 473, 345, 380, 6, 375, 571, 1005, 871, 367, 379, 376, 378, 372, 373, 374]. Fishermen [26, 127]. Fit [949]. Fitness [639]. Fitted [691, 1007]. Fitting [139, 416, 646, 981, 174, 619, 1014, 285, 540, 1047, 456, 88, 315, 156, 514, 60, 149, 521]. Five [798]. Fixed [1043, 924, 812, 683]. Flamant [385]. Flieger [991]. Flour [283, 461, 840]. Flow [961]. Follow [165, 86, 157]. Follow-Up [165, 86, 157]. Fondamentaux [501]. Food [522]. Forest [415, 13, 800, 297, 320, 975, 963]. Forestry [400, 412]. Foreword [366]. Form [239]. Formal [609]. Formation [368]. Formula [1044, 887, 132]. Formulae [171]. Forstliche [962]. Fortran [862]. Foundation [481]. Foundations [704, 1025]. Four [786, 509]. Four-way [786]. Fractional [392]. Fractions [581, 499]. Fraga [653]. Franceschetti [893]. Frank [600, 721]. Free [251, 316, 648, 755]. Freedom [65, 188]. French [760, 501, 560, 668, 841, 624]. Fréquences [841]. Frequencies [736, 639, 457, 149, 917, 82]. Frequency [70, 466, 841]. Freudenthal [741]. Freund [960]. Front [18, 37, 57, 81, 100, 121, 142, 166, 185, 205, 206, 246, 273, 291, 312, 331, 364, 390, 419, 452, 453, 495, 526, 550, 583, 615, 637, 665, 688, 722, 747, 771, 810, 839, 870, 905, 935, 971, 1001, 1029].

Fruit [881]. Full [980]. Function [981, 699, 1036, 570, 1047, 856, 708, 1005]. Functional [108]. Functions [164, 591, 67, 473, 953, 539, 226, 594, 304, 817, 240, 577, 1049]. Fundamental [501, 427]. Fundamentals [957]. Further [505, 571, 424, 178, 148, 296, 553].

G [990, 486, 444, 860, 966, 354, 575, 515, 706, 802, 958, 1049, 356, 96, 480, 1021, 1048, 181, 266, 577, 77, 382, 763, 220, 353, 801, 481]. Gain [678, 543]. Galton [821]. Gamma [542, 689, 846]. Gaussian [696, 944]. Gavrilova [304]. Gelfand [1049]. Gene [639, 841, 509]. General [1044, 751, 860, 732, 679, 256, 589, 887, 642]. Generalised [1035, 694, 541]. Generalization [139, 539, 303, 88, 876, 537, 1013]. Generalized [591, 586, 239, 212, 90, 294, 125]. Generating [325, 106, 888]. Generation [758]. Generations [409]. Genes [247, 841]. Genetic [1058, 302, 508, 942, 679, 368, 352, 459, 937, 41, 841, 640, 733, 788, 599, 130, 24, 123, 919, 1025, 1009, 31, 748, 32]. Genetical [1038, 821]. Genetics [328, 751, 428, 504, 939, 571, 376, 233, 1025, 814, 624, 115, 479, 857, 704]. Génétique [841, 624]. Genome [1008]. Genotype [399, 177]. Genotype-Environment [399]. Genotypes [824]. Génotypiques [974]. Geometric [149]. Geometry [861]. George [32]. Gerald [134]. German [888]. Gilman [573]. Given [7, 888]. Glagoleva [1049]. Gleichzeitige [899]. Gleser [858]. Goitre [818]. Goldberg [116]. Goldberger [516]. Goldstein [574]. Goncalves [653]. Good [707]. Goodman [306]. Goodness [949]. Goodness-of-Fit [949]. Goon [957, 897]. Gordon [793]. Gradient [85]. Graduation [70]. Graphical [155, 1038, 648]. Graphically [729]. Graphs [1049]. Grasshoppers [4]. Grassland [829]. Graveyard [323]. Green [708]. Gregor [655]. Grizzle [816]. Group [849, 3, 800, 594, 558]. Grouped [1046, 356, 977, 454, 250, 440]. Grouping [626]. Groups [924, 307, 434, 378, 430]. Grouse [820]. Growth [1057, 4, 945, 591, 13, 1032, 979, 828, 435, 725, 454, 102, 1, 936, 596, 157, 882, 823, 1013, 122, 522]. Growth-Invariant [591]. Gumbel [94]. Gupta [957, 897, 957].

H [706, 713, 805, 711, 762, 804, 961, 992, 995, 158, 900, 741, 609, 159, 32, 181, 266, 383, 77, 740, 996, 410, 654]. Hacking [742]. Haiger [898]. Half [302, 985]. Half-Sib [302]. Half-Table [985]. Hamilton [96]. Handbook [762, 864]. Hans [832]. Hanson [486]. Hardy [446]. Harmonic [434, 513]. Harvests [227]. Haskell [793]. having [849]. Hays [959]. Health [714, 715, 572, 520, 803, 716, 445]. Healthy [415, 297, 320]. Heinisch [601, 632]. Hellman [96]. Helvey [76]. Henderson [907]. Henry [1050]. Heritability [492, 302, 339, 568, 131, 986, 146, 874, 1039, 179, 39, 215, 737]. Hermite [304]. Herrendorfer [899]. Heterogeneity [129]. Heterogeneous [434, 229]. Heuristic [30]. Hierarchal [683, 231]. Hierarchical [232, 952]. High [764]. Highly [54, 5]. Higley [77]. Histocompatibility [972]. Historical [711]. History [964]. Hobson [703]. Hodges [442]. Hoffman [353]. Homogeneity [201, 393, 950, 455, 212, 167, 65, 198]. Horst [798].

Horticultural [227]. Host [64, 729, 125]. Howard [657]. Hsu [381]. Humaines [384]. Human [513, 334, 378]. Huntsberger [265]. Hybrid [195]. Hybridisation [655]. Hybrids [398, 650]. Hyperbola [118, 646, 66]. Hypergeometric [134, 294]. Hypersensitivity [84]. Hypotheses [878, 421, 124, 853, 153, 135]. Hypothesis [860].

Ian [742]. IBM [28, 198]. IBM-650 [28]. Identification [1020, 535]. Identified [508]. II [764, 171, 507, 846, 483]. II. [506, 733]. III [710, 86]. III. [165]. Immediate [126]. Improved [402]. Improving [1016]. Inbred [1009]. Inbreeding [993, 28, 247, 509, 537, 917]. Incidence [556, 628]. Including [659]. Incompatibility [592, 7]. Incomplete [947, 63, 677, 670, 20, 783, 1011, 253, 89, 275, 758, 558, 626, 785, 667, 1059]. Incompletely [391, 734]. Independent [333, 299]. Index [496, 465, 647, 954, 251, 572, 680, 335, 925, 974, 177, 213]. India [628]. Indices [779, 918]. Indirect [496, 923, 538]. Indirect-Adjusted [923]. Indistinguishable [697]. Individual [168, 937, 978, 65, 188, 276]. Individuals [286, 248, 656]. Industry [341]. Ineffective [260]. Inefficient [154]. Infected [415, 297, 320]. Infection [587, 1008]. Infectious [834]. Inference [391, 443, 483, 734, 1002, 742, 801, 606]. Inferring [500]. Influence [236]. Information [58, 143, 224, 313, 420, 551, 666, 772, 906, 1030, 728, 585, 555, 914, 889, 667]. Inheritance [853]. Initial [1014]. Injections [191]. Insect [52]. Insects [234, 122]. Instance [323]. Institute [488]. Instrumentation [478]. Integral [304]. Intensities [59]. Inter [928, 63, 723]. Inter-Block [63]. Inter-Plant [928, 723]. Interaction [816, 878, 462, 399, 536, 275, 319, 46]. Interactions [421, 876, 756]. Interactive [315]. International [488, 800]. Interpretation [23, 638, 841]. Interrelations [586]. Interval [299, 68, 885]. Intervals [302, 986, 92, 774, 854]. Intervertebral [77]. Intra [63]. Intra- [63]. Intrablock [73]. Intraclass [430]. Introduction [470, 506, 445, 305, 220, 471, 507, 116, 658, 482, 705, 115, 604, 357, 352, 859]. Introductory [574]. Introgression [641]. Invariant [591]. Inverse [594, 921]. Inversion [864]. Inversions [873]. Investigation [534]. Investigations [716]. Invited [470, 506, 972, 748]. Involved [701]. Involving [391, 218, 335, 458, 734, 430]. Iosifescu [1022]. Irradiated [668]. Irradiés [668]. Irregular [581, 499]. Irreversible [675]. Isogenic [650]. Isolation [639]. Issue [760]. Iterated [298]. Iterative [1043, 812, 1040]. Ivan [830]. Iyer [799].

J [858, 359, 1020, 134, 180, 707, 446, 573, 703, 803, 709, 765, 829, 961, 960, 352, 609, 483, 94, 445, 181, 381, 385, 305, 708, 220, 410, 442]. Jackknife [755]. Jacob [441]. Jerome [443]. Joan [864]. John [865]. Jonge [158]. Jöreskog [481]. Jr [442]. Jr. [653, 410, 801]. Judith [573]. July [706]. Junctions [337]. June [706, 760].

Kärber [781]. Keilson [708]. Kendall [990, 354, 575, 958]. Keyfitz [991]. Khabaza [603]. Kiefer [1020]. Kinds [877, 207]. kinetic [294]. Kinetics [190, 294]. Kish [1023]. Kitov [355]. Kloek [833]. knowledge [186]. Knox [876]. Knut [412]. Kodlin [1062]. Kolmogorov [926]. Kotz [517]. Krinitskii [355]. Krishna [799]. Kulldorf [356]. Kurt [1024]. Kurtz [992]. Kutz [411].

L [858, 180, 572, 703, 804, 993, 900, 914, 159, 712, 96, 606, 14, 793, 266, 959, 442]. Lag [46]. Laha [577]. L’Analyse [501, 560]. Landwirtschaftliche [965]. Landwirtschaftsverlag [965]. Lange [357]. Large [873]. Latin [126, 72, 163, 424, 259, 325, 276, 883]. Latscha [381]. Lattice [2]. Law [1013]. Layout [503]. Lazarfeld [1050]. Least [880, 254]. LeClerg [266]. Lederman [576]. Ledermann [738]. Lee [473, 314]. Leenders [833]. Lehman [442]. Lellouch [385]. Leonard [266]. Leser [901]. Leslie [1023]. L’Essai [698]. Leukemia [405]. Level [208, 392]. levels [345]. Li [443, 482, 483, 160]. Liaisons [501, 560]. Lieberman [134]. Life [164, 165, 67, 86, 989, 692, 596]. Likelihood [521, 732, 879, 649, 41, 880, 668, 255, 700, 456, 210, 917]. Liljedahl [993]. Limited [1008]. Limits [475, 647, 955, 542]. Lindgren [484]. Lindley [604]. Line [728, 43, 820, 976, 540, 51]. Linear [1042, 1056, 948, 860, 617, 843, 174, 884, 108, 1004, 911, 880, 754, 856, 864, 514, 882, 410, 90, 431, 1015, 618, 1005]. Linearity [514, 254]. Lines [153]. Linkage [352]. Linked [929, 826, 247]. Links [501, 560]. Listing [510]. Literature [354, 575, 958]. Littauer [705]. Little [445]. Local [27]. Locus [929, 826]. log [881]. Logarithmic [324, 6, 172]. Logic [742]. Logistic [139, 124, 88, 239, 1013]. Logit [816, 541]. Logit-Normal [541]. Logits [980]. Lognormally [476]. Lognormally-Distributed [476]. Long [107]. Longitudinal [619, 229]. Look [907]. Looking [964]. Loss [509, 396, 257, 260]. Love [96]. Lowell [660]. Lukacs [577]. Lung [556]. Lysine [588].

M [990, 199, 1020, 354, 358, 575, 703, 711, 803, 798, 829, 958, 1049, 609, 384, 445, 793, 1021, 957, 1051, 381, 1022, 897, 657, 763, 603, 963]. MAA [764]. Machine [28]. Macnaughton [656]. Macnaughton-Smith [656]. Main [756]. Mainland [518]. Maintained [7]. Majeurs [841]. Major [885, 841]. Maki [706]. Malaria [766]. Mammalian [1007]. Management [975]. Manner [1]. MANOVA [169]. Manquants [698]. Mantegazza [743]. Manual [479, 523, 857]. Marginal [82]. Marine [383]. Mark [563]. Markov [938, 822, 825, 675]. Mason [382]. Mass [538]. Matched [534, 845, 209, 433]. Matching [775, 293, 978, 327]. Matematicheskaya [360]. Material [317]. Materials [477, 76]. Matern [800]. Mathematical [928, 468, 750, 945, 27, 572, 352, 877, 529, 191, 340, 723, 1048, 1050, 1007, 268, 1051]. Mathematics [523, 682, 220, 964, 830, 833, 863]. Mathematische [712]. Mating [459, 914, 938]. Matrices [201, 927, 850, 886, 167, 1005].

Matrix [749, 864, 975, 659]. Matter [18, 37, 57, 81, 100, 121, 142, 166, 185, 205, 206, 246, 273, 291, 312, 331, 389, 364, 390, 419, 452, 453, 495, 526, 550, 583, 615, 637, 665, 688, 722, 747, 771, 810, 839, 870, 905, 935, 971, 1001, 1029]. Maximum [409, 521, 214, 732, 879, 649, 41, 880, 668, 255, 700, 456, 210, 917]. mays [914]. McCarthy [709]. McCulloch [605]. McIntosh [358]. McPartland [703]. Mean [1016, 508, 953, 987, 510, 920, 262, 695]. Meaning [702, 552, 761]. Means [796, 333, 218, 474, 393, 792, 673, 476, 186]. Measure [827]. Measurement [554, 1003]. Measurements [676, 62, 674]. Measures [986, 943]. Measuring [25, 236]. Medgyessy [240]. Medical [518, 1036, 96, 425, 200, 487, 353]. Medicina [1022]. Medicine [414, 371, 220, 382]. Medikov [360]. mehreren [899]. Memoriam [365]. Memories [379]. Men [76]. Mendel [655]. Mendelian [429]. Mendenhall [859]. Merit [918]. Merkmalen [899]. Messick [1051]. Method [72, 163, 521, 696, 1038, 174, 884, 20, 502, 464, 1014, 820, 976, 124, 701, 912, 179, 191, 73, 247, 774, 668, 755, 700, 445, 456, 503, 325, 1, 595, 563, 156, 514, 654, 917, 781, 276, 888]. Méthode [668]. Methoden [898, 159]. Méthodes [384, 95]. Methodology [602, 148]. Methods [894, 699, 716, 766, 829, 158, 773, 435, 692, 39, 754, 136, 733, 172, 708, 509, 907, 316, 553, 648, 30, 170, 705, 707, 966, 743]. Mice [993]. Michail [962]. Microbial [438]. Microbiological [38]. Microbiology [383]. Microorganismes [668]. Microorganisms [92, 668]. Middle [107]. Migration [639, 631, 69]. Miller [960, 801]. Mind [605]. Mineral [956]. Mingling [776]. Minimum [816, 889]. Misclassification [774, 910, 263, 257]. Missing [392, 916, 911, 472, 652, 1061]. Mite [628]. Mitscherlich [436]. Mixed [794, 1043, 872, 645, 812, 683, 1012, 498, 150, 813]. Mixtures [416, 468, 401, 1003, 315, 690]. Mode [556]. Model [72, 283, 163, 794, 928, 1043, 750, 467, 557, 780, 812, 1019, 679, 683, 256, 421, 61, 8, 124, 1012, 253, 559, 556, 985, 723, 840, 498, 9, 150, 1007, 530, 545, 631, 975, 319, 69, 190, 109, 188, 813, 125]. Models [416, 1056, 468, 391, 199, 470, 23, 972, 625, 531, 1004, 261, 734, 977, 590, 229, 673, 536, 296, 315, 690, 275, 1002, 432, 254, 918, 471, 444, 572, 180]. Modern [707, 657]. Modification [926]. Modified [1016, 404, 8, 1017]. Mokken [609]. Moment [875]. Moments [700]. Monofactorial [853]. Monte [173, 171, 1018]. Mood [703]. Morbidity [716]. Morowitz [961]. Morphological [32]. Morphologically [697]. Morris [829]. Mortality [795, 562, 527, 256, 59, 955, 461, 281, 249]. Mosaic [776]. Mosaics [561]. Most [912]. Mosteller [606, 519, 8]. Multi [351, 328, 929, 233, 594, 821]. Multi-Allelic [929]. Multi-Dimensional [351, 233, 328]. Multi-Factor [594]. Multi-Type [821]. Multiallelic [826]. Multicompartment [69]. Multifactor [643]. Multinomial [344]. Multiple [171, 585, 113, 395, 321, 597, 112, 253, 911, 565, 749, 910, 978, 24, 114, 145, 410, 71, 752, 685]. Multiple-Decision [171]. Multivariate [54, 4, 83, 676, 5, 300, 625, 135, 982, 672, 912, 303, 372, 936, 169].

Multivariate-Covariance [912]. Multiway [82]. Mutagenesis [559]. Mutants [402]. Mutation [7].

N [352, 355, 831, 267, 740, 220, 991, 431, 409]. nach [899]. Nagai [706]. Nahikian [657]. Nations [520]. Natural [894, 422, 234]. Nature [610]. Nearest [408]. Necessary [910]. Negative [847, 874, 730, 949]. Neighbour [408]. Neil [1050]. Nested [930, 322, 987, 852]. Network [431]. News [17, 36, 56, 80, 99, 120, 162, 184, 204, 223, 244, 272, 290, 311, 330, 363, 388, 418, 451, 494, 525, 549, 582, 614, 636, 664, 687, 720, 746, 770, 809, 838, 869, 904, 934, 970, 1000, 1028, 1055, 141]. Newspaper [609]. Neyman [211]. Nicolas [964]. Ninth [489]. Niven [830]. No [816, 878]. Nomography [485]. Non [818, 394, 344, 1011, 1018, 880, 498, 754, 856, 315, 553, 514, 1040, 254, 90, 1009]. Non-Asymptotic [856]. Non-Constant [880]. Non-Directional [1011]. Non-Inbred [1009]. Non-Interactive [315]. Non-Linear [754, 856, 514, 90]. Non-Linearity [514, 254]. Non-Normality [394]. Non-Orthogonal [1040]. Non-Orthogonality [498]. Non-Parametric [344, 1018, 553]. Non-Toxic [818]. None [845, 978, 260]. Noninteractive [416]. Nonlinear [939, 539]. Nonorthogonal [651]. Nonparametric [563]. Normal [791, 171, 333, 473, 799, 299, 68, 885, 734, 792, 854, 541, 318]. Normality [394]. Norman [307]. North [628]. North-East [628]. Norwegian [412]. Notation [816]. Note [72, 74, 1016, 264, 439, 49, 324, 986, 12, 645, 1019, 50, 217, 569, 1014, 1018, 987, 259, 132, 112, 982, 261, 828, 178, 110, 45, 73, 984, 1017, 111, 985, 102, 437, 263, 325, 983, 230, 739, 51, 262, 1013, 1015, 440, 260, 351, 409, 791, 163, 797, 929, 930, 406, 439, 475, 759, 736, 301, 438, 302, 826, 196, 349, 473, 891, 825, 154, 699, 218, 953, 175, 350, 954, 683, 407, 474, 197, 569, 571, 347, 597, 757, 651, 827, 649, 285, 570, 701, 828, 178, 28, 408, 889, 179, 925, 303, 650, 92, 955, 890]. Note [792, 14, 286, 598, 629, 755, 700, 926, 854, 1041, 853, 856, 735, 239, 476, 630, 758, 215, 348, 756, 855, 29, 219, 472, 198, 599, 10, 156, 514, 737, 346, 545, 177, 631, 852, 888, 30, 31, 440, 93, 652, 952, 824]. Notes [378]. Nuclear [76]. Number [1008, 7]. Numbers [568, 349, 228, 322, 643, 701, 458, 814, 160]. Numerical [656, 887, 152, 864, 1044, 603].

O [487, 304]. Obituary [441, 600, 653, 314, 242, 893, 793, 738, 632]. Objects [697]. Observation [598, 693, 430]. Observational [842, 995]. Observations [64, 908, 643, 988, 983, 167, 652, 260, 201]. Obtained [492, 339, 215]. Obtaining [1014, 874, 774]. Occurrences [621]. Occurring [253]. October [800]. Offspring [131, 152]. Oils [956]. One [3, 171, 508, 467, 798, 129, 598, 235, 65, 430]. One-Way [129, 65]. Ones [297]. Only [508]. Operating [336]. Operation [197]. Oppenheimer [383]. Opportunity [703]. Optimal [321, 426]. Optimum [85, 48, 132, 151, 29, 737, 177, 823]. Orchard [458]. Orchards [783, 1011, 458]. Order [264, 12, 196, 500, 126]. Ordered [403]. Organism [102]. Organisms [454, 528]. Organization [714, 715, 520, 716]. Oriented [729]. Origin [51]. Orthogonal [1044, 474, 887, 731, 1040]. Orthogonality [498]. Osborne [32]. Oskar [357]. Other [656, 539, 863, 126]. Otto [485]. Ottokar [601, 632]. Outline [412]. outlook [1031]. Overall [827]. Overs [349]. Owen [134].

P [656, 802, 709, 799, 384, 240, 381, 1022, 743, 200, 304]. P.B.I.B. [2]. Pain [148]. Pair [209]. Paired [1019, 20, 8, 274, 536, 654, 567]. Paired-Comparison [8]. Pairing [260]. Pairs [922, 20, 573, 845, 752]. Pairwise [948]. Paleontologic [201, 167]. Paleontological [936]. Paper [470, 506, 972, 748, 440]. Papers [760, 804, 265, 469]. Parabolas [464]. Paradox [982]. Parallel [728]. Parameter [596, 667, 22, 318, 781]. Parameters [1014, 879, 124, 335, 255, 856, 788, 819, 298, 848]. Parametric [344, 1018, 553]. Parent [131, 508, 152]. Parents [1010]. Part [288, 105]. Partial [182, 1061, 11, 278, 639, 916, 103, 733]. Partially [356, 558]. Partitioning [13, 45]. Pascal [144]. Passage [777]. Patchily [297]. Patchworks [342]. Paternal [302]. Path [871, 675, 90, 21, 46]. Path-Probability [675]. pathogen [125]. Patients [434]. Patil [802]. Pattern [439, 342]. Patterns [4, 461, 102, 561, 610]. Paul [1050]. Pawlik [1024]. Peak [535]. Pearce [658]. Pearson [804, 473, 804]. Peas [301]. Peng [994]. Pèpe [384]. Percentage [568]. Percentages [82, 573]. Percentiles [694]. Period [1060, 511, 229, 815]. Periods [535]. Perrier [384]. person [621]. Personal [379]. Personnel [858]. Pests [396]. Pflanzen [899]. Pflanzen- [899]. Phage [463, 296, 190]. Pharmacology [158]. Phase [342, 153]. Phases [776]. Phenomenon [789]. Phenotype [917]. Phenotypic [130]. Phosphate [436]. Physical [529]. Physiological [19, 446]. Piccinini [743]. Pillai [135]. Pitfalls [323]. Planning [622, 607, 75]. Plans [208]. Plant [928, 750, 501, 560, 277, 691, 655, 723, 235]. Plants [891]. Plating [701]. Plochinsky [267]. Plot [585, 91, 266, 348, 785]. Plusieurs [974]. Pockets [639]. Points [798]. Poisson [406, 817, 22, 47, 953, 318, 589, 370, 649, 144, 170, 778, 60]. Poissonian [848]. Polychotomous [42, 590]. Polynomial [927, 850]. Polynomials [1044, 474, 887, 594, 921, 731, 304]. Ponderation [974]. Population [409, 328, 924, 199, 214, 52, 562, 428, 522, 803, 939, 820, 1032, 335, 540, 749, 454, 498, 233, 281, 1006, 991, 973, 234, 122, 7]. Populations [3, 913, 171, 886, 638, 299, 827, 914, 915, 792, 640, 873, 476, 235, 195, 509, 1025, 938, 52]. Populationsgenetik [159]. Positive [791, 416, 50, 730, 315]. Possession [286]. Potato [891]. Potencies [1003]. Potency [622, 464, 316, 648]. Power [1042, 843, 212, 1034, 260]. Practical [154, 211]. Practice [486]. Pre [958]. Pre-1940 [958]. Precision [954, 436, 229, 284]. Predicted [335]. Prediction [1010, 865]. Preliminary [391, 753, 38, 545]. Presence [923]. Present [462]. Prevalent [323]. Principal [300, 937, 624]. principales [624]. Principes [501]. Principles [486, 501, 1021, 181, 763]. Prior [728]. priori [186]. Probabilistic [444, 972]. probabilités [710]. Probabilities [797, 707, 825, 681, 555, 910, 790, 890, 629, 873]. Probability [164, 134, 604, 256, 519, 762, 874, 680, 247, 725, 774, 675, 708, 831, 859, 567, 442, 377, 67, 116, 960, 741]. Probit [700]. Problem [173, 816, 428, 405, 764, 940, 340, 592, 936, 431]. Problems [83, 972, 446, 519, 618, 211, 901, 487]. Procedure [1043, 171, 812, 393, 261, 630, 26, 403]. Procedures [794, 479, 317, 693, 181, 336, 150, 608, 648, 857, 1020]. Proceedings [488, 489, 520]. Procese [1022]. Process [351, 409, 214, 675, 334, 821]. Processes [328, 233, 781, 713]. Processing [28]. Prodan [962, 963]. Production [236, 191, 608]. Products [477]. Professor [367]. Program [981, 1047, 198]. Programme [642]. Programmes [1010, 29]. Programming [306]. Progress [335]. Progression [149]. Properties [609]. Proportion [234, 93]. Proportions [1042, 843, 1015]. Prudnikov [304]. Psychological [858]. Public [716, 803]. Publications [721]. Publiés [448]. Pupal [840]. Purpose [470, 251, 471, 932].

Q [703]. Quadratic [285, 918]. Qualitative [83, 61]. Quality [341, 133]. Quantal [416, 554, 468, 622, 42, 401, 754, 921, 782, 315, 690, 82, 781]. Quantiles [955]. Quantitatifs [841]. Quantitative [844, 115, 587, 479, 368, 841, 341, 857, 814, 158]. Quarter [110, 189]. Quasi [683]. Quasi- [683]. Query [155, 13, 113, 11, 790, 238, 544, 951, 176, 157, 114, 216]. Question [50].

R [865, 660, 1020, 895, 607, 573, 703, 609, 443, 483, 306, 32, 181, 381, 385, 75, 577, 862, 740, 864, 304, 710, 996, 353, 659, 410]. R. [414, 225, 371, 571, 379, 376, 373]. Radiation [76]. Radio [891]. Radio-Active [891]. Radioactivity [191]. Rahman [831]. Raiffa [996]. Random [799, 261, 776]. Randomised [325, 652]. Randomization [280]. Randomness [344, 561]. Range [597, 112, 71, 778]. Rank [1046, 475, 20, 236, 282, 145, 250, 440]. Ranking [1020, 251, 693]. Rao [607, 607, 982]. Rapid [105, 890, 778, 288]. rapports [624]. Rashevsky [964]. Rate [509, 543, 7]. Rates [1032, 112, 434, 923, 249, 114, 621]. Ratio [797, 406, 475, 299, 985, 543, 106, 629, 621]. Ratios [288, 429, 105]. Raw [477]. Ray [689, 846]. Reader [517]. Readings [765, 1050]. Realized [737]. Recapture [563]. Received [447, 490, 546, 578, 611, 661, 717, 767, 806, 835, 866, 902, 967, 997, 1052]. Recessive [736]. Recherche [841]. Reciprocal [46]. Recognisable [377]. Recombination [368, 92]. Records [562, 146, 881, 107]. Recoveries [249]. Recruitment [795, 527, 281]. Recruits [563]. Rectangular [118, 646, 66, 2]. Recurrence [1044, 887]. Red [191, 628]. Reducing [218]. redundancy [888]. Reexamination [7]. Reference [52, 1038, 43, 38, 915, 181]. Regions [43]. Regression [927, 131, 70, 850, 48, 43, 585, 497, 1033, 178, 880, 39, 14, 111, 255, 152, 754, 921, 689, 846, 788, 65, 740, 156, 153, 51, 90, 262, 1015, 106, 284, 294, 410].

Regressions [21]. Regularly [750]. Reject [336]. Related [638, 1039, 248]. Relating [750, 531]. Relation [168, 108, 910]. Relations [398, 500]. Relationship [301, 1019, 914, 556, 755, 781, 624]. Relationships [948, 292, 617, 322, 882]. Relative [309, 131, 464, 237, 316, 648, 888]. relativer [888]. Relatives [397, 248, 824]. Relief [148]. Removal [562]. Removing [842]. Rényi [926]. Reparameterization [539]. Repeat [459]. Repeatability [146]. Repeated [676, 62, 191]. Repeated-Measurements [62]. Repetitions [20]. Replicates [349, 228, 110, 189]. Replication [332, 652, 709]. Replications [626]. Report [706]. Representation [529]. Representations [1019]. Reproducing [509]. Reproduction [296, 334]. Reproductive [1007]. Requirements [639]. Research [894, 506, 402, 19, 489, 323, 507, 486, 829]. Residual [423, 759, 126]. Resolution [696]. Respect [251, 235, 297]. Response [245, 1057, 554, 844, 1037, 188, 187, 407, 531, 896, 979, 42, 436, 726, 1010, 590, 754, 594, 921, 782, 215, 530]. Responses [416, 468, 884, 401, 845, 978, 315, 690]. Restricted [598, 693, 1009]. Restrictions [925]. Results [171, 689, 846, 460, 533]. Résumés [448]. Retained [1010]. Retention [881]. Retrospective [209]. Retunnelling [283]. Review [518, 990, 858, 865, 574, 482, 486, 602, 660, 705, 894, 994, 268, 199, 52, 359, 448, 1020, 115, 444, 134, 656, 180, 604, 707, 895, 860, 411, 517, 966, 116, 354, 357, 412, 446, 358, 477, 575, 519, 522, 478, 572, 479, 523, 488, 489, 520, 515, 573, 706, 713, 805, 802, 703, 711, 762, 800, 764, 803, 709, 716, 765, 766, 798, 804, 799, 829, 834, 958, 956, 961, 1049, 992, 993, 995, 989, 965, 158, 960, 352, 742, 360, 898, 900, 899, 896, 307, 741, 265, 658, 609, 443, 483, 159, 135, 306, 384, 655, 356, 712, 96, 355, 160, 32, 606, 94, 480, 484, 901]. Review [240, 33, 445, 1021, 830, 833, 863, 957, 1048, 1051, 1050, 861, 136, 181, 381, 385, 485, 75, 266, 1022, 897, 305, 577, 657, 708, 831, 832, 1023, 383, 267, 95, 743, 859, 76, 77, 576, 862, 200, 382, 487, 608, 740, 763, 864, 959, 964, 133, 304, 710, 220, 605, 996, 353, 659, 603, 410, 962, 991, 654, 610, 857, 1025, 516, 704, 442, 963, 801, 481, 995, 411]. Reviews [607, 1024]. Revolution [616]. Richard [863]. Richards [981, 1047]. Richter [832]. Ridge [9]. Ridge-Count [9]. Ridley [803]. Risk [209]. Risks [165, 309, 86, 237, 977]. Risser [710]. Robertson [523]. Robinson [758]. robusta [640]. Robustness [394, 455]. Role [506, 19, 507]. Ronald [270, 375, 367, 372, 374, 365]. Root [891]. Rouquette [385]. Routine [728, 254]. Roving [127]. Row [1045, 584]. Roy [900, 159]. Royal [522]. Rules [336, 1006]. Runs [235, 320, 415]. Rupert [801]. Russian [517, 517].

S [486, 268, 199, 115, 116, 805, 703, 766, 804, 799, 658, 135, 833, 133, 304, 605, 659, 516, 391]. S. [314]. Salkind [764]. Salmon [486]. Salvage [197]. Same [825]. Sample [1046, 942, 953, 108, 299, 68, 570, 885, 910, 792, 735, 151, 396, 567, 250, 440]. Samples [736, 534, 5, 1018, 298, 433, 318, 186, 356]. Sampling [1058, 171, 279, 724, 38, 477, 976, 458, 880, 689, 846, 26, 599, 628, 338, 426, 669, 123, 400, 1023, 515, 716, 136]. Samuel [517]. Scale [788]. Scaling [403]. Schemes [44, 426, 669]. Schick [573]. Schlaifer [996]. Schmitt [487]. School [764]. Schumann [31]. Schwartz [385, 179]. Science [602, 305, 1050]. Sciences [448, 384, 833, 863, 1048, 659, 894, 445, 1051, 181, 897]. Scientific [994]. Scientists [660]. Scoring [917]. Screening [1046, 465, 402, 317, 1036, 730, 295, 425, 558, 336, 250, 440]. Search [841]. Searle [659]. Sebastian [705]. Seber [860]. Second [616, 252]. Seed [783, 1011]. Selected [804, 1010]. Selecting [171, 467, 912]. Selection [791, 929, 496, 826, 251, 317, 335, 28, 925, 748, 1010, 941, 215, 29, 974, 678, 177, 338, 823, 919, 213, 918, 938, 538, 593]. Selective [340]. Selektion [899]. Self [872, 940, 592, 7]. Self- [872]. Self-Incompatibility [592, 7]. Self-Sterility [940]. Selfing [733]. Semiconductors [478]. Seminar [520]. Sensitivities [350]. Sensitivity [394, 754]. Sensory [292, 104]. Separation [54, 5]. September [800]. Sequence [344]. Sequences [671, 620, 218, 888]. Sequential [1046, 193, 171, 155, 884, 1006, 250, 440, 822, 192, 1020, 200]. serial [87]. Serially [229]. Series [64, 324, 787, 713, 370, 212, 172, 943]. Serology [784]. ses [624, 710]. Session [488]. set [252]. Sets [791, 1045, 674, 393, 584, 377, 276]. Several [171, 251, 322, 1032, 460]. Sewell [1025]. Sex [929, 826]. Sex-Linked [929, 826]. Sexlinked [397]. Shelled [301]. Sheps [803]. Shnol [1049]. Short [984]. si [1022]. Sib [302, 568, 938]. Sibships [61]. Sielaff [765]. Sigmoid [539]. sign [752]. Signed [475]. Significance [54, 288, 930, 406, 391, 5, 798, 370, 105, 274, 982, 381, 852, 345]. Simha [799]. Similar [394, 674]. Similarity [680]. Simple [696, 349, 557, 780, 175, 464, 985, 102, 156, 431, 276, 888]. Simulation [976, 909]. Simultaneous [948, 49, 854, 733, 801]. Single [795, 64, 527, 398, 219, 514, 254]. Single-Tagging [795, 527]. Singly [473, 298, 318]. Sir [270, 375, 372, 374]. Sites [1008]. Situation [912]. Situations [1043, 812, 651]. Size [409, 214, 299, 651, 68, 910, 792, 151, 281, 422, 29, 1006, 882, 7]. Skinnemoen [412]. Slope [108, 43, 885]. Slopes [13]. Small [54, 5, 108, 1018, 885, 457, 735, 949, 627, 346, 260]. Smirnov [926]. Smith [656, 861, 862, 740]. Smoking [556]. Sneddon [220]. Sobel [1020]. Social [236, 833, 863, 1050]. Society [180, 16, 35, 79, 98, 119, 140, 161, 183, 203, 222, 243, 271, 289, 310, 362, 387, 417, 450, 548, 580, 613, 635, 663, 686, 719, 745, 769, 808, 837, 933, 969, 999, 1054, 1027, 329, 903, 55, 522, 524, 493, 868]. Soil [436]. Solution [173, 864]. Solutions [939, 519]. Some [849, 521, 292, 83, 908, 787, 724, 337, 531, 259, 773, 539, 672, 734, 170, 813, 379, 456, 233, 102, 689, 846, 172, 89, 584, 788, 378, 784, 983, 564, 553, 192, 153, 145, 432, 128, 284, 373, 817, 186, 328, 1045, 988, 656, 265]. Son [624]. Space [405, 446, 876]. Space-Time [876]. Spaced [750]. Spaces [567]. Spacing [347, 207]. Spatial [439, 914, 342]. Spearman [557, 87]. Special [52, 1038, 43, 38, 181]. Species [173, 872, 501, 560, 235, 1009, 319]. Specified [391, 299, 68, 734]. Spectrometer [689, 846]. Spectrum [431]. Spider [628].

Split [980, 91, 348, 785]. Split-Plot [91, 348]. Sponsored [520]. Spread [924, 945, 512]. Square [72, 163, 288, 1042, 843, 424, 105, 259, 45, 457, 510, 883, 756, 252]. Square-Summing [756]. Squares [796, 508, 13, 987, 259, 261, 880, 325, 920, 126, 254, 786, 695, 276]. Stability [50, 939]. Stable [973]. Stage [467, 946, 231, 840, 336]. Stages [454, 596]. Standard [218]. Standardisation [434, 84]. Stanley [359]. State [825, 639]. States [989]. Statistic [72, 163]. Statistical [818, 118, 705, 656, 517, 891, 908, 966, 501, 560, 405, 972, 354, 575, 522, 488, 829, 958, 625, 742, 66, 988, 295, 853, 641, 305, 9, 127, 563, 133, 996, 697, 801, 813, 602, 894, 444, 766, 992, 443, 483, 135, 484, 481]. Statisticians [800]. Statistics [714, 715, 604, 886, 19, 711, 375, 960, 658, 755, 957, 181, 897, 659, 373, 616, 109, 518, 990, 482, 268, 411, 517, 607, 762, 765, 741, 265, 96, 33, 1021, 831, 859, 763, 959, 442, 660, 358, 765, 896, 94, 353]. Statistik [712]. Statistika [360]. Statistique [501, 560, 480, 710]. Statistiques [384, 95]. Statistische [159]. Status [651]. Steady [639]. Steel [181]. Steepest [85]. Stepwise [300]. Stereoscopic [625]. Sterility [940]. Sternberg [833]. Steven [14]. Stevens [178, 111]. Stochastic [283, 164, 165, 52, 67, 86, 713, 463, 531, 556, 840, 673, 334, 530, 319, 199]. Stockholm [800]. Stohastice [1022]. Stone [863]. Straight [540, 51]. Strains [461]. Strategy [1002]. Structural [948, 49]. Structure [725]. Stuart [990]. Students [523, 445, 897, 353]. Studies [842, 28, 148, 487, 209, 882, 433, 993]. Study [164, 165, 942, 1038, 67, 86, 501, 560, 1018, 976, 454, 777, 157, 956]. Studying [675, 1]. Sturt [640]. Sub [377]. Sub-Sets [377]. Subclass [568, 814]. Subclassification [842]. Subgroup [73]. Subject [554, 146]. subobscura [640]. Subpopulations [791]. Success [434]. Successive [1036]. Sufficient [321]. Suggested [727]. Sully [576, 738]. Sum [13, 145]. Summaries [760]. Summation [630]. Summing [756]. Sums [261]. Superpositions [240]. Supplement [798]. Supplies [522]. Surgical [256]. Survey [562, 458, 972, 1023]. Survey-Removal [562]. Surveys [716, 136, 396, 709]. Survie [668]. Survival [280, 555, 879, 668, 873, 667]. Susceptibles [877]. Suspension [190]. Swaroop [766]. Sylvain [705]. symbol [888]. Symbolfolgen [888]. Symbols [992]. Symmetrical [426]. Symposia [180]. Symposium [383]. Synchrony [779]. Syndrome [77]. Synthesi [796, 920]. Synthesis [4, 695]. System [468, 465, 125]. Systematic [347, 207, 400]. Systems [1038, 247].

T [52, 764, 765, 352, 265, 833, 76]. Table [164, 165, 851, 67, 86, 175, 798, 692, 985, 31, 381]. Tables [797, 844, 878, 980, 947, 473, 395, 762, 798, 135, 178, 45, 455, 890, 14, 111, 590, 629, 735, 263, 460, 210, 432, 257, 82, 194, 989, 134, 573, 799, 381, 304]. Tabulating [96]. Tag [249]. Tagging [795, 527]. Tail [43]. Taste [25, 341]. Tautu [1022]. Taxonomy [1036, 226]. Tea [396, 628]. Teacher [358].

Teaching [908, 811, 265, 988, 892]. Technique [1016, 1033, 481, 266]. Techniques [818, 4, 444, 656, 677, 1036, 901, 628, 515, 608]. Tenderometer [301]. Tenderometer-Yield [301]. Terms [517, 992]. Terrill [608]. Test [54, 288, 391, 264, 344, 475, 404, 155, 12, 196, 623, 175, 5, 683, 569, 105, 597, 112, 457, 733, 776, 1034, 198, 257, 377, 345, 71, 778, 752]. Tested [421]. Testing [930, 884, 972, 798, 393, 124, 950, 381, 949, 104, 65, 44, 545, 852, 426, 669, 697, 489]. Tests [1042, 1046, 858, 394, 292, 843, 280, 729, 135, 982, 455, 926, 853, 425, 536, 427, 145, 250, 440]. Tetrasomic [872]. Text [574]. Th [609]. Their [245, 328, 187, 586, 233]. Theoretical [398, 102]. Theoretically [154]. Theory [416, 990, 602, 27, 368, 352, 337, 401, 315, 127, 973, 107, 895, 860, 484, 708, 996, 516]. Therapeutic [465]. Therapeutique [698]. thérapeutiques [385]. There [421]. Thinking [1051]. Thomas [411, 446]. Thöni [995]. Those [640]. Three [288, 644, 398, 322, 105, 110, 189, 147, 275, 336]. Three-Dimensional [275]. Three-Part [288, 105]. Three-Quarter [110]. Three-Way [644, 398, 322]. Threshold [25]. Thurstone [8]. Tied [343]. Tied-Double-Change-Over [343]. Tierproduktion [898]. Tierzucht [900]. Tierzüchtung [899]. Ties [8]. Time [245, 405, 187, 251, 407, 789, 713, 531, 726, 229, 876, 943]. Time-Response [245, 187, 407]. Times [530]. Timing [227]. Tippett [804]. Tisserand [384]. Tisserand-Perrier [384]. Tolerances [416, 315]. Tomato [38]. Tome [710]. Torii [52]. Torrie [181]. Total [918]. Toxic [818]. Tracer [891, 23]. Track [107]. Traité [710]. Traits [251, 841]. Transect [820, 976]. Transects [415, 235, 320]. Transformation [301, 881, 932]. Transformations [788, 995]. Transformed [568]. Transition [825]. Transmission [397]. Trapping [682]. Travaux [448]. Traynard [710]. Treating [247]. Treatment [652, 46, 260]. Treatments [849, 1045, 620, 467, 218, 113, 424, 259, 651, 89, 584, 252, 786, 752]. Tree [347]. Trees [415, 297, 320]. Trends [1042, 843, 168, 175, 884, 459, 276]. Trial [347]. Trials [327, 1060, 344, 884, 946, 332, 293, 511, 983, 731, 533, 822, 192, 200]. Triangle [404]. Triple [104, 2]. Tromp [805]. Tropical [347]. true [932]. Truncated [736, 822, 473, 649, 540, 93, 848, 47, 318]. Tuberculin [84]. Tukey [72, 163]. Tunnelling [283]. Twinning [824]. Two [208, 54, 416, 794, 1045, 1046, 1060, 947, 467, 228, 699, 1019, 922, 946, 48, 5, 392, 464, 322, 877, 299, 511, 950, 434, 692, 503, 498, 733, 342, 315, 690, 584, 336, 150, 153, 697, 250, 440, 260, 815, 1046]. Two-Dimensional [947]. Two-Level [208, 392]. Two-Period [1060, 511, 815]. Two-Phase [342]. Two-Sample [1046, 250, 440]. Two-Stage [467, 946]. Two-Treatment [260]. Two-Way [794, 322, 950, 503, 498, 150]. Type [435, 458, 211, 821].

Überla [1024]. Ulrich [242]. Unbalanced [930, 875, 683, 232, 261, 219, 852]. Unbiased [774, 26, 258, 106, 284]. Underlying [470, 471]. Unequal [1035, 568, 886, 349, 20, 643, 651, 458, 1005, 814]. Unequally [474]. Unified [416, 401, 315]. Uniform [122]. United [520, 989]. Units [128, 815].

Univariate [147]. University [811, 892]. Unknown [171, 880]. Unnested [683]. Unreplicated [392]. Unusual [197, 466]. Upon [1032, 639]. Urbakh [360]. Urquhart [711]. Use [1060, 391, 728, 753, 954, 589, 347, 783, 1011, 511, 730, 186, 226, 694, 754, 784, 432, 667, 624]. Used [995, 621]. Useful [317, 594]. Using [475, 23, 8, 436, 755, 648, 778]. utilisation [624]. Utilizing [914, 558].

V [604, 799, 360, 265, 32, 901, 833]. Vaccinations [834]. Vaccine [193, 533]. Vaccines [426, 669]. Valeurs [974]. Validity [427]. Value [789, 694, 731, 538]. Values [1061, 392, 597, 307, 916, 911, 71]. Variability [64, 729, 827, 650, 777]. Variable [406, 497, 598]. Variables [791, 753, 617, 501, 560, 922, 300, 995, 941, 667]. Variance [849, 138, 492, 794, 930, 1043, 1058, 171, 644, 875, 339, 397, 676, 505, 812, 953, 300, 101, 421, 322, 393, 231, 462, 232, 68, 733, 500, 176, 215, 219, 198, 599, 150, 907, 169, 346, 430, 1040, 852, 1009, 1005]. variance-covariance [1005]. Variances [796, 1035, 333, 644, 875, 299, 950, 880, 348, 920, 695]. Variate [953]. Variates [1033, 937, 618, 817]. Variation [492, 339, 277, 41, 542, 32]. Varietal [321]. Variety [332, 638]. Veb [965]. Vectorial [640]. Vectors [234]. Végétales [501, 560]. Vegetation [342]. Verfahren [888]. Verhaltens [1024]. Verlag [965]. Vessereau [95]. Viewpoint [604]. Viral [1008]. Viruses [532]. Vital [714, 33, 991]. Vitamin [1037]. Voiding [587]. Vol [990, 307, 306, 957, 861, 1025]. Volume [58, 143, 224, 313, 420, 551, 666, 772, 906, 1030, 43, 805]. vorgegebener [888]. Vraisemblance [668].

W [180, 479, 515, 805, 762, 484, 1048, 1050, 266, 382, 605, 991]. W. [14]. Waerden [712]. Wahrscheinlichkeitstheorie [832]. Wallace [606]. Walter [242, 857, 441]. Walter-Ulrich [242]. Ward [410]. Washburne [307]. Watson [821]. Way [794, 644, 875, 398, 129, 322, 950, 503, 498, 65, 150, 786]. Wearden [179]. Weight [251]. Weight-Free [251]. Weighted [474, 880, 921]. Weighting [219]. Weihe [805]. Weinfeld [703]. Westlake [864]. Wetherill [966]. where [651]. Which [554]. WHO [834]. Width [299, 68]. Wijngaarden [609]. Wilcoxon [475, 600, 721]. Wilks [268]. William [859, 959]. Williams [610]. Wine [660]. within [791, 924, 922, 585, 570]. within- [585]. within-Sample [570]. without [46]. Wolberg [865]. Wold [713]. Work [376]. Working [851]. World [714, 715, 520, 716, 991]. Wright [871, 1025].

Yates [757, 136, 756]. year [621]. Yield [301, 750]. Yields [627]. York [703]. Youden [252, 786, 259]. Yu [360]. Yule [573].

Zea [914]. Zeitschriften [965]. Zero [93]. Zero-Truncated [93]. zur [888]. Zygosis [9].

References

Pearce:1960:MSM

[1] S. C. Pearce. A method of studying manner of growth. Biometrics, 16(1): 1–6, March 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527950.

Pasternack:1960:ACT

[2] Bernard S. Pasternack. Analysis of covariance for a 3×4 triple rectangular lattice design (3 associate P.B.I.B.). Biometrics, 16(1):7–18, March 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527951.

Andersen:1960:CPC

[3] F. Søgaard Andersen. Competition in populations consisting of one age group. Biometrics, 16(1):19–27, March 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527952.

Blackith:1960:SMT

[4] R. E. Blackith. A synthesis of multivariate techniques to distinguish patterns of growth in grasshoppers. Biometrics, 16(1):28–40, March 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527953.

Dempster:1960:STS

[5] A. P. Dempster. A significance test for the separation of two highly multivariate small samples. Biometrics, 16(1):41–50, March 1960. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527954. See correction [54].

Darwin:1960:EDA

[6] J. H. Darwin. An ecological distribution akin to Fisher’s logarithmic distribution. Biometrics, 16(1):51–60, March 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527955.

Wright:1960:NSI

[7] Sewall Wright. On the number of self-incompatibility alleles maintained in equilibrium by a given mutation rate in a population of given size: a reexamination. Biometrics, 16(1):61–85, March 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527956.

Glenn:1960:TPC

[8] W. A. Glenn and H. A. David. Ties in paired-comparison experiments using a modified Thurstone–Mosteller model. Biometrics, 16(1):86–109, March 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527957.

Richter:1960:SMD

[9] Donald L. Richter and Seymour Geisser. A statistical model for diagnosing zygosis by ridge-count. Biometrics, 16(1):110–114, March 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527958.

Shah:1960:NFD

[10] B. V. Shah. 143. Note: On a 5 × 2^{2} factorial design. Biometrics, 16(1): 115–118, March 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527959. See

correction [74].

F:1960:QCP

[11] D. J. F. 144. Query: On combining partial correlation coefficients. Biometrics, 16(1):118, March 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527960.

Chassan:1960:NTO

[12] J. B. Chassan. 145 note: On a test for order. Biometrics, 16(1):119– 121, March 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527961.

Cormack:1960:QPB

[13] R. M. Cormack. 146. Query: Partitioning the “Between Slopes” sum of squares for forest growth data. Biometrics, 16(1):121–124, March 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527962.

Linhart:1960:NTW

[14] H. Linhart. 147. Note: Tables for W. L. Steven’s “Asymptotic Regression”. Biometrics, 16(1):125, March 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527963.

Anonymous:1960:Aa

[15] Anonymous. Abstracts. Biometrics, 16(1):126–137, March 1960. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527964.

Anonymous:1960:BSa

[16] Anonymous. The Biometric Society. Biometrics, 16(1):139–150, March 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527965.

Anonymous:1960:NAa

[17] Anonymous. News and announcements. Biometrics, 16(1):151–158, March 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527966.

Anonymous:1960:FMa

[18] Anonymous. Front matter. Biometrics, 16(1):??, March 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527949.

Emmens:1960:RSP

[19] C. W. Emmens. The role of statistics in physiological research. Biometrics, 16(2):161–175, June 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527549.

Dykstra:1960:RAI

[20] Otto Dykstra, Jr. Rank analysis of incomplete block designs: a method of paired comparisons employing unequal repetitions on pairs. Biometrics, 16(2):176–188, June 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527550.

Wright:1960:PCP

[21] Sewall Wright. Path coefficients and path regressions: Alternative or complementary concepts? Biometrics, 16(2):189–202, June 1960. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527551.

Cohen:1960:EPC

[22] A. Clifford Cohen, Jr. Estimating the parameter in a conditional Poisson distribution. Biometrics, 16(2):203–211, June 1960. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527552.

Cornfield:1960:MIE

[23] Jerome Cornfield, Jesse Steinfeld, and Samuel W. Greenhouse. Models for the interpretation of experiments using tracer compounds. Biometrics, 16(2):212–234, June 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527553.

Stanton:1960:GCM

[24] R. G. Stanton. Genetic correlations with multiple alleles. Biometrics, 16(2):235–244, June 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527554.

Harris:1960:AEM

[25] Eugene K. Harris. Analysis of experiments measuring threshold taste. Biometrics, 16(2):245–260, June 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527555.

Robson:1960:USE

[26] D. S. Robson. An unbiased sampling and estimation procedure for creel censuses of fishermen. Biometrics, 16(2):261–277, June 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527556.

Chakravarti:1960:MTB

[27] N. K. Chakravarti. The mathematical theory of biological assay of a local anaesthetic. Biometrics, 16(2):278–291, June 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527557.

Hoen:1960:NCI

[28] K. Hoen and A. H. E. Grandage. 148. Note: Calculation of inbreeding in family selection studies on the IBM-650 data processing machine. Biometrics, 16(2):292–296, June 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527558.

Robertson:1960:NOF

[29] Alan Robertson. 149. Note: On optimum family size in selection programmes. Biometrics, 16(2):296–298, June 1960. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527559.

Turner:1960:NHE

[30] Malcolm E. Turner. 150. Note: On heuristic estimation methods. Biometrics, 16(2):299–301, June 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527560.

Wearden:1960:NGA

[31] Stanley Wearden. 151. Note: a genetic application of the Schumann– Bradley table. Biometrics, 16(2):301–303, June 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527561.

Karn:1960:BRG

[32] M. N. Karn. Book review: Genetic Basis of Morphological Variation, by R. H. Osborne, F. V. De George. Biometrics, 16(2):304–307, June 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527562.

Martin:1960:BRE

[33] W. J. Martin. Book review: Elements of Vital Statistics, by B. Benjamin. Biometrics, 16(2):308–309, June 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527563.

Anonymous:1960:Ab

[34] Anonymous. Abstracts. Biometrics, 16(2):310–316, June 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527564.

Anonymous:1960:BSb

[35] Anonymous. The Biometric Society. Biometrics, 16(2):317–332, June 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527565.

Anonymous:1960:NAb

[36] Anonymous. News and announcements. Biometrics, 16(2):333–335, June 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527566.

Anonymous:1960:FMb

[37] Anonymous. Front matter. Biometrics, 16(2):??, June 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527548.

Elteto:1960:PSS

[38] Ö. Éltetö and K. Sarkadi. On preliminary sampling with special reference to the microbiological control of tomato concentrates. Biometrics, 16(3):339–347, September 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527685.

Latter:1960:EDE

[39] B. D. H. Latter and Alan Robertson. Experimental design in the estimation of heritability by regression methods. Biometrics, 16(3):348–353, September 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527686.

Chapman:1960:ACC

[40] D. G. Chapman and D. S. Robson. The analysis of a catch curve. Biometrics, 16(3):354–368, September 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527687.

Hayman:1960:MLE

[41] B. I. Hayman. Maximum likelihood estimation of genetic components of variation. Biometrics, 16(3):369–381, September 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527688.

Gurland:1960:PQR

[42] John Gurland, Ilbok Lee, and Paul A. Dahm. Polychotomous quantal response in biological assay. Biometrics, 16(3):382–398, September 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527689.

Dudzinski:1960:ERS

[43] M. L. Dudzinski. Estimation of regression slope from tail regions with special reference to the volume line. Biometrics, 16(3):399–407, September 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2527690.

Smith:1960:EAT

[44] Charles Smith. Efficiency of animal testing schemes. Biometrics, 16 (3):408–415, September 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527691.

Kastenbaum:1960:NAP

[45] Marvin A. Kastenbaum. A note on the additive partitioning of chi-square in contingency tables. Biometrics, 16(3):416–422, September 1960. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527692.

Wright:1960:TRI

[46] Sewall Wright. The treatment of reciprocal interaction, with or without lag, in path analysis. Biometrics, 16(3):423–445, September 1960. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527693.

Cohen:1960:ETP

[47] A. Clifford Cohen, Jr. An extension of a truncated Poisson distribution. Biometrics, 16(3):446–450, September 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527694.

DeBaun:1960:OAR

[48] Robert M. De Baun and Victor Chew. Optimum allocation in regression experiments with two components of error. Biometrics, 16(3):451–463, September 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527695.

Basmann:1960:ENE

[49] R. L. Basmann. An expository note on estimation of simultaneous structural equations. Biometrics, 16(3):464–480, September 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527696.

Dempster:1960:NQS

[50] Everett R. Dempster. 152 note: The question of stability with positive feedback. Biometrics, 16(3):481–483, September 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527697.

Turner:1960:NSL

[51] Malcolm E. Turner. 153 note: Straight line regression through the origin. Biometrics, 16(3):483–485, September 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527698.

Blackith:1960:BRS

[52] R. E. Blackith. Book review: The Stochastic Approach in Field Population Ecology with Special Reference to Field Insect Populations by T. Torii. Biometrics, 16(3):486–487, September 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527699.

Anonymous:1960:Ac

[53] Anonymous. Abstracts. Biometrics, 16(3):488–491, September 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527700.

Anonymous:1960:CST

[54] Anonymous. Correction: A significance test for the separation of two highly multivariate small examples. Biometrics, 16(3):492, September 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527701. See [5].

Darmois:1960:BS

[55] G. Darmois. The Biometric Society. Biometrics, 16(3):493–499, September 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2527702.

Anonymous:1960:NAc

[56] Anonymous. News and announcements. Biometrics, 16(3):500–501, September 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527703.

Anonymous:1960:FMc

[57] Anonymous. Front matter. Biometrics, 16(3):??, September 1960. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527684.

Anonymous:1960:VI

[58] Anonymous. Volume information. Biometrics, 16(4):i–xvii, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527756.

Kimball:1960:EMI

[59] A. W. Kimball. Estimation of mortality intensities in animal experiments. Biometrics, 16(4):505–521, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527758.

Shumway:1960:FPB

[60] Robert Shumway and John Gurland. Fitting the Poisson binomial distribution. Biometrics, 16(4):522–533, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527759.

Gittelsohn:1960:MAD

[61] A. M. Gittelsohn. A model for the analysis of the distribution of qualitative characters in sibships. Biometrics, 16(4):534–546, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527760.

Danford:1960:ARM

[62] M. B. Danford, Harry M. Hughes, and R. C. McNee. On the analysis of repeated-measurements experiments. Biometrics, 16(4):547–565, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527761.

Bradley:1960:IIB

[63] R. A. Bradley, R. E. Walpole, and C. Y. Kramer. Intra- and inter-block analysis for factorials in incomplete block designs. Biometrics, 16(4):566– 581, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527762.

Armitage:1960:DHV

[64] P. Armitage and G. E. Bartsch. The detection of host variability in a dilution series with single observations. Biometrics, 16(4):582–592, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527763.

Robson:1960:IDF

[65] D. S. Robson and G. F. Atkinson. Individual degrees of freedom for testing homogeneity of regression coefficients in a one-way analysis of covariance. Biometrics, 16(4):593–605, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527764.

Hey:1960:SER

[66] E. N. Hey and M. H. Hey. The statistical estimation of a rectangular hyperbola. Biometrics, 16(4):606–617, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527765. See correction [118].

Chiang:1960:SSL

[67] Chin Long Chiang. A stochastic study of the life table and its applications: I. Probability distributions of the biometric functions. Biometrics, 16(4):618–635, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2527766. See corrections [164].

Graybill:1960:SSS

[68] Franklin A. Graybill and Robert D. Morrison. Sample size for a specified width confidence interval on the variance of a normal distribution. Biometrics, 16(4):636–641, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527767.

Wiggins:1960:MMM

[69] Alvin D. Wiggins. On a multicompartment migration model with chronic feeding. Biometrics, 16(4):642–658, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527768.

Buehler:1960:ARF

[70] Robert J. Buehler. An application of regression to frequency graduation. Biometrics, 16(4):659–670, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527769.

Harter:1960:CVD

[71] H. Leon Harter. Critical values for Duncan’s new multiple range test. Biometrics, 16(4):671–685, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527770.

Abraham:1960:NAM

[72] John K. Abraham. 154 note: On an alternative method of computing Tukey’s statistic for the Latin square model. Biometrics, 16(4):686– 691, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527771. See

corrections [163].

Katti:1960:NDM

[73] S. K. Katti. 155 note: a direct method for constructing the intrablock subgroup. Biometrics, 16(4):691–694, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527772.

Anonymous:1960:CNF

[74] Anonymous. Correction to note 143 “On a 5 × 2^{2} factorial design”. Biometrics, 16(4):695, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527773. See [10].

Patterson:1960:BRP

[75] H. D. Patterson. Book review: Planning of Experiments by D. R. Cox. Biometrics, 16(4):696, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527774.

S:1960:BRE

[76] J. G. S. Book review: Effects of Nuclear Radiation on Men and Materials, by T. C. Helvey. Biometrics, 16(4):697, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527775.

S:1960:BRI

[77] J. G. S. Book review: The Intervertebral Disc Syndrome, by H. G. Higley. Biometrics, 16(4):697, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527776.

Anonymous:1960:Ad

[78] Anonymous. Abstracts. Biometrics, 16(4):698, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527777.

Anonymous:1960:BSc

[79] Anonymous. The Biometric Society. Biometrics, 16(4):699–705, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527778.

Anonymous:1960:NAd

[80] Anonymous. News and announcements. Biometrics, 16(4):706–709, December 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527779.

Anonymous:1960:FMd

[81] Anonymous. Front matter. Biometrics, 16(4):??, December 1960. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527757.

Yates:1961:MPM

[82] F. Yates. Marginal percentages in multiway tables of quantal data with disproportionate frequencies. Biometrics, 17(1):1–9, March 1961. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527492.

Cochran:1961:SCP

[83] William G. Cochran and Carl E. Hopkins. Some classification problems with multivariate qualitative data. Biometrics, 17(1):10–32, March 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527493.

Stone:1961:STH

[84] M. Stone and R. A. Bruce. The standardisation of tuberculin hypersensitivity. Biometrics, 17(1):33–47, March 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527494.

Brooks:1961:OEG

[85] Samuel H. Brooks and M. Ray Mickey. Optimum estimation of gradient direction in steepest ascent experiments. Biometrics, 17(1):48–56, March 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527495.

Chiang:1961:SSL

[86] Chin Long Chiang. A stochastic study of the life table and its applications. III. The follow-up study with the consideration of competing risks. Biometrics, 17(1):57–78, March 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2527496. See corrections [165].

Johnson:1961:SES

[87] Eugene A. Johnson and Byron Wm. Brown, Jr. The Spearman estimator for serial dilution assays. Biometrics, 17(1):79–88, March 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527497.

Nelder:1961:FGL

[88] J. A. Nelder. The fitting of a generalization of the logistic curve. Biometrics, 17(1):89–110, March 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2527498. See correction [139].

Pavate:1961:CAB

[89] M. V. Pavate. Combined analysis of balanced incomplete block designs with some common treatments. Biometrics, 17(1):111–119, March 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527499.

Turner:1961:GAR

[90] Malcolm E. Turner, Robert J. Monroe, and Henry L. Lucas, Jr. Generalized asymptotic regression and non-linear path analysis. Biometrics, 17(1):120–143, March 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527500.

Harter:1961:ASP

[91] H. Leon Harter. On the analysis of split-plot experiments. Biometrics, 17(1):144–149, March 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527501.

Kimball:1961:NCI

[92] A. W. Kimball. 156. Note: Confidence intervals for recombination experiments with microorganisms. Biometrics, 17(1):150–153, March 1961.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527502.

Wilkinson:1961:NEP

[93] G. N. Wilkinson. 157. Note: Estimation of proportion from zerotruncated binomial data. Biometrics, 17(1):153–159, March 1961. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527503.

Kimball:1961:BRS

[94] B. F. Kimball. Book review: Statistics of Extremes, by E. J. Gumbel. Biometrics, 17(1):160–161, March 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527504.

Robinson:1961:BRM

[95] P. Robinson. Book review: Méthodes Statistiques en Biologie et en Agronomie, by A. Vessereau. Biometrics, 17(1):162–163, March 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527505.

Holland:1961:BRT

[96] W. W. Holland. Book review: Tabulating Equipment and Army Medical Statistics, by A. G. Love, E. L. Hamilton, I. L. Hellman. Biometrics, 17 (1):163, March 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527506.

Anonymous:1961:Aa

[97] Anonymous. Abstracts. Biometrics, 17(1):164–189, March 1961. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527507.

Anonymous:1961:BSa

[98] Anonymous. The Biometric Society. Biometrics, 17(1):190–195, March 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527508.

Anonymous:1961:NAa

[99] Anonymous. News and announcements. Biometrics, 17(1):196–204, March 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527509.

Anonymous:1961:FMa

[100] Anonymous. Front matter. Biometrics, 17(1):??, March 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527491.

Elston:1961:AAV

[101] Robert C. Elston. On additivity in the analysis of variance. Biometrics, 17(2):209–219, June 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2527987. See correction [138].

Nelder:1961:NSG

[102] J. A. Nelder. A note on some growth patterns in a simple theoretical organism. Biometrics, 17(2):220–228, June 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527988.

Kempthorne:1961:PDC

[103] O. Kempthorne and R. N. Curnow. The partial diallel cross. Biometrics, 17(2):229–250, June 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2527989. See corrections [182].

Park:1961:STT

[104] G. T. Park. Sensory testing by triple comparisons. Biometrics, 17(2): 251–260, June 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527990.

Gates:1961:RCS

[105] Charles E. Gates and Benjamin H. Beard. Rapid chi-square test of significance for three-part ratios. Biometrics, 17(2):261–266, June 1961. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527991. See correction [288].

Williams:1961:GUR

[106] W. H. Williams. Generating unbiased ratio and regression estimators. Biometrics, 17(2):267–274, June 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527992.

Turner:1961:BTM

[107] Malcolm E. Turner and Eleanor D. Campbell. A biometric theory of middle and long distance track records. Biometrics, 17(2):275–282, June 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527993.

Dorff:1961:SSB

[108] Martin Dorff and John Gurland. Small sample behavior of slope estimators in a linear functional relation. Biometrics, 17(2):283–298, June 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527994.

vanWoerkom:1961:SDM

[109] Adrianus J. van Woerkom and Keeve Brodman. Statistics for a diagnostic model. Biometrics, 17(2):299–318, June 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527995.

John:1961:NTQ

[110] Peter W. M. John. 159 note: Three-quarter replicates of 2 4 and 2 5 designs. Biometrics, 17(2):319–321, June 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527996.

Lipton:1961:NES

[111] S. Lipton. 160 note: On the extension of Stevens’ tables for asymptotic regression. Biometrics, 17(2):321, June 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527997.

Harter:1961:NCE

[112] H. Leon Harter. 161 note: Corrected error rates for Duncan’s new multiple range test. Biometrics, 17(2):321–324, June 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527998.

Dunnett:1961:QMC

[113] C. W. Dunnett. 162. Query: Multiple comparisons between treatments and a control. Biometrics, 17(2):324–326, June 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527999.

Steel:1961:QER

[114] R. G. D. Steel. 163. Query: Error rates in multiple comparisons. Biometrics, 17(2):326–328, June 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528000.

Bodmer:1961:BRI

[115] W. F. Bodmer. Book review: Introduction to Quantitative Genetics, by D. S. Falconer. Biometrics, 17(2):329–330, June 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528001.

David:1961:BRP

[116] F. N. David. Book review: Probability: an Introduction, by S. Goldberg. Biometrics, 17(2):330, June 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528002.

Anonymous:1961:Ab

[117] Anonymous. Abstracts. Biometrics, 17(2):331–333, June 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528003.

Anonymous:1961:CSE

[118] Anonymous. Correction: The statistical estimation of a rectangular hyperbola. Biometrics, 17(2):333, June 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528004. See [66].

Anonymous:1961:BSb

[119] Anonymous. The Biometric Society. Biometrics, 17(2):334–339, June 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528005.

Anonymous:1961:NAb

[120] Anonymous. News and announcements. Biometrics, 17(2):340–343, June 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528006.

Anonymous:1961:FMb

[121] Anonymous. Front matter. Biometrics, 17(2):??, June 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527986.

Williams:1961:GAD

[122] E. J. Williams. The growth and age-distribution of a population of insects under uniform conditions. Biometrics, 17(3):349–358, September 1961.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527830.

VanVleck:1961:ESE

[123] L. D. VanVleck and C. R. Henderson. Empirical sampling estimates of genetic correlations. Biometrics, 17(3):359–371, September 1961. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527831.

Grizzle:1961:NMT

[124] James E. Grizzle. A new method of testing hypotheses and estimating parameters for the logistic model. Biometrics, 17(3):372–385, September 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527832.

Mode:1961:GMH

[125] C. J. Mode. A generalized model of a host–pathogen system. Biometrics, 17(3):386–404, September 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527833.

Sheehe:1961:LSB

[126] Paul R. Sheehe and Irwin D. J. Bross. Latin squares to balance immediate residual, and other order, effects. Biometrics, 17(3):405–414, September 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527834.

Robson:1961:STR

[127] D. S. Robson. On the statistical theory of a roving creel census of fishermen. Biometrics, 17(3):415–437, September 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527835.

Williams:1961:AEU

[128] J. S. Williams. Allocation of experimental units in some elementary balanced designs. Biometrics, 17(3):438–446, September 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527836.

Federer:1961:ADO

[129] Walter T. Federer. Augmented designs with one-way elimination of heterogeneity. Biometrics, 17(3):447–473, September 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527837.

Searle:1961:PGE

[130] S. R. Searle. Phenotypic, genetic and environmental correlations. Biometrics, 17(3):474–480, September 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527838.

Bohren:1961:REH

[131] B. B. Bohren, H. E. McKean, and Yukio Yamada. Relative efficiencies of heritability estimates based on regression of offspring on parent. Biometrics, 17(3):481–491, September 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527839.

Gomes:1961:NFE

[132] F. Pimentel Gomes. 164 note: On a formula for the estimation of the optimum dressing of a fertilizer. Biometrics, 17(3):492–494, September 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527840.

Shenton:1961:BRS

[133] L. R. Shenton. Book review: The Statistical Basis of Quality Control Charts, by S. K. Ekambaram. Biometrics, 17(3):495, September 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527841.

Bradley:1961:BRT

[134] R. A. Bradley. Book review: Tables of the Hypergeometric Probability Distribution, by Gerald J. Lieberman, Donald B. Owen. Biometrics, 17(3):495–496, September 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527842.

Healy:1961:BRS

[135] M. J. R. Healy. Book review: Statistical Tables for Tests of Multivariate Hypotheses, by K. C. S. Pillai. Biometrics, 17(3):496, September 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527843.

Mountford:1961:BRS

[136] M. D. Mountford. Book review: Sampling Methods for Censuses and Surveys. 3rd Edition, by F. Yates. Biometrics, 17(3):497, September 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527844.

Anonymous:1961:Ac

[137] Anonymous. Abstracts. Biometrics, 17(3):498–509, September 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527845.

Anonymous:1961:CAA

[138] Anonymous. Correction: On additivity in the analysis of variance. Biometrics, 17(3):509, September 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2527847. See [101].

Anonymous:1961:CFG

[139] Anonymous. Correction: The fitting of a generalization of the logistic curve. Biometrics, 17(3):509, September 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527846. See [88].

Anonymous:1961:BSc

[140] Anonymous. The Biometric Society. Biometrics, 17(3):510–519, September 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2527848.

Linder:1961:NA

[141] A. Linder, H. L. LeRoy, S. Sagoroff, and L. Schmetterer. News and announcements. Biometrics, 17(3):520–521, September 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527849.

Anonymous:1961:FMc

[142] Anonymous. Front matter. Biometrics, 17(3):??, September 1961. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527829.

Anonymous:1961:VI

[143] Anonymous. Volume information. Biometrics, 17(4):i–xviii, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527851.

Katti:1961:PPD

[144] S. K. Katti and John Gurland. The Poisson Pascal distribution. Biometrics, 17(4):527–538, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527853.

Steel:1961:SRS

[145] Robert G. D. Steel. Some rank sum multiple comparisons tests. Biometrics, 17(4):539–552, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527854.

Curnow:1961:ERH

[146] R. N. Curnow. The estimation of repeatability and heritability from records subject to culling. Biometrics, 17(4):553–566, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527855.

Khatri:1961:TCU

[147] C. G. Khatri and I. R. Patel. Three classes of univariate discrete distributions. Biometrics, 17(4):567–575, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527856.

Meier:1961:FCM

[148] Paul Meier and Spencer M. Free, Jr. Further consideration of methodology in studies of pain relief. Biometrics, 17(4):576–583, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527857.

Williams:1961:FGP

[149] E. J. Williams. Fitting a geometric progression to frequencies. Biometrics, 17(4):584–606, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527858.

Searle:1961:CPE

[150] S. R. Searle and C. R. Henderson. Computing procedures for estimating components of variance in the two-way classification, mixed model. Biometrics, 17(4):607–616, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2527859. See correction [794].

Nordskog:1961:OSS

[151] A. W. Nordskog, H. T. David, and H. B. Eisenberg. Optimum sample size in animal disease control. Biometrics, 17(4):617–625, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527860.

McKean:1961:NAR

[152] H. E. McKean and B. B. Bohren. Numerical aspects of the regression of offspring on parent. Biometrics, 17(4):626–633, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527861.

Sprent:1961:SHC

[153] P. Sprent. Some hypotheses concerning two phase regression lines. Biometrics, 17(4):634–645, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527862.

Cox:1961:NPA

[154] C. P. Cox. 165. Note: a practical application of a theoretically inefficient design. Biometrics, 17(4):646–649, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527863.

Bross:1961:QGS

[155] I. D. J. Bross. 166. Query: On a graphical sequential test. Biometrics, 17(4):649–651, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527864.

Shah:1961:NSM

[156] B. K. Shah. 167. Note: a simple method of fitting the regression curve y = α + δx + βρx. Biometrics, 17(4):651–653, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527865.

Sampford:1961:QFS

[157] M. R. Sampford. 168. Query: On a follow-up study of the growth of children. Biometrics, 17(4):653–656, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527866.

Finney:1961:BRQ

[158] D. J. Finney. Book review: Quantitative Methods in Pharmacology, by H. De Jonge. Biometrics, 17(4):657–658, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527867.

Harte:1961:BRS

[159] C. Harte. Book review: Statistische Methoden der Populationsgenetik, by H. L. Le Roy. Biometrics, 17(4):658–659, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527868.

Jowett:1961:BRN

[160] G. H. Jowett. Book review: Numbers from Experiments, by C. C. Li. Biometrics, 17(4):659–660, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527869.

Anonymous:1961:BSd

[161] Anonymous. The Biometric Society. Biometrics, 17(4):661–666, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527870.

Anonymous:1961:NAc

[162] Anonymous. News and announcements. Biometrics, 17(4):667–668, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527871.

Anonymous:1961:CNA

[163] Anonymous. Corrections: 154. Note: On an alternative method of computing Tukey’s statistic for the Latin square model. Biometrics, 17(4): 669, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527874. See

[72].

Anonymous:1961:CSSa

[164] Anonymous. Corrections: A stochastic study of the life table and its applications: I. probability distributions of the biometric functions. Biometrics, 17(4):669, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2527872. See [67].

Anonymous:1961:CSSb

[165] Anonymous. Corrections: A stochastic study of the life table and its applications: III. the follow-up study with the consideration of competing risks. Biometrics, 17(4):669, December 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527873. See [86].

Anonymous:1961:FMd

[166] Anonymous. Front matter. Biometrics, 17(4):??, December 1961. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527852.

Reyment:1962:OHC

[167] R. A. Reyment. Observations on homogeneity of covariance matrices in paleontologic biometry. Biometrics, 18(1):1–11, March 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527706. See corrections [201].

Cox:1962:RBC

[168] C. Philip Cox. The relation between covariance and individual curvature analyses of experiments with background trends. Biometrics, 18(1):12– 21, March 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527707.

Smith:1962:MAV

[169] H. Smith, R. Gnanadesikan, and J. B. Hughes. Multivariate analysis of variance (MANOVA). Biometrics, 18(1):22–41, March 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527708.

Katti:1962:SME

[170] S. K. Katti and John Gurland. Some methods of estimation for the Poisson binomial distribution. Biometrics, 18(1):42–51, March 1962. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527709.

Bechhofer:1962:SMD

[171] Robert E. Bechhofer and Saul Blumenthal. A sequential multipledecision procedure for selecting the best one of several normal populations with a common unknown variance, II: Monte Carlo sampling results and new computing formulae. Biometrics, 18(1):52–67, March 1962. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527710.

Patil:1962:SME

[172] G. P. Patil. Some methods of estimation for the logarithmic series distribution. Biometrics, 18(1):68–75, March 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527711.

Barnett:1962:MCS

[173] V. D. Barnett. The Monte Carlo solution of a competing species problem. Biometrics, 18(1):76–103, March 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527712.

Cornell:1962:MFL

[174] Richard G. Cornell. A method for fitting linear combinations of exponentials. Biometrics, 18(1):104–113, March 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527713.

Daly:1962:NST

[175] C. Daly. 169. Note: a simple test for trends in a contingency table. Biometrics, 18(1):114–119, March 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527714.

Plackett:1962:QVC

[176] R. L. Plackett. 170. Query: On variance components and confounding. Biometrics, 18(1):119–120, March 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527715.

Tallis:1962:NSI

[177] G. M. Tallis. 171. Note: a selection index for optimum genotype. Biometrics, 18(1):120–122, March 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527716.

Hiorns:1962:NFN

[178] R. W. Hiorns. 172. Note: a further note on the extension of Stevens’ tables for asymptotic regression. Biometrics, 18(1):123, March 1962.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527717.

James:1962:NSW

[179] J. W. James. 173. Note: On Schwartz and Wearden’s method of estimating heritability. Biometrics, 18(1):123–125, March 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527718.

Brian:1962:BRM

[180] M. V. Brian. Book review: Models and Analogues in Biology — Symposia of the Society for Experimental Biology by J. W. L. Beament. Biometrics, 18(1):126, March 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527719.

Mountford:1962:BRP

[181] M. D. Mountford. Book review: Principles and Procedures of Statistics with Special Reference to the Biological Sciences by R. G. D. Steel, J. H. Torrie. Biometrics, 18(1):127, March 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527720.

Anonymous:1962:CPD

[182] Anonymous. Corrections: The partial diallel cross. Biometrics, 18(1): 128, March 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527721. See [103].

Anonymous:1962:BSa

[183] Anonymous. The Biometric Society. Biometrics, 18(1):129–132, March 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527722.

Anonymous:1962:NAa

[184] Anonymous. News and announcements. Biometrics, 18(1):133–138, March 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527723.

Anonymous:1962:FMa

[185] Anonymous. Front matter. Biometrics, 18(1):??, March 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527705.

Katti:1962:USP

[186] S. K. Katti. Use of some a priori knowledge in the estimation of means from double samples. Biometrics, 18(2):139–147, June 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527452.

Elston:1962:ETR

[187] R. C. Elston and J. E. Grizzle. Estimation of time-response curves and their confidence bands. Biometrics, 18(2):148–159, June 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527453. See correction [245].

Bose:1962:RMC

[188] R. C. Bose and R. L. Carter. Response model coefficients and the individual degrees of freedom of a factorial design. Biometrics, 18(2):160–171, June 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2527454.

John:1962:TQR

[189] Peter W. M. John. Three quarter replicates of 2^{n} designs. Biometrics, 18(2):172–184, June 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527455.

Yassky:1962:MKP

[190] D. Yassky. A model for the kinetics of phage attachment to bacteria in suspension. Biometrics, 18(2):185–191, June 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527456.

Jowett:1962:MME

[191] G. H. Jowett. A mathematical method of estimating the production of red blood cells by repeated injections of radioactivity. Biometrics, 18(2): 192–202, June 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527457.

Spicer:1962:SNC

[192] C. C. Spicer. Some new closed sequential designs for clinical trials. Biometrics, 18(2):203–211, June 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527458.

Atkinson:1962:SEV

[193] G. F. Atkinson and D. S. Robson. Sequential estimation of vaccine effectiveness. Biometrics, 18(2):212–223, June 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527459.

Kincaid:1962:CCT

[194] W. M. Kincaid. The combination of 2 × m contingency tables. Biometrics, 18(2):224–228, June 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527460.

Rawlings:1962:ADC

[195] J. O. Rawlings and C. Clark Cockerham. Analysis of double cross hybrid populations. Biometrics, 18(2):229–244, June 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527461.

Chassan:1962:NET

[196] J. B. Chassan. 174. Note: An extension of a test for order. Biometrics, 18(2):245–247, June 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527462.

Finney:1962:NUS

[197] D. J. Finney. 175. Note: An unusual salvage operation. Biometrics, 18(2): 247–250, June 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527463.

Roof:1962:NIC

[198] Sharon D. Roof and George D. Frazier. 176. Note: An IBM 704 computer program for computing Bartlett’s test of homogeneity of variance. Biometrics, 18(2):251–252, June 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527464.

Bharucha-Reid:1962:BRS

[199] A. T. Bharucha-Reid. Book review: Stochastic Population Models in Ecology and Epidemiology, by M. S. Bartlett. Biometrics, 18(2):253– 254, June 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527465.

Schneiderman:1962:BRS

[200] M. A. Schneiderman. Book review: Sequential Medical Trials, by P. Armitage. Biometrics, 18(2):254–255, June 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527466.

Anonymous:1962:COH

[201] Anonymous. Corrections: Observations on homogeneity of covariance matrices in paleontologic biometry. Biometrics, 18(2):256, June 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527467. See [167].

Anonymous:1962:Aa

[202] Anonymous. Abstracts. Biometrics, 18(2):257–268, June 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527468.

Anonymous:1962:BSb

[203] Anonymous. The Biometric Society. Biometrics, 18(2):269–278, June 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527469.

Anonymous:1962:NAb

[204] Anonymous. News and announcements. Biometrics, 18(2):279–280, June 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527470.

Anonymous:1962:FMb

[205] Anonymous. Front matter. Biometrics, 18(2):??, June 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527451.

Anonymous:1962:FMc

[206] Anonymous. Front matter. Biometrics, 18(3):281–282, September 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527472.

Nelder:1962:NKS

[207] J. A. Nelder. New kinds of systematic designs for spacing experiments. Biometrics, 18(3):283–307, September 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527473.

Addelman:1962:AFP

[208] Sidney Addelman. Augmenting factorial plans to accommodate additional two-level factors. Biometrics, 18(3):308–322, September 1962. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527474.

Sheehe:1962:DRA

[209] Paul R. Sheehe. Dynamic risk analysis in retrospective matched pair studies of disease. Biometrics, 18(3):323–341, September 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527475.

Tallis:1962:MLE

[210] G. M. Tallis. The maximum likelihood estimation of correlation from contingency tables. Biometrics, 18(3):342–353, September 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527476.

Martin:1962:ANT

[211] Donald C. Martin and S. K. Katti. Approximations to the Neyman Type A distribution for practical problems. Biometrics, 18(3):354–364, September 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527477.

Patil:1962:HCE

[212] G. P. Patil. On homogeneity and combined estimation for the generalized power series distribution and certain applications. Biometrics, 18(3):365– 374, September 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527478.

Williams:1962:ESI

[213] J. S. Williams. The evaluation of a selection index. Biometrics, 18 (3):375–393, September 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527479.

Bishir:1962:MPS

[214] John Bishir. Maximum population size in a branching process. Biometrics, 18(3):394–403, September 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527480.

Prout:1962:NEV

[215] Timothy Prout. 177. Note: The error variance of the heritability estimate obtained from selection response. Biometrics, 18(3):404–407, September 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527481.

Wilkinson:1962:Q

[216] G. N. Wilkinson. 178. Query. Biometrics, 18(3):407–409, September 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527482.

Finney:1962:EN

[217] D. J. Finney. Editorial note. Biometrics, 18(3):409, September 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527483.

Curnow:1962:NAC

[218] R. N. Curnow and E. Sharpe. 179. Note: The analysis of covariance as a means of reducing standard errors in certain experiments involving sequences of treatments. Biometrics, 18(3):410–413, September 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527484.

Robertson:1962:NWE

[219] Alan Robertson. 180. Note: Weighting in the estimation of variance components in the unbalanced single classification. Biometrics, 18(3):413– 417, September 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527485.

Skellam:1962:BRI

[220] J. G. Skellam. Book review: An Introduction to the Mathematics of Medicine and Biology, by J. G. Defares, I. N. Sneddon. Biometrics, 18(3):418–419, September 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527486.

Anonymous:1962:Ab

[221] Anonymous. Abstracts. Biometrics, 18(3):420–429, September 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527487.

Anonymous:1962:BSc

[222] Anonymous. The Biometric Society. Biometrics, 18(3):430–433, September 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2527488.

Anonymous:1962:NAc

[223] Anonymous. News and announcements. Biometrics, 18(3):434–436, September 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527489.

Anonymous:1962:VI

[224] Anonymous. Volume information. Biometrics, 18(4):i–xxi, December 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527891.

Bliss:1962:RFA

[225] C. I. Bliss, J. A. Fraser Roberts, F. Yates, D. J. Finney, and A. R. G. Owen. R. A. Fisher. Appreciations. Biometrics, 18(4):437–454, December 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527893.

Lubischew:1962:UDF

[226] Alexander A. Lubischew. On the use of discriminant functions in taxonomy. Biometrics, 18(4):455–477, December 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527894.

Cooke:1962:THH

[227] D. Cooke. The timing of harvests in horticultural experiments. Biometrics, 18(4):478–498, December 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527895.

Corsten:1962:BBD

[228] L. C. A. Corsten. Balanced block designs with two different numbers of replicates. Biometrics, 18(4):499–519, December 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527896.

McHugh:1962:EPT

[229] Richard B. McHugh and Francis J. Wall. Estimating the precision of time period effects in longitudinal models with serially correlated and heterogeneous errors. Biometrics, 18(4):520–528, December 1962. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527897.

S:1962:EN

[230] M. R. S. Editorial note. Biometrics, 18(4):529, December 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527898.

Gates:1962:AVS

[231] Charles E. Gates and Cherng-Jiann Shiue. The analysis of variance of the S-stage hierarchal classification. Biometrics, 18(4):529–536, December 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527899.

Gower:1962:VCE

[232] J. C. Gower. Variance component estimation for unbalanced hierarchical classifications. Biometrics, 18(4):537–542, December 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527900.

Mode:1962:SMD

[233] C. J. Mode. Some multi-dimensional birth and death processes and their applications in population genetics. Biometrics, 18(4):543–567, December 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2527901. See correction

[328].

Thompson:1962:EPV

[234] Keith H. Thompson. Estimation of the proportion of vectors in a natural population of insects. Biometrics, 18(4):568–578, December 1962. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527902.

Pielou:1962:ROS

[235] E. C. Pielou. Runs of one species with respect to another in transects through plant populations. Biometrics, 18(4):579–593, December 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527903.

James:1962:MIS

[236] J. W. James. On measuring the influence of social rank on production. Biometrics, 18(4):594–600, December 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527904.

Gart:1962:CRR

[237] John J. Gart. On the combination of relative risks. Biometrics, 18 (4):601–610, December 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2527905. See corrections [309].

Mantel:1962:Q

[238] Nathan Mantel. 181. Query. Biometrics, 18(4):611–614, December 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527906.

Nelder:1962:NAF

[239] J. A. Nelder. 182. Note: an alternative form of a generalized logistic equation. Biometrics, 18(4):614–616, December 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527907.

Mallows:1962:BRD

[240] C. L. Mallows. Book review: Decomposition of Superpositions of Distribution Functions, by P. Medgyessy. Biometrics, 18(4):617, December 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527908.

Anonymous:1962:Ac

[241] Anonymous. Abstracts. Biometrics, 18(4):618–633, December 1962. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527909.

Heinisch:1962:ODW

[242] O. Heinisch. Obituary: Dr. Walter-Ulrich Behrens. Biometrics, 18(4): 634, December 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527910.

Anonymous:1962:BSd

[243] Anonymous. The Biometric Society. Biometrics, 18(4):635–636, December 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527911.

Anonymous:1962:NAd

[244] Anonymous. News and announcements. Biometrics, 18(4):637–643, December 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527912.

Anonymous:1962:CET

[245] Anonymous. Correction: Estimation of time-response curves and their confidence bands. Biometrics, 18(4):643, December 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527913. See [187].

Anonymous:1962:FMd

[246] Anonymous. Front matter. Biometrics, 18(4):??, December 1962. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527892.

Kimura:1963:PMT

[247] Motoo Kimura. A probability method for treating inbreeding systems, especially with linked genes. Biometrics, 19(1):1–17, March 1963. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527569.

Willham:1963:CBR

[248] R. L. Willham. The covariance between relatives for characters composed of components contributed by related individuals. Biometrics, 19(1):18– 27, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527570.

Paulik:1963:EMR

[249] G. J. Paulik. Estimates of mortality rates from tag recoveries. Biometrics, 19(1):28–57, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527571.

Wilcoxon:1963:TST

[250] Frank Wilcoxon, L. J. Rhodes, and Ralph A. Bradley. Two sequential two-sample grouped rank tests with applications to screening experiments. Biometrics, 19(1):58–77 + 79–84, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527572. See corrections [1046].

Elston:1963:WFI

[251] R. C. Elston. A weight-free index for the purpose of ranking or selection with respect to several traits at a time. Biometrics, 19(1):85–97, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527573.

Clarke:1963:SST

[252] G. M. Clarke. A second set of treatments in a Youden square design. Biometrics, 19(1):98–104, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527574.

Hinkelmann:1963:COI

[253] Klaus Hinkelmann. A commonly occurring incomplete multiple classification model. Biometrics, 19(1):105–117, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527575.

Tootill:1963:RLS

[254] J. P. R. Tootill. Routine least squares estimation from models containing a single non-linearity. Biometrics, 19(1):118–143, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527576.

Lipton:1963:MLE

[255] S. Lipton and C. McGilchrist. Maximum likelihood estimators of parameters in double exponential regression. Biometrics, 19(1):144–151, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527577.

Eisenberg:1963:GPM

[256] Herbert B. Eisenberg, Randolph R. M. Geoghagen, and John E. Walsh. A general probability model for binomial events, with application to surgical mortality. Biometrics, 19(1):152–157, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527578.

Walsh:1963:LTE

[257] John E. Walsh. Loss in test efficiency due to misclassification for 2 × 2 tables. Biometrics, 19(1):158–162, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527579.

Seshadri:1963:CUE

[258] V. Seshadri. Combining unbiased estimators. Biometrics, 19(1):163–170, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527580.

Giri:1963:NCA

[259] N. C. Giri. 183 note: On the combined analysis of Youden squares and of Latin square designs with some common treatments. Biometrics, 19(1): 171–174, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527581.

Youkeles:1963:NLP

[260] Lee Herrera Youkeles. 184 note: Loss of power through ineffective pairing of observations in small two-treatment all-or-none experiments. Biometrics, 19(1):175–180, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527582.

Hickman:1963:NAP

[261] C. G. Hickman. 185 note: Abbreviated procedure for deriving expectations of sums of squares for random models of unbalanced designs. Biometrics, 19(1):180–183, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527583.

Turner:1963:NME

[262] Malcolm E. Turner. 186 note: Mean exponential regression. Biometrics, 19(1):183–186, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527584.

Newell:1963:NMT

[263] D. J. Newell. 187 note: Misclassification in 2×2 tables. Biometrics, 19(1): 187–188, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527585.

Bartholomew:1963:NCT

[264] D. J. Bartholomew. 188 note: On Chassan’s test for order. Biometrics, 19(1):188–191, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527586.

Greenberg:1963:BRS

[265] B. G. Greenberg. Book review: Some Papers Concerning the Teaching of Statistics, by T. A. Bancroft, D. V. Huntsberger. Biometrics, 19(1):192– 193, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527587.

Patterson:1963:BRF

[266] H. D. Patterson. Book review: Field Plot Technique, by E. L. LeClerg, W. H. Leonard, A. G. Clark. Biometrics, 19(1):193, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527588.

Reyment:1963:BRB

[267] R. A. Reyment. Book review: Biometria, by N. A. Plochinsky. Biometrics, 19(1):193–194, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527589.

Barton:1963:BRM

[268] D. E. Barton. Book review: Mathematical Statistics, by S. S. Wilks. Biometrics, 19(1):194–196, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527590.

Anonymous:1963:Aa

[269] Anonymous. Abstracts. Biometrics, 19(1):197–199, March 1963. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527591.

Cornish:1963:SRA

[270] E. A. Cornish. Sir Ronald Aylmer Fisher, F.R.S. Activities in Australia, 1959–62. Biometrics, 19(1):200–201, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527592.

Anonymous:1963:BSa

[271] Anonymous. The Biometric Society. Biometrics, 19(1):202–206, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527593.

Anonymous:1963:NAa

[272] Anonymous. News and announcements. Biometrics, 19(1):207–212, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527594.

Anonymous:1963:FMa

[273] Anonymous. Front matter. Biometrics, 19(1):??, March 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527568.

Gridgeman:1963:SAP

[274] N. T. Gridgeman. Significance and adjustment in paired comparisons. Biometrics, 19(2):213–228, June 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527811.

Potthoff:1963:TDI

[275] Richard F. Potthoff. Three-dimensional incomplete block designs for interaction models. Biometrics, 19(2):229–263, June 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527812.

Ogilvie:1963:SME

[276] John C. Ogilvie. A simple method for the elimination of individual trends in the analysis of balanced sets of Latin squares. Biometrics, 19(2):264– 272, June 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527813.

Freeman:1963:CEE

[277] G. H. Freeman. The combined effect of environmental and plant variation. Biometrics, 19(2):273–277, June 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527814.

Fyfe:1963:PDC

[278] J. L. Fyfe and N. Gilbert. Partial diallel crosses. Biometrics, 19(2):278– 286, June 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527815.

Curnow:1963:SDC

[279] R. N. Curnow. Sampling the diallel cross. Biometrics, 19(2):287–306, June 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2527816.

Cox:1963:RTC

[280] D. F. Cox and O. Kempthorne. Randomization tests for comparing survival curves. Biometrics, 19(2):307–317, June 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527817.

Parker:1963:EPS

[281] Richard A. Parker. On the estimation of population size, mortality, and recruitment. Biometrics, 19(2):318–323, June 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527818.

Mantel:1963:ABR

[282] Nathan Mantel and Max Halperin. Analyses of birth-rank data. Biometrics, 19(2):324–340, June 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527819.

Ahmed:1963:SMT

[283] Mohamad Salahuddin Ahmed. A stochastic model for the tunnelling and retunnelling of the flour beetle. Biometrics, 19(2):341–351, June 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527820.

Williams:1963:PSU

[284] W. H. Williams. The precision of some unbiased regression estimators. Biometrics, 19(2):352–361, June 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527821.

Healy:1963:NFQ

[285] M. J. R. Healy. 189. Note: Fitting a quadratic. Biometrics, 19(2):362– 363, June 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527822.

Macnaughton-Smith:1963:NCI

[286] P. Macnaughton-Smith. 190. Note: The classification of individuals by the possession of attributes associated with a criterion. Biometrics, 19(2): 364–366, June 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527823.

Anonymous:1963:Ab

[287] Anonymous. Abstracts. Biometrics, 19(2):367–369, June 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527824.

Anonymous:1963:CRC

[288] Anonymous. Correction: Rapid chi-square test of significance for threepart ratios. Biometrics, 19(2):369, June 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527825. See [105].

Anonymous:1963:BSb

[289] Anonymous. The Biometric Society. Biometrics, 19(2):370–382, June 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527826.

Anonymous:1963:NAb

[290] Anonymous. News and announcements. Biometrics, 19(2):383–384, June 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527827.

Anonymous:1963:FMb

[291] Anonymous. Front matter. Biometrics, 19(2):??, June 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527810.

Bradley:1963:SRA

[292] Ralph A. Bradley. Some relationships among sensory difference tests. Biometrics, 19(3):385–397, September 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527928.

Gridgeman:1963:MTD

[293] N. T. Gridgeman. Matching trials by double criteria. Biometrics, 19 (3):398–405, September 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2527929. See correction [327].

Turner:1963:GKR

[294] Malcolm E. Turner, Robert J. Monroe, and Louis D. Homer. Generalized kinetic regression analysis: Hypergeometric kinetics. Biometrics, 19(3):406–428, September 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527930.

King:1963:SDD

[295] E. P. King. A statistical design for drug screening. Biometrics, 19(3):429– 440, September 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527931.

Ohlsen:1963:FMP

[296] Sally Ohlsen. Further models for phage reproduction in a bacterium. Biometrics, 19(3):441–449, September 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527932.

Pielou:1963:DDT

[297] E. C. Pielou. The distribution of diseased trees with respect to healthy ones in a patchily infected forest. Biometrics, 19(3):450–459, September 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527933.

Winer:1963:EPI

[298] P. Winer. The estimation of the parameters of the iterated exponential distribution from singly censored samples. Biometrics, 19(3):460–464, September 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527934.

Goldman:1963:SSS

[299] Aaron Goldman. Sample size for a specified width confidence interval on the ratio of variances from two independent normal populations. Biometrics, 19(3):465–477, September 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527935.

Dempster:1963:SMA

[300] A. P. Dempster. Stepwise multivariate analysis of variance based on principal variables. Biometrics, 19(3):478–490, September 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527936.

Berry:1963:NTT

[301] G. Berry. 191. Note: a transformation for the tenderometer-yield relationship in shelled peas. Biometrics, 19(3):491–494, September 1963.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527937.

Bogyo:1963:NEC

[302] Thomas P. Bogyo and Walter A. Becker. 192. Note: Exact confidence intervals for genetic heritability estimated from paternal half-sib correlations. Biometrics, 19(3):494–496, September 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527938.

Jolicoeur:1963:NMG

[303] Pierre Jolicoeur. 193. Note: The multivariate generalization of the allometry equation. Biometrics, 19(3):497–499, September 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527939.

Shenton:1963:BRT

[304] L. R. Shenton. Book review: Tables of Integral Error Functions and Hermite Polynomials, by O. S. Berlyand, R. I. Gavrilova, A. P. Prudnikov. Biometrics, 19(3):500–501, September 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527940.

Proctor:1963:BRI

[305] J. R. Proctor. Book review: An Introduction to Statistical Science in Agriculture, by D. J. Finney. Biometrics, 19(3):501, September 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527941.

Healy:1963:BRA

[306] M. J. R. Healy. Book review: Annual Review in Automatic Programming, Vol. 3, by R. Goodman. Biometrics, 19(3):501–502, September 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527942.

Good:1963:BRD

[307] I. J. Good. Book review: Decisions, Values and Groups, Vol. 2, by Norman F. Washburne. Biometrics, 19(3):502, September 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527943.

Anonymous:1963:Ac

[308] Anonymous. Abstracts. Biometrics, 19(3):503–509, September 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527944.

Anonymous:1963:CCR

[309] Anonymous. Corrections: On the combination of relative risks. Biometrics, 19(3):509, September 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2527945. See [237].

Anonymous:1963:BSc

[310] Anonymous. The Biometric Society. Biometrics, 19(3):510–511, September 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2527946.

Anonymous:1963:NAc

[311] Anonymous. News and announcements. Biometrics, 19(3):512–513, September 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527947.

Anonymous:1963:FMc

[312] Anonymous. Front matter. Biometrics, 19(3):??, September 1963. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527927.

Anonymous:1963:VI

[313] Anonymous. Volume information. Biometrics, 19(4):i–xxii, December 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527528.

Duncan:1963:OLC

[314] D. B. Duncan. Obituary: S. Lee Crump. Biometrics, 19(4):515–516, December 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527530.

Plackett:1963:UTQ

[315] R. L. Plackett and P. S. Hewlett. A unified theory for quantal responses to mixtures of drugs: The fitting to data of certain models for two non-interactive drugs with complete positive correlation of tolerances. Biometrics, 19(4):517–531, December 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527531.

Sen:1963:ERP

[316] Pranab Kumar Sen. On the estimation of relative potency in dilution (-direct) assays by distribution-free methods. Biometrics, 19(4):532–552, December 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527532.

Federer:1963:PDU

[317] Walter T. Federer. Procedures and designs useful for screening material in selection and allocation, with a bibliography. Biometrics, 19(4):553– 587, December 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527533.

Doss:1963:EBA

[318] S. A. D. C. Doss. On the efficiency of best asymptotically normal estimates of the Poisson parameter based on singly and doubly truncated or censored samples. Biometrics, 19(4):588–594, December 1963. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527534.

Weiss:1963:CDS

[319] George H. Weiss. Comparison of a deterministic and a stochastic model for interaction between antagonistic species. Biometrics, 19(4):595–602, December 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527535.

Pielou:1963:RHD

[320] E. C. Pielou. Runs of healthy and diseased trees in transects through an infected forest. Biometrics, 19(4):603–614, December 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527536. See corrections [415].

Elandt:1963:OSA

[321] Regina C. Elandt. Optimal and sufficient allocation of multiple varietal experiments. Biometrics, 19(4):615–628, December 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527537.

Federer:1963:RBT

[322] W. T. Federer. Relationships between a three-way classification disproportionate numbers analysis of variance and several two-way classification and nested analyses. Biometrics, 19(4):629–637, December 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527538.

Myers:1963:IPP

[323] Robert J. Myers. An instance of the pitfalls prevalent in graveyard research. Biometrics, 19(4):638–650, December 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527539.

Birch:1963:NAL

[324] M. W. Birch. 194 note: an algorithm for the logarithmic series distribution. Biometrics, 19(4):651–652, December 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527540.

OCarroll:1963:NMG

[325] F. M. O’Carroll. 195 note: a method of generating randomised Latin squares. Biometrics, 19(4):652–653, December 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527541.

Anonymous:1963:Ad

[326] Anonymous. Abstracts. Biometrics, 19(4):654–666, December 1963. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527542.

Anonymous:1963:CMT

[327] Anonymous. Correction: Matching trials by double criteria. Biometrics, 19(4):667, December 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2527543. See [293].

Anonymous:1963:CSM

[328] Anonymous. Correction: Some multi-dimensional birth and death processes and their applications in population genetics. Biometrics, 19(4): 667, December 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527544. See

[233].

Campbell:1963:BS

[329] R. C. Campbell, M. J. R. Healy, and C. I. Bliss. The Biometric Society. Biometrics, 19(4):668–674, December 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527545.

Anonymous:1963:NAd

[330] Anonymous. News and announcements. Biometrics, 19(4):675–679, December 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527546.

Anonymous:1963:FMd

[331] Anonymous. Front matter. Biometrics, 19(4):??, December 1963. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527529.

Finney:1964:RVT

[332] D. J. Finney. The replication of variety trials. Biometrics, 20(1):1–15, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527614.

Behrens:1964:CMI

[333] W.-U. Behrens. The comparison of means of independent normal distributions with different variances. Biometrics, 20(1):16–27, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527615.

Perrin:1964:HRS

[334] Edward B. Perrin and Mindel C. Sheps. Human reproduction: a stochastic process. Biometrics, 20(1):28–45, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527616.

Harris:1964:EPP

[335] Dewey L. Harris. Expected and predicted progress from index selection involving estimates of population parameters. Biometrics, 20(1):46–72, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527617.

Roseberry:1964:OCC

[336] Thomas D. Roseberry and Edmund A. Gehan. Operating characteristic curves and accept–reject rules for two and three stage screening procedures. Biometrics, 20(1):73–84, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527618.

Gale:1964:SAT

[337] J. S. Gale. Some applications of the theory of junctions. Biometrics, 20(1):85–117, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527619.

Tallis:1964:SEA

[338] G. M. Tallis. Sampling errors associated with family selection. Biometrics, 20(1):118–121, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527620.

Bogyo:1964:CVH

[339] Thomas P. Bogyo. Coefficients of variation of heritability estimates obtained from variance analyses. Biometrics, 20(1):122–129, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527621. See correction [492].

MacQueen:1964:MAP

[340] J. MacQueen. A mathematical approach to the problem of achieving selective biological effects. Biometrics, 20(1):130–142, March 1964. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527622.

Read:1964:QAC

[341] D. R. Read. A quantitative approach to the comparative assessment of taste quality in the confectionery industry. Biometrics, 20(1):143–155, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527623.

Pielou:1964:SPT

[342] E. C. Pielou. The spatial pattern of two-phase patchworks of vegetation. Biometrics, 20(1):156–167, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527624.

Federer:1964:TDC

[343] W. T. Federer and G. F. Atkinson. Tied-double-change-over designs. Biometrics, 20(1):168–181, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527625.

Bennett:1964:NPT

[344] B. M. Bennett. A non-parametric test for randomness in a sequence of multinomial trials. Biometrics, 20(1):182–190, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527626.

Cochran:1964:ASL

[345] William G. Cochran. Approximate significance levels of the Behrens– Fisher test. Biometrics, 20(1):191–195, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527627.

Sprent:1964:NAV

[346] P. Sprent. 196. Note: Analysis of variance with small experiments. Biometrics, 20(1):196–199, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527628.

Freeman:1964:NUS

[347] G. H. Freeman. 197. Note: The use of a systematic design for a spacing trial with a tropical tree crop. Biometrics, 20(1):200–203, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527629.

Ramachandran:1964:NAC

[348] G. Ramachandran. 198. Note: Analysis of covariance for the split-plot design: Average variances. Biometrics, 20(1):204–206, March 1964. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527630.

Chassan:1964:NAS

[349] J. B. Chassan. 199. Note: On the analysis of simple cross-overs with unequal numbers of replicates. Biometrics, 20(1):206–208, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527631.

Dar:1964:NCS

[350] S. N. Dar. 200. Note: Comparison of the sensitivities of dependent experiments. Biometrics, 20(1):209–212, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527632.

Adke:1964:NMD

[351] S. R. Adke. 201. Note: a multi-dimensional birth and death process. Biometrics, 20(1):212–216, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527633.

Gale:1964:BRI

[352] J. S. Gale. Book review: Introduction to the Mathematical Theory of Genetic Linkage by N. T. J. Bailey. Biometrics, 20(1):217, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527634.

Spicer:1964:BRS

[353] C. C. Spicer. Book review: Statistics for Medical Students by R. G. Hoffman. Biometrics, 20(1):217–218, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527635.

F:1964:BRB

[354] W. T. F. Book review: Bibliography of Statistical Literature, 1950– 1958 by M. G. Kendall, A. G. Doig. Biometrics, 20(1):218, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527636.

Jeffers:1964:BRE

[355] J. N. R. Jeffers. Book review: Electronic Computers by A. I. Kitov, N. A. Krinitskii. Biometrics, 20(1):218–219, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527637.

Holgate:1964:BRE

[356] P. Holgate. Book review: Estimation from Grouped and Partially Grouped Samples by G. Kulldorf. Biometrics, 20(1):219–220, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527638.

F:1964:BRI

[357] W. T. F. Book review: Introduction to Econometrics by Oskar Lange. Biometrics, 20(1):220, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527639.

F:1964:BRS

[358] W. T. F. Book review: Statistics for the Teacher by Douglas M. McIntosh. Biometrics, 20(1):220, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527640.

Blackith:1964:BRE

[359] R. E. Blackith. Book review: The Essence of Biometry by J. Stanley. Biometrics, 20(1):220–221, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527641.

Gilbert:1964:BRM

[360] N. Gilbert. Book review: Matematicheskaya Statistika dlya Biologov i Medikov by V. Yu Urbakh. Biometrics, 20(1):221, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527642.

Anonymous:1964:Aa

[361] Anonymous. Abstracts. Biometrics, 20(1):222–225, March 1964. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527643.

Anonymous:1964:BSa

[362] Anonymous. The Biometric Society. Biometrics, 20(1):226–232, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527644.

Anonymous:1964:NAa

[363] Anonymous. News and announcements. Biometrics, 20(1):233–235, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527645.

Anonymous:1964:FMa

[364] Anonymous. Front matter. Biometrics, 20(1):??, March 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527613.

Anonymous:1964:MRA

[365] Anonymous. In memoriam: Ronald Aylmer Fisher, 1890–1962. Biometrics, 20(2):i–iii, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2528390. Special issue: In memoriam: Ronald Aylmer Fisher, 1890– 1962.

Finney:1964:F

[366] D. J. Finney. Foreword. Biometrics, 20(2):237, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:/ /www.jstor.org/stable/2528391. Special issue: In memoriam: Ronald Aylmer Fisher, 1890–1962.

Mahalanobis:1964:PRA

[367] P. C. Mahalanobis. Professor Ronald Aylmer Fisher. Biometrics, 20(2): 238–252, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528392. Spe-

cial issue: In memoriam: Ronald Aylmer Fisher, 1890–1962.

Fisher:1964:QTG

[368] R. A. Fisher. A quantitative theory of genetic recombination and chiasma formation. Biometrics, 20(2):253–260, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www. jstor.org/stable/2528393. Special issue: In memoriam: Ronald Aylmer Fisher, 1890–1962.

Fisher:1964:B

[369] R. A. Fisher. Biometry. Biometrics, 20(2):261–264, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:/ /www.jstor.org/stable/2528394. Special issue: In memoriam: Ronald Aylmer Fisher, 1890–1962.

Fisher:1964:SDE

[370] R. A. Fisher. The significance of deviations from expectation in a Poisson series. Biometrics, 20(2):265–272, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2528395. Special issue: In memoriam: Ronald Aylmer Fisher, 1890–1962.

Bliss:1964:RFC

[371] C. I. Bliss. R. A. Fisher’s contributions to medicine and bioassay. Biometrics, 20(2):273–285, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2528396. Special issue: In memoriam: Ronald Aylmer Fisher, 1890– 1962. See corrections [414].

Rao:1964:SRA

[372] C. Radhakrishna Rao. Sir Ronald Aylmer Fisher — the architect of multivariate analysis. Biometrics, 20(2):286–300, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:/ /www.jstor.org/stable/2528397. Special issue: In memoriam: Ronald Aylmer Fisher, 1890–1962.

Williams:1964:SEB

[373] C. B. Williams. Some experiences of a biologist with R. A. Fisher and statistics. Biometrics, 20(2):301–306, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www. jstor.org/stable/2528398. Special issue: In memoriam: Ronald Aylmer Fisher, 1890–1962.

Yates:1964:SRF

[374] F. Yates. Sir Ronald Fisher and the design of experiments. Biometrics, 20(2):307–321, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2528399. Special issue: In memoriam: Ronald Aylmer Fisher, 1890– 1962.

Finney:1964:SRF

[375] D. J. Finney. Sir Ronald Fisher’s contributions to biometric statistics. Biometrics, 20(2):322–329, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/ stable/2528400. Special issue: In memoriam: Ronald Aylmer Fisher, 1890–1962.

Mather:1964:RFW

[376] K. Mather. R. A. Fisher’s work in genetics. Biometrics, 20(2):330– 342, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528401. Special

issue: In memoriam: Ronald Aylmer Fisher, 1890–1962.

Yates:1964:FPR

[377] F. Yates. Fiducial probability, recognisable sub-sets and Behrens’ test. Biometrics, 20(2):343–360, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/ stable/2528402. Special issue: In memoriam: Ronald Aylmer Fisher, 1890–1962.

Race:1964:SNF

[378] R. R. Race. Some notes on Fisher’s contributions to human blood groups. Biometrics, 20(2):361–367, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/ stable/2528403. Special issue: In memoriam: Ronald Aylmer Fisher, 1890–1962.

Mahalanobis:1964:SPM

[379] P. C. Mahalanobis. Some personal memories of R. A. Fisher. Biometrics, 20(2):368–371, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2528404. Special issue: In memoriam: Ronald Aylmer Fisher, 1890– 1962.

Cornish:1964:FAA

[380] E. A. Cornish. Fisher’s activities in Australia, 1959–62. Biometrics, 20(2):372–373, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2528405. Special issue: In memoriam: Ronald Aylmer Fisher, 1890– 1962.

Mountford:1964:BRT

[381] M. D. Mountford. Book review: Tables for Testing Significance in a 2 × 2 Contingency Table, by D. J. Finney, R. Latscha, B. M. Bennett, P. Hsu. Biometrics, 20(2):374, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/ stable/2528406. Special issue: In memoriam: Ronald Aylmer Fisher, 1890–1962.

Searle:1964:BRC

[382] S. R. Searle. Book review: Computer Applications in Medicine, by E. E. Mason, W. G. Bulgren. Biometrics, 20(2):374–375, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:/ /www.jstor.org/stable/2528407. Special issue: In memoriam: Ronald Aylmer Fisher, 1890–1962.

Regier:1964:BRS

[383] H. A. Regier. Book review: Symposium on Marine Microbiology, by C. H. Oppenheimer. Biometrics, 20(2):375, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www. jstor.org/stable/2528408. Special issue: In memoriam: Ronald Aylmer Fisher, 1890–1962.

Healy:1964:BRM

[384] M. J. R. Healy. Book review: Méthodes Statistiques dans les Sciences Humaines, by P. Pèpe, M. Tisserand-Perrier. Biometrics, 20(2):375– 376, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528409. Special

issue: In memoriam: Ronald Aylmer Fisher, 1890–1962.

Nelder:1964:BRE

[385] J. A. Nelder. Book review: Les Essais thérapeutiques cliniques, by D. Schwartz, R. Flamant, J. Lellouch, C. Rouquette. Biometrics, 20(2): 376, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528410. Special

issue: In memoriam: Ronald Aylmer Fisher, 1890–1962.

Anonymous:1964:Ab

[386] Anonymous. Abstracts. Biometrics, 20(2):377–402, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:/ /www.jstor.org/stable/2528411. Special issue: In memoriam: Ronald Aylmer Fisher, 1890–1962.

Anonymous:1964:BSb

[387] Anonymous. The Biometric Society. Biometrics, 20(2):403–421, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528412. Special issue:

In memoriam: Ronald Aylmer Fisher, 1890–1962.

Anonymous:1964:NAb

[388] Anonymous. News and announcements. Biometrics, 20(2):422–425, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528413. Special issue:

In memoriam: Ronald Aylmer Fisher, 1890–1962.

Anonymous:1964:BM

[389] Anonymous. Back matter. Biometrics, 20(2):426–xii, June 1964. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528414. Special issue: In memo-

riam: Ronald Aylmer Fisher, 1890–1962.

Anonymous:1964:FMb

[390] Anonymous. Front matter. Biometrics, 20(2):??, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:/ /www.jstor.org/stable/2528389. Special issue: In memoriam: Ronald Aylmer Fisher, 1890–1962.

Bancroft:1964:AII

[391] T. A. Bancroft. Analysis and inference for incompletely specified models involving the use of preliminary test(s) of significance. Biometrics, 20(3):427–442, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528486.

Draper:1964:EMV

[392] Norman R. Draper and David M. Stoneman. Estimating missing values in unreplicated two-level factorial and fractional factorial designs. Biometrics, 20(3):443–458, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528487.

Gabriel:1964:PTH

[393] K. R. Gabriel. A procedure for testing the homogeneity of all sets of means in analysis of variance. Biometrics, 20(3):459–477, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528488.

Atiqullah:1964:RNN

[394] M. Atiqullah. Robustness to non-normality of tests for sensitivity in similar experiments. Biometrics, 20(3):478–481, September 1964. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528489.

Dunnett:1964:NTM

[395] C. W. Dunnett. New tables for multiple comparisons with a control. Biometrics, 20(3):482–491, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528490.

Sen:1964:ELC

[396] A. R. Sen and R. P. Chakrabarty. Estimation of loss of crop from pests and diseases of tea from sample surveys. Biometrics, 20(3):492–504, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528491.

Bohidar:1964:DEV

[397] Neeti R. Bohidar. Derivation and estimation of variance and covariance components associated with covariance between relatives under sexlinked transmission. Biometrics, 20(3):505–521, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528492.

Eberhart:1964:TRA

[398] S. A. Eberhart. Theoretical relations among single, three-way, and double cross hybrids. Biometrics, 20(3):522–539, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528493.

Hanson:1964:GEI

[399] W. D. Hanson. Genotype-environment interaction concepts for field experimentation. Biometrics, 20(3):540–552, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528494.

Zinger:1964:SSF

[400] Alexis Zinger. Systematic sampling in forestry. Biometrics, 20(3):553– 565, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528495.

Hewlett:1964:UTQ

[401] P. S. Hewlett and R. L. Plackett. A unified theory for quantal responses to mixtures of drugs: Competitive action. Biometrics, 20(3):566–575, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528496. See

corrections [416].

Davies:1964:SIM

[402] O. L. Davies. Screening for improved mutants in antibiotic research. Biometrics, 20(3):576–591, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528497.

Snell:1964:SPO

[403] E. J. Snell. A scaling procedure for ordered categorical data. Biometrics, 20(3):592–607, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528498.

Bradley:1964:MTT

[404] Ralph A. Bradley and Tom J. Harmon. The modified triangle test. Biometrics, 20(3):608–625, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528499.

Ederer:1964:SPS

[405] Fred Ederer, Max H. Myers, and Nathan Mantel. A statistical problem in space and time: Do leukemia cases come in clusters? Biometrics, 20(3):626–638, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528500.

Bailar:1964:NSF

[406] John C. Bailar, III and Fred Ederer. 202. Note: Significance factors for the ratio of a Poisson variable to its expectation. Biometrics, 20(3):639– 643, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528501.

Elston:1964:NET

[407] R. C. Elston. 203. Note: On estimating time-response curves. Biometrics, 20(3):643–647, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528502.

Holgate:1964:NEN

[408] P. Holgate. 204. Note: The efficiency of nearest neighbour estimators. Biometrics, 20(3):647–649, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528503.

Adke:1964:NMP

[409] S. R. Adke. 205. Note: The maximum population size in the first N generations of a branching process. Biometrics, 20(3):649–651, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528504.

Steel:1964:BRA

[410] R. G. D. Steel. Book review: Applied Multiple Linear Regression, by R. A. Bottenberg, J. H. Ward Jr. Biometrics, 20(3):652–653, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528505.

Choi:1964:BRR

[411] Keewhan Choi. Book review: Review of Basic Statistics, by Thomas E. Kutz. Biometrics, 20(3):653–654, September 1964. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528506.

F:1964:BRO

[412] W. T. F. Book review: An Outline of Norwegian Forestry, by Knut Skinnemoen. Biometrics, 20(3):654, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528507.

Anonymous:1964:Ac

[413] Anonymous. Abstracts. Biometrics, 20(3):655–665, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528508.

Anonymous:1964:CRF

[414] Anonymous. Corrections: R. A. Fisher’s contributions to medicine and bioassay. Biometrics, 20(3):666, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528511. See [371].

Anonymous:1964:CRH

[415] Anonymous. Corrections: Runs of healthy and diseased trees in transects through an infected forest. Biometrics, 20(3):666, September 1964. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528509. See [320].

Anonymous:1964:CUT

[416] Anonymous. Corrections: A unified theory for quantal responses to mixtures of drugs: The fitting to date of certain models for two noninteractive drugs with complete positive correlation of tolerances. Biometrics, 20(3):666, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2528510. See [401].

Anonymous:1964:BSc

[417] Anonymous. The Biometric Society. Biometrics, 20(3):667–677, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528512.

Anonymous:1964:NAc

[418] Anonymous. News and announcements. Biometrics, 20(3):678–680, September 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528513.

Anonymous:1964:FMc

[419] Anonymous. Front matter. Biometrics, 20(3):??, September 1964. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528485.

Anonymous:1964:VI

[420] Anonymous. Volume information. Biometrics, 20(4):i–x, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528120.

Elston:1964:HCT

[421] R. C. Elston and N. Bush. The hypotheses that can be tested when there are interactions in an analysis of variance model. Biometrics, 20(4):681– 698, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528122.

Pearce:1964:EBN

[422] S. C. Pearce. Experimenting with blocks of natural size. Biometrics, 20(4):699–706, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528123.

Berenblut:1964:CDC

[423] I. I. Berenblut. Change-over designs with complete balance for first residual effects. Biometrics, 20(4):707–712, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528124.

Freeman:1964:AFT

[424] G. H. Freeman. The addition of further treatments to Latin square designs. Biometrics, 20(4):713–729, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528125.

Nissen-Meyer:1964:EST

[425] Sven Nissen-Meyer. Evaluation of screening tests in medical diagnosis. Biometrics, 20(4):730–755, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528126.

Tootill:1964:OSD

[426] J. P. R. Tootill. Optimal symmetrical double sampling schemes for testing the antigenicity of vaccines. Biometrics, 20(4):756–769, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528127.

Sen:1964:TVF

[427] Pranab Kumar Sen. Tests for the validity of the fundamental assumption in dilution (-direct) assays. Biometrics, 20(4):770–784, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528128.

Cormack:1964:BPA

[428] R. M. Cormack. A boundary problem arising in population genetics. Biometrics, 20(4):785–793, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528129.

Evans:1964:DMR

[429] D. A. Evans and U. Philip. On the distribution of Mendelian ratios. Biometrics, 20(4):794–817, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528130.

Swiger:1964:VIC

[430] L. A. Swiger, W. R. Harvey, D. O. Everson, and K. E. Gregory. The variance of intraclass correlation involving groups with one observation. Biometrics, 20(4):818–826, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528131.

Turner:1964:SLN

[431] Malcolm E. Turner. The spectrum of a linear network: a simple approach to the n compartment problem. Biometrics, 20(4):827–831, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528132.

Tallis:1964:UMA

[432] G. M. Tallis. The use of models in the analysis of some classes of contingency tables. Biometrics, 20(4):832–839, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528133.

Worcester:1964:MSE

[433] Jane Worcester. Matched samples in epidemiologic studies. Biometrics, 20(4):840–848, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528134.

Jowett:1964:HSC

[434] G. H. Jowett. The harmonic standardisation of comparisons between success rates in two heterogeneous groups of patients. Biometrics, 20(4):849– 858, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528135.

Hoel:1964:MCG

[435] Paul G. Hoel. Methods for comparing growth type curves. Biometrics, 20(4):859–872, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528136.

Inkson:1964:PES

[436] R. H. E. Inkson. The precision of estimates of the soil content of phosphate using the Mitscherlich response equation. Biometrics, 20(4):873– 882, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528137.

Nelder:1964:EN

[437] J. A. Nelder. Editorial note. Biometrics, 20(4):883, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528138.

Bliss:1964:NAA

[438] C. I. Bliss and David A. Fleischer. 206. Note: The analysis in angles of a microbial experiment. Biometrics, 20(4):883–891, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528139.

Bartlett:1964:NNS

[439] M. S. Bartlett. 207. Note: a note on spatial pattern. Biometrics, 20(4):891–892, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528140.

Wilcoxon:1964:NNP

[440] Frank Wilcoxon and Ralph A. Bradley. 208. Note: a note on the paper two sequential two-sample grouped rank tests with application to screening experiments. Biometrics, 20(4):892–895, December 1964. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528141.

Anonymous:1964:OWC

[441] Anonymous. Obituary: Walter C. Jacob. Biometrics, 20(4):896, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528142.

Walpole:1964:BRB

[442] R. E. Walpole. Book review: Basic Concepts of Probability and Statistics, by J. L. Hodges, Jr., E. L. Lehman. Biometrics, 20(4):897, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528143.

Harshbarger:1964:BRS

[443] Boyd Harshbarger. Book review: Statistical Inference I, by C. R. Li. Jerome. Biometrics, 20(4):898, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528144.

Bozovich:1964:BRS

[444] Helen Bozovich. Book review: Statistical Techniques Based on Probabilistic Models, by G. A. Baker. Biometrics, 20(4):898–899, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528145.

Martin:1964:BRI

[445] M. P. Martin. Book review: An Introduction to the Experimental Method for Students of Biology and the Health Sciences, by J. M. Little. Biometrics, 20(4):899, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528146.

F:1964:BRP

[446] W. T. F. Book review: Physiological Problems in Space Exploration, by J. D. Hardy, Charles C. Thomas. Biometrics, 20(4):899–900, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528147.

Anonymous:1964:BR

[447] Anonymous. [Books received]. Biometrics, 20(4):900, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528149.

Bliss:1964:BRR

[448] C. I. Bliss. Book review: Résumés des Travaux Publiés par l’Academie des Sciences de Bulgarie en 1944–50, 51–52, 53–55, 56–57. Biometrics, 20(4):900, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528148.

Anonymous:1964:Ad

[449] Anonymous. Abstracts. Biometrics, 20(4):901–915, December 1964. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528150.

Anonymous:1964:BSd

[450] Anonymous. The Biometric Society. Biometrics, 20(4):916–927, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528151.

Anonymous:1964:NAd

[451] Anonymous. News and announcements. Biometrics, 20(4):928–932, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528152.

Anonymous:1964:FMd

[452] Anonymous. Front matter. Biometrics, 20(4):??, December 1964. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528121.

Anonymous:1965:FMa

[453] Anonymous. Front matter. Biometrics, 21(1):i–ii, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528347.

Lefkovitch:1965:SPG

[454] L. P. Lefkovitch. The study of population growth in organisms grouped by stages. Biometrics, 21(1):1–18, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528348.

Lewontin:1965:RHT

[455] R. C. Lewontin and J. Felsenstein. The robustness of homogeneity tests in 2 × N tables. Biometrics, 21(1):19–33, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528349.

Martin:1965:FCC

[456] D. C. Martin and S. K. Katti. Fitting of certain contagious distributions to some available data by the maximum likelihood method. Biometrics, 21(1):34–48, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2528350. See correction [521].

Katti:1965:BES

[457] S. K. Katti and A. N. Sastry. Biological examples of small expected frequencies and the chi-square test. Biometrics, 21(1):49–54, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528351.

Holland:1965:SEO

[458] D. A. Holland. Sampling errors in an orchard survey involving unequal numbers of orchards of distinct type. Biometrics, 21(1):55–62, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528352.

Giesbrecht:1965:ERM

[459] F. Giesbrecht and O. Kempthorne. Examination of a repeat mating design for estimating environmental and genetic trends. Biometrics, 21(1): 63–85, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528353.

Radhakrishna:1965:CRS

[460] S. Radhakrishna. Combination of results from several 2 × 2 contingency tables. Biometrics, 21(1):86–98, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528354.

Mertz:1965:MPE

[461] David B. Mertz, Thomas Park, and W. J. Youden. Mortality patterns in eight strains of flour beetles. Biometrics, 21(1):99–114, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528355.

Gosslee:1965:AVD

[462] David G. Gosslee and Henry L. Lucas. Analysis of variance of disproportionate data when interaction is present. Biometrics, 21(1):115–133, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528356.

Gani:1965:SPA

[463] J. Gani. Stochastic phage attachment to bacteria. Biometrics, 21(1):134– 139, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528357.

Elston:1965:SME

[464] R. C. Elston. A simple method of estimating relative potency from two parabolas. Biometrics, 21(1):140–149, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528358.

Brock:1965:SSA

[465] Norbert Brock and Berthold Schneider. A screening system for anticancer agents based on the therapeutic index. Biometrics, 21(1):150– 158, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528359.

Newell:1965:UFD

[466] D. J. Newell. Unusual frequency distributions. Biometrics, 21(1):159– 168, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528360.

Colton:1965:TSM

[467] Theodore Colton. A two-stage model for selecting one of two treatments. Biometrics, 21(1):169–180, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528361.

Ashford:1965:ASC

[468] J. R. Ashford and C. S. Smith. An alternative system for the classification of mathematical models for quantal responses to mixtures of drugs in biological assay. Biometrics, 21(1):181–188, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528362.

Pearce:1965:EPB

[469] S. C. Pearce. Expository papers for biometrics. Biometrics, 21(1):189, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528363.

Cattell:1965:BIPa

[470] Raymond B. Cattell. A biometrics invited paper. factor analysis: an introduction to essentials i. the purpose and underlying models. Biometrics, 21(1):190–215, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528364.

Cattell:1965:FAIa

[471] Raymond B. Cattell. Factor analysis: an introduction to essentials. I. The purpose and underlying models. Biometrics, 21(1):190–215, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Robinson:1965:NAD

[472] P. Robinson. 209. Note: The analysis of a diallel crossing experiment with certain crosses missing. Biometrics, 21(1):216–219, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528365.

Cicchinelli:1965:NTP

[473] A. L. Cicchinelli. 210. Note: Tables of Pearson–Lee–Fisher functions of singly truncated normal distributions. Biometrics, 21(1):219–226, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528366.

Emerson:1965:NOP

[474] Phillip L. Emerson. 211. Note: Orthogonal polynomials for unequally weighted means. Biometrics, 21(1):226–230, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528367.

Bennett:1965:NCL

[475] B. M. Bennett. 212. Note: Confidence limits for a ratio using Wilcoxon’s signed rank test. Biometrics, 21(1):231–234, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528368.

Oldham:1965:NEA

[476] P. D. Oldham. 213. Note: On estimating the arithmetic means of lognormally-distributed populations. Biometrics, 21(1):235–239, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528369.

F:1965:BRBa

[477] W. T. F. Book review: Bibliography on Sampling of Raw Materials and Products in Bulk, by C. A. Bicking. Biometrics, 21(1):240, March 1965.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528371.

F:1965:BRIa

[478] W. T. F. Book review: Instrumentation with Semiconductors, by C. C. Brown. Biometrics, 21(1):240, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528372.

F:1965:BRMa

[479] W. T. F. Book review: Manual of Procedures in Quantitative Genetics, by W. A. Becker. Biometrics, 21(1):240, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528370.

Ledermann:1965:BRC

[480] Sully Ledermann. Book review: Cours de Statistique Descriptive, by G. Calot. Biometrics, 21(1):241, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528373.

Williams:1965:BRS

[481] E. J. Williams. Book review: Statistical Estimation in Factor Analysis. A New Technique and Its Foundation, by K. G. Jöreskog. Biometrics, 21(1): 241–242, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528374.

Balaam:1965:BRI

[482] L. N. Balaam. Book review: Introduction to Experimental Statistics, by C. C. Li. Biometrics, 21(1):242–243, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528375.

Harshbarger:1965:BRS

[483] Boyd Harshbarger. Book review: Statistical Inference II, by J. C. R. Li. Biometrics, 21(1):243, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528376.

Link:1965:BRS

[484] R. F. Link. Book review: Statistical Theory, by B. W. Lindgren. Biometrics, 21(1):244, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528377.

OMathuna:1965:BRN

[485] Diarmuid O’Mathuna. Book review: Nomography, by Edward Otto. Biometrics, 21(1):244–245, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528378.

Balaam:1965:BRP

[486] L. N. Balaam. Book review: The Principles and Practice of Agricultural Research, by S. G. Salmon, A. A. Hanson. Biometrics, 21(1):246, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528379.

Searle:1965:BRE

[487] S. R. Searle. Book review: Electronic and Computer-Assisted Studies of Bio-Medical Problems, by O. E. Schmitt, C. A. Cacenes. Biometrics, 21(1):246–247, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528380.

F:1965:BRPa

[488] W. T. F. Book review: Proceedings of the 34th Session of the International Statistical Institute. Biometrics, 21(1):247, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528381.

F:1965:BRPb

[489] W. T. F. Book review: Proceedings of the Ninth Conference on the Design of Experiments in Army Research Development and Testing. Biometrics, 21(1):247, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528382.

Anonymous:1965:BRa

[490] Anonymous. [Books received]. Biometrics, 21(1):248, March 1965. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528383.

Anonymous:1965:Aa

[491] Anonymous. Abstracts. Biometrics, 21(1):249–264, March 1965. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528384.

Anonymous:1965:CCV

[492] Anonymous. Correction: Coefficients of variation of heritability estimates obtained from variance analyses. Biometrics, 21(1):265, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528385. See [339].

Tucker:1965:BS

[493] Henry Tucker, C. I. Bliss, and Erwin L. LeClerg. The Biometric Society. Biometrics, 21(1):266–285, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528386.

Anonymous:1965:NAa

[494] Anonymous. News and announcements. Biometrics, 21(1):286–290, March 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528387.

Anonymous:1965:FMb

[495] Anonymous. Front matter. Biometrics, 21(2):i–ii, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528089.

Binet:1965:CII

[496] F. E. Binet. On the construction of an index for indirect selection. Biometrics, 21(2):291–299, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528090.

Glasser:1965:RAD

[497] Marvin Glasser. Regression analysis with dependent variable censored. Biometrics, 21(2):300–307, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528091.

Mielke:1965:NOT

[498] Paul W. Mielke, Jr. and Richard B. McHugh. Non-orthogonality in the two-way classification for the mixed effects finite population model. Biometrics, 21(2):308–323, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528092.

Westlake:1965:CDB

[499] W. J. Westlake. Composite designs based on irregular fractions of factorials. Biometrics, 21(2):324–336, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2528093. See correction [581].

Peritz:1965:IOR

[500] Eric Peritz. On inferring order relations in analysis of variance. Biometrics, 21(2):337–344, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528094.

Dagnelie:1965:LCVa

[501] Pierre Dagnelie. L’Étude des communautés végétales par L’Analyse statistique des liaisons entre les espèces et les variables Écologiques: Principes fondamentaux. (French) [The study of plant communities by statistical analysis of links between species and ecological variables: Fundamental principles]. Biometrics, 21(2):345–361, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528095.

Edwards:1965:MCA

[502] A. W. F. Edwards and L. L. Cavalli-Sforza. A method for cluster analysis. Biometrics, 21(2):362–375, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528096.

Mathai:1965:AMA

[503] A. M. Mathai. An approximate method of analysis for a two-way layout. Biometrics, 21(2):376–385, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528097.

Ewens:1965:ADA

[504] W. J. Ewens. The adequacy of the diffusion approximation to certain distributions in genetics. Biometrics, 21(2):386–394, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528098.

Compton:1965:FEC

[505] W. A. Compton, C. O. Gardner, and J. H. Lonnquist. Further evidence on the consistency of estimates of variance components. Biometrics, 21(2):395–404, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528099.

Cattell:1965:BIPb

[506] Raymond B. Cattell. A biometrics invited paper. factor analysis: an introduction to essentials II. the role of factor analysis in research. Biometrics, 21(2):405–435, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528100.

Cattell:1965:FAIb

[507] Raymond B. Cattell. Factor analysis: an introduction to essentials. II. The role of factor analysis in research. Biometrics, 21(2):405–435, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bohren:1965:EMS

[508] B. B. Bohren, H. E. McKean, and G. W. Friars. The expected mean squares in genetic experiments when only one parent is identified. Biometrics, 21(2):436–446, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528101.

Robinson:1965:EEI

[509] P. Robinson and D. F. Bray. Expected effects on the inbreeding coefficient and rate of gene loss of four methods of reproducing finite diploid populations. Biometrics, 21(2):447–458, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528102.

Lentner:1965:LEM

[510] M. M. Lentner. Listing expected mean square components. Biometrics, 21(2):459–466, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528103.

Grizzle:1965:TPC

[511] James E. Grizzle. The two-period change-over design and its use in clinical trials. Biometrics, 21(2):467–480, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528104. See corrections [1060].

Weiss:1965:SEC

[512] George H. Weiss. On the spread of epidemics by carriers. Biometrics, 21(2):481–490, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528105.

Lu:1965:HAH

[513] K. H. Lu. Harmonic analysis of the human face. Biometrics, 21(2):491– 505, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528106.

Shah:1965:NMF

[514] B. K. Shah. 214. Note: a method of fitting a non-linear curve containing a single non-linearity. Biometrics, 21(2):506–507, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528107.

F:1965:BRS

[515] W. T. F. Book review: Sampling Techniques by W. G. Cochran. Biometrics, 21(2):508, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528108.

Wallace:1965:BRE

[516] T. D. Wallace. Book review: Econometric Theory by A. S. Goldberger. Biometrics, 21(2):509–510, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528109.

Choi:1965:BRR

[517] K. Choi. Book review: Russian–English Dictionary of Statistical Terms and Expressions and Russian Reader in Statistics by Samuel Kotz. Biometrics, 21(2):511, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528110.

Anderson:1965:BRE

[518] Paul S. Anderson, Jr. Book review: Elementary Medical Statistics by D. Mainland. Biometrics, 21(2):511–512, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528111.

F:1965:BRFa

[519] W. T. F. Book review: Fifty Challenging Problems in Probability with Solutions by F. Mosteller. Biometrics, 21(2):512, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528112.

F:1965:BRPc

[520] W. T. F. Book review: Proceedings of the Seminar Sponsored by the United Nations and the World Health Organization. Biometrics, 21(2): 512–514, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528113.

Anonymous:1965:CFS

[521] Anonymous. Correction: Fitting of some contagious distributions to some available data by the maximum likelihood method. Biometrics, 21 (2):514, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528116. See

[456].

F:1965:BRFb

[522] W. T. F. Book review: Food Supplies and Population Growth by Royal Statistical Society. Biometrics, 21(2):514, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528115.

F:1965:BRMb

[523] W. T. F. Book review: Manual of Mathematics for Students of Agriculture by F. Robertson. Biometrics, 21(2):514, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528114.

LeRoy:1965:BS

[524] H. L. LeRoy. The Biometric Society. Biometrics, 21(2):515–525, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528117.

Anonymous:1965:NAb

[525] Anonymous. News and announcements. Biometrics, 21(2):526–528, June 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528118.

Anonymous:1965:FMc

[526] Anonymous. Front matter. Biometrics, 21(3):i–ii, September 1965. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528538.

Chapman:1965:EMR

[527] Douglas G. Chapman. The estimation of mortality and recruitment from a single-tagging experiment. Biometrics, 21(3):529–542, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528539. See correction [795].

Watson:1965:DO

[528] Geoffrey S. Watson. The distribution of organisms. Biometrics, 21 (3):543–550, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528540.

Horner:1965:MRB

[529] Theodore W. Horner. Mathematical representation of the biological and physical decay of chamber aerosols. Biometrics, 21(3):551–561, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528541.

Shortley:1965:SMD

[530] George Shortley. A stochastic model for distributions of biological response times. Biometrics, 21(3):562–582, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528542.

Gart:1965:SSM

[531] John J. Gart. Some stochastic models relating time and dosage in response curves. Biometrics, 21(3):583–599, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528543.

Roberts:1965:ECV

[532] E. A. Roberts and G. G. Coote. The estimation of concentration of viruses and bacteria from dilution counts. Biometrics, 21(3):600–615, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528544.

Tootill:1965:CRC

[533] J. P. R. Tootill. On combining the results from clinical trials of a vaccine. Biometrics, 21(3):616–622, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528545.

Billewicz:1965:EMS

[534] W. Z. Billewicz. The efficiency of matched samples: an empirical investigation. Biometrics, 21(3):623–644, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528546.

David:1965:IAP

[535] H. A. David and D. J. Newell. The identification of annual peak periods for a disease. Biometrics, 21(3):645–650, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528547.

Ogilvie:1965:PCM

[536] John C. Ogilvie. Paired comparison models with tests for interaction. Biometrics, 21(3):651–664, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528548.

Shikata:1965:GIC

[537] Morikazu Shikata. A generalization of the inbreeding coefficient. Biometrics, 21(3):665–681, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528549.

Searle:1965:VIS

[538] S. R. Searle. The value of indirect selection. I. Mass selection. Biometrics, 21(3):682–707, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528550.

Grosenbaugh:1965:GRS

[539] L. R. Grosenbaugh. Generalization and reparameterization of some sigmoid and other nonlinear functions. Biometrics, 21(3):708–714, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528551.

Holgate:1965:FSL

[540] P. Holgate. Fitting a straight line to data from a truncated population. Biometrics, 21(3):715–720, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528552.

Mead:1965:GLN

[541] R. Mead. A generalised logit-normal distribution. Biometrics, 21(3):721– 732, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528553.

Linhart:1965:ACL

[542] H. Linhart. Approximate confidence limits for the coefficient of variation of gamma distributions. Biometrics, 21(3):733–738, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528554.

Sutherland:1965:CBF

[543] T. M. Sutherland. The correlation between feed efficiency and rate of gain, a ratio and its denominator. Biometrics, 21(3):739–749, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528555.

Nelson:1965:Q

[544] August E. Nelson and Cedric A. B. Smith. 215. Query. Biometrics, 21(3):750–752, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528556.

Srivastava:1965:NEA

[545] S. R. Srivastava and V. P. Gupta. 216. Note: Estimation after preliminary testing in anova model I. Biometrics, 21(3):752–758, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528557.

Anonymous:1965:BRb

[546] Anonymous. [Book received]. Biometrics, 21(3):759–760, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528558.

Anonymous:1965:Ab

[547] Anonymous. Abstracts. Biometrics, 21(3):761–777, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528559.

Anonymous:1965:BSa

[548] Anonymous. The Biometric Society. Biometrics, 21(3):778–779, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528560.

Anonymous:1965:NAc

[549] Anonymous. News and announcements. Biometrics, 21(3):780–784, September 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528561.

Anonymous:1965:FMd

[550] Anonymous. Front matter. Biometrics, 21(4):i–ii, December 1965. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528243.

Anonymous:1965:VI

[551] Anonymous. Volume information. Biometrics, 21(4):i–xxvii, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528242.

Finney:1965:MB

[552] D. J. Finney. The meaning of bioassay. Biometrics, 21(4):785–798, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528244. See com-

ments [702, 761].

Sen:1965:SFA

[553] Pranab Kumar Sen. Some further applications of non-parametric methods in dilution (-direct) assays. Biometrics, 21(4):799–810, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528245.

Ashford:1965:AQR

[554] J. R. Ashford and C. S. Smith. An analysis of quantal response data in which the measurement of response is subject to error. Biometrics, 21(4):811–825, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528246.

Feigl:1965:EES

[555] Polly Feigl and Marvin Zelen. Estimation of exponential survival probabilities with concomitant information. Biometrics, 21(4):826–838, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528247.

Kreyberg:1965:ERL

[556] H. J. A. Kreyberg. Empirical relationship of lung cancer incidence to cigarette smoking and a stochastic model for the mode of action of carcinogens. Biometrics, 21(4):839–857, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528248.

Cornell:1965:SES

[557] Richard G. Cornell. Spearman estimation for a simple exponential model. Biometrics, 21(4):858–864, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528249.

Redman:1965:GSU

[558] C. E. Redman and E. P. King. Group screening utilizing balanced and partially balanced incomplete block designs. Biometrics, 21(4):865–874, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528250.

Kimball:1965:MCM

[559] A. W. Kimball. A model for chemical mutagenesis in bacteriophage. Biometrics, 21(4):875–889, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528251.

Dagnelie:1965:LCVb

[560] Pierre Dagnelie. L’Étude des communautés végétales par l’Analyse statistique des liaisons entre les espèces et les variables Écologiques: Un exemple. (French) [The study of plant communities through statistical analysis of links between species and ecological variables: an example]. Biometrics, 21(4):890–907, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528252.

Pielou:1965:CRP

[561] E. C. Pielou. The concept of randomness in the patterns of mosaics. Biometrics, 21(4):908–920, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528253.

Chapman:1965:EMP

[562] Douglas G. Chapman and Garth I. Murphy. Estimates of mortality and population from survey-removal records. Biometrics, 21(4):921–935, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528254.

Robson:1965:NSM

[563] D. S. Robson and W. A. Flick. A nonparametric statistical method for culling recruits from a mark–recapture experiment. Biometrics, 21(4):936–947, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528255.

Sardana:1965:CAS

[564] M. G. Sardana and M. N. Das. On the construction and analysis of some confounded asymmetrical factorial designs. Biometrics, 21(4):948–956, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528256.

Katti:1965:MCA

[565] S. K. Katti. Multiple covariate analysis. Biometrics, 21(4):957–974, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528257. See cor-

rection [685].

McGilchrist:1965:ACE

[566] C. A. McGilchrist. Analysis of competition experiments. Biometrics, 21(4):975–985, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528258.

Trawinski:1965:EPD

[567] B. J. Trawinski. An exact probability distribution over sample spaces of paired comparisons. Biometrics, 21(4):986–1000, December 1965. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528259.

Bogyo:1965:EHT

[568] Thomas P. Bogyo and Walter A. Becker. Estimates of heritability from transformed percentage sib data with unequal subclass numbers. Biometrics, 21(4):1001–1007, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528260.

Fleiss:1965:NNC

[569] Joseph L. Fleiss. 215. Note: a note on Cochran’s Q test. Biometrics, 21(4):1008–1010, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528261.

Healy:1965:NCD

[570] M. J. R. Healy. 216. Note: Computing a discriminant function from within-sample dispersions. Biometrics, 21(4):1011–1012, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528262.

Foulkes:1965:NFA

[571] H. O. Foulkes. 217. Note: Further analysis of R. A. Fisher’s enumerations in genetics. Biometrics, 21(4):1012–1015, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528263.

F:1965:BRIb

[572] W. T. F. Book review: An Index of Health: Mathematical Models by C. L. Chiang. Biometrics, 21(4):1016, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528264.

F:1965:BRT

[573] W. T. F. Book review: Tables for Yule’s Q Association Coefficient for Pairs of Percentages by J. A. Davis, R. Gilman, Judith Schick. Biometrics, 21(4):1016, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528265.

Balaam:1965:BRB

[574] L. N. Balaam. Book review: Biostatistics — an Introductory Text by Avram Goldstein. Biometrics, 21(4):1016–1018, December 1965. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528266.

F:1965:BRBb

[575] W. T. F. Book review: Bibliography of Statistical Literature, 1940–1949 by M. G. Kendall, A. G. Diog. Biometrics, 21(4):1018, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528267.

Schaffer:1965:BRA

[576] Henry E. Schaffer. Book review: Alcool Alcoolisme Alcoolisation by Sully Lederman. Biometrics, 21(4):1018, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528268.

Raktoe:1965:BRA

[577] B. L. Raktoe. Book review: Application of Characteristic Functions by E. Lukacs, R. G. Laha. Biometrics, 21(4):1019, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528269.

Anonymous:1965:BRc

[578] Anonymous. [Books received]. Biometrics, 21(4):1019–1020, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528270.

Anonymous:1965:Ac

[579] Anonymous. Abstracts. Biometrics, 21(4):1021–1031, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528271.

Anonymous:1965:BSb

[580] Anonymous. The Biometric Society. Biometrics, 21(4):1032–1036, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528272.

Anonymous:1965:CCD

[581] Anonymous. Correction: Composite designs based on irregular fractions of factorials. Biometrics, 21(4):1036, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528273. See [499].

Anonymous:1965:NAd

[582] Anonymous. News and announcements. Biometrics, 21(4):1037–1041, December 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528274.

Anonymous:1966:FMa

[583] Anonymous. Front matter. Biometrics, 22(1):i–ii, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528209.

Preece:1966:SRC

[584] D. A. Preece. Some row and column designs for two sets of treatments. Biometrics, 22(1):1–25, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2528210. See corrections [1045].

Duncan:1966:MRC

[585] David B. Duncan and Mackenzie Walser. Multiple regression combining within- and between-plot information. Biometrics, 22(1):26–43, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528211.

Katti:1966:IAG

[586] S. K. Katti. Interrelations among generalized distributions and their components. Biometrics, 22(1):44–52, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528212.

Boen:1966:QDE

[587] J. R. Boen and David Sylwester. A quantitative discussion of the effectiveness of voiding as a defense against bladder infection. Biometrics, 22(1):53–57, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528213.

Campbell:1966:CAL

[588] R. C. Campbell. The chick assay of lysine. Biometrics, 22(1):58–73, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528214.

Eisenberg:1966:GUP

[589] Herbert B. Eisenberg, Randolph R. M. Geoghagen, and John E. Walsh. A general use of the Poisson approximation for binomial events, with application to bacterial endocarditis data. Biometrics, 22(1):74–82, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528215.

Mantel:1966:MCC

[590] Nathan Mantel. Models for complex contingency tables and polychotomous dosage response curves. Biometrics, 22(1):83–95, March 1966. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528216.

Burnaby:1966:GID

[591] T. P. Burnaby. Growth-invariant discriminant functions and generalized distances. Biometrics, 22(1):96–110, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528217.

Mayo:1966:PSI

[592] O. Mayo. On the problem of self-incompatibility alleles. Biometrics, 22(1):111–120, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528218.

Tallis:1966:EUS

[593] G. M. Tallis. Equilibria under selection for k alleles. Biometrics, 22(1): 121–127, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528219.

Nelder:1966:IPU

[594] J. A. Nelder. Inverse polynomials, a useful group of multi-factor response functions. Biometrics, 22(1):128–141, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528220.

Pike:1966:MAC

[595] M. C. Pike. A method of analysis of a certain class of experiments in carcinogenesis. Biometrics, 22(1):142–161, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528221.

Riffenburgh:1966:GPE

[596] Robert H. Riffenburgh. On growth parameter estimation for early life stages. Biometrics, 22(1):162–178, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528222.

Gebhardt:1966:NAC

[597] Friedrich Gebhardt. 218. Note: Approximations to the critical values for Duncan’s multiple range test. Biometrics, 22(1):179–182, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528223.

Mantel:1966:NCC

[598] Nathan Mantel. 219. Note: Corrected correlation coefficients when observation on one variable is restricted. Biometrics, 22(1):182–187, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528224.

Scheinberg:1966:NSV

[599] Eliyahu Scheinberg. 220. Note: The sampling variance of the correlation coefficients estimated in genetic experiments. Biometrics, 22(1): 187–191, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528225. See

corrections [1058].

Bradley:1966:OFW

[600] Ralph A. Bradley. Obituary: Frank Wilcoxon. Biometrics, 22(1):192– 194, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528226.

Anonymous:1966:AOH

[601] Anonymous. Appreciation: Ottokar heinisch. Biometrics, 22(1):195, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528227.

Balaam:1966:BRSa

[602] L. N. Balaam. Book review: Statistical Theory and Methodology in Science and Engineering by K. A. Brownlee. Biometrics, 22(1):196, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528228.

Sprent:1966:BRN

[603] P. Sprent. Book review: Numerical Analysis by I. M. Khabaza. Biometrics, 22(1):196–197, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528229.

Campbell:1966:BRI

[604] R. C. Campbell. Book review: Introduction to Probability and Statistics from a Bayesian Viewpoint by D. V. Lindley. Biometrics, 22(1):197– 199, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528230.

Skellam:1966:BRE

[605] J. G. Skellam. Book review: Embodiments of Mind by W. S. McCulloch. Biometrics, 22(1):199–200, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528231.

Kendall:1966:BRI

[606] M. G. Kendall. Book review: Inference and Disputed Authorship: The Federalist by F. Mosteller, D. L. Wallace. Biometrics, 22(1):200–202, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528232.

Clatworthy:1966:BRC

[607] W. H. Clatworthy. Book reviews: Contributions to Statistics by C. R. Rao; Essays on Econometrics and Planning, by C. R. Rao. Biometrics, 22(1):202–205, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528233.

Searle:1966:BRT

[608] S. R. Searle. Book review: Techniques and Procedures in Animal Production by Clair E. Terrill. Biometrics, 22(1):205–206, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528234.

Harding:1966:BRF

[609] E. F. Harding. Book review: Formal Properties of Newspaper Dutch by J. A. Th. M. van Berckel, H. B. Corstius, R. J. Mokken, A. van Wijngaarden. Biometrics, 22(1):206, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528235.

Urquhart:1966:BRP

[610] N. Scott Urquhart. Book review: Patterns in the Balance of Nature by C. B. Williams. Biometrics, 22(1):206–207, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528236.

Anonymous:1966:BRa

[611] Anonymous. [Books received]. Biometrics, 22(1):207–208, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528237.

Anonymous:1966:Aa

[612] Anonymous. Abstracts. Biometrics, 22(1):209–212, March 1966. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528238.

Anonymous:1966:BSa

[613] Anonymous. The Biometric Society. Biometrics, 22(1):213–229, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528239.

Anonymous:1966:NAa

[614] Anonymous. News and announcements. Biometrics, 22(1):230–232, March 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528240.

Anonymous:1966:FMb

[615] Anonymous. Front matter. Biometrics, 22(2):i–ii, June 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528515.

Yates:1966:CSR

[616] F. Yates. Computers, the second revolution in statistics. Biometrics, 22(2):233–251, June 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528516.

Carlson:1966:LRB

[617] F. D. Carlson, E. Sobel, and G. S. Watson. Linear relationships between variables affected by errors. Biometrics, 22(2):252–267, June 1966. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528517.

Weiner:1966:EVL

[618] John M. Weiner and Olive Jean Dunn. Elimination of variates in linear discrimination problems. Biometrics, 22(2):268–275, June 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528518.

Day:1966:FCL

[619] N. E. Day. Fitting curves to longitudinal data. Biometrics, 22(2):276– 291, June 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528519.

Atkinson:1966:DST

[620] G. F. Atkinson. Designs for sequences of treatments with carry-over effects. Biometrics, 22(2):292–309, June 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528520.

Sheps:1966:CRU

[621] Mindel C. Sheps. Characteristics of a ratio used to estimate failure rates: Occurrences per person year of exposure. Biometrics, 22(2):310–321, June 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528521.

Brown:1966:PQA

[622] Byron Wm. Brown, Jr. Planning a quantal assay of potency. Biometrics, 22(2):322–329, June 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528522.

Cormack:1966:TEC

[623] R. M. Cormack. A test for equal catchability. Biometrics, 22(2):330– 342, June 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528523.

Rouvier:1966:LCP

[624] R. Rouvier. L’analyse en composantes principales: Son utilisation en génétique et ses rapports avec l’analyse discriminatoire. (French) [Principal component analysis: Its use in genetics and its relationship to discriminatory analysis]. Biometrics, 22(2):343–357, June 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528524.

Fraser:1966:SMM

[625] A. R. Fraser and M. Kovats. Stereoscopic models of multivariate statistical data. Biometrics, 22(2):358–367, June 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528525.

Robinson:1966:BIB

[626] J. Robinson. Balanced incomplete block designs with double grouping of blocks into replications. Biometrics, 22(2):368–373, June 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528526.

Panse:1966:ECY

[627] V. G. Panse, M. Rajagopalan, and S. S. Pillai. Estimation of crop yields for small areas. Biometrics, 22(2):374–384, June 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528527.

Sen:1966:STE

[628] A. R. Sen, R. P. Chakrabarty, and A. R. Sarkar. Sampling techniques for estimation of incidence of red spider mite on tea crop in North-East India. Biometrics, 22(2):385–403, June 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528528.

Mantel:1966:NRP

[629] Nathan Mantel. 218. Note: F -ratio probabilities from binomial tables. Biometrics, 22(2):404–407, June 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/ stable/2528529. See corrections [797].

Pasternack:1966:NDS

[630] Bernard S. Pasternack and Nathan Mantel. 219. Note: a deficiency in the summation of chi procedure. Biometrics, 22(2):407–409, June 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528530.

Tallis:1966:NMM

[631] G. M. Tallis. 220. Note: a migration model. Biometrics, 22(2):409– 412, June 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528531.

Weiling:1966:OOH

[632] F. Weiling. Obituary: Ottokar Heinisch. Biometrics, 22(2):413–414, June 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528532.

Anonymous:1966:C

[633] Anonymous. Correction. Biometrics, 22(2):414, June 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528533.

Anonymous:1966:Ab

[634] Anonymous. Abstracts. Biometrics, 22(2):415–421, June 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528534.

Anonymous:1966:BSb

[635] Anonymous. The Biometric Society. Biometrics, 22(2):422–435, June 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528535.

Anonymous:1966:NAb

[636] Anonymous. News and announcements. Biometrics, 22(2):436–438, June 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528536.

Anonymous:1966:FMc

[637] Anonymous. Front matter. Biometrics, 22(3):i–ii, September 1966. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528180.

Gardner:1966:AIV

[638] C. O. Gardner and S. A. Eberhart. Analysis and interpretation of the variety cross diallel and related populations. Biometrics, 22(3):439–452, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528181.

Hanson:1966:EPI

[639] W. D. Hanson. Effects of partial isolation (distance), migration, and different fitness requirements among environmental pockets upon steady state gene frequencies. Biometrics, 22(3):453–468, September 1966. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528182.

Misra:1966:VAG

[640] R. K. Misra. Vectorial analysis for genetic clines in body dimensions in populations of Drosophila subobscura Coll. and a comparison with those of D. robusta Sturt. Biometrics, 22(3):469–487, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528183.

Namkoong:1966:SAI

[641] Gene Namkoong. Statistical analysis of introgression. Biometrics, 22 (3):488–502, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528184.

Yates:1966:GCP

[642] F. Yates and A. J. B. Anderson. A general computer programme for the analysis of factorial experiments. Biometrics, 22(3):503–524, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528185.

Federer:1966:AMC

[643] W. T. Federer and M. Zelen. Analysis of multifactor classifications with unequal numbers of observations. Biometrics, 22(3):525–552, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528186.

Blischke:1966:VEV

[644] W. R. Blischke. Variances of estimates of variance components in a threeway classification. Biometrics, 22(3):553–565, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528187.

Cohen:1966:NCD

[645] A. Clifford Cohen, Jr. A note on certain discrete mixed distributions. Biometrics, 22(3):566–572, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528188.

Bliss:1966:FRH

[646] C. I. Bliss and A. T. James. Fitting the rectangular hyperbola. Biometrics, 22(3):573–602, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528189.

Dubey:1966:DCL

[647] Satya D. Dubey. On the determination of confidence limits of an index. Biometrics, 22(3):603–609, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528190.

Shorack:1966:GPU

[648] Galen R. Shorack. Graphical procedures for using distribution-free methods in the estimation of relative potency in dilution (-direct) assays. Biometrics, 22(3):610–619, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528191.

Haight:1966:NML

[649] Frank A. Haight and Gary R. Fisher. 221. Note: Maximum likelihood estimation for the truncated Poisson. Biometrics, 22(3):620–623, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528192.

Jones:1966:NEV

[650] Marshall B. Jones. 222. Note: On the evaluation of variability in isogenic hybrids. Biometrics, 22(3):623–628, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528193.

Goodchild:1966:NAN

[651] N. A. Goodchild. 223. Note: Applications of nonorthogonal designs to situations where treatments or blocks are of unequal status or size. Biometrics, 22(3):629–631, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528194.

Williams:1966:NET

[652] D. A. Williams. 224. Note: Errors of treatment comparisons when observations are missing from a randomised block experiment with additional replication of a control treatment. Biometrics, 22(3):632–633, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528195.

Conagin:1966:OCG

[653] A. Conagin. Obituary: Constantino Goncalves Fraga, Jr.: 1907–66. Biometrics, 22(3):634–635, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528196.

Thompson:1966:BRM

[654] W. A. Thompson, Jr. Book review: The Method of Paired Comparisons by H. A. David. Biometrics, 22(3):636, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528197.

Hewlett:1966:BRE

[655] P. S. Hewlett. Book review: Experiments in Plant Hybridisation by Gregor Mendel. Biometrics, 22(3):636–638, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528198.

Bradley:1966:BRS

[656] Ralph A. Bradley. Book review: Some Statistical and Other Numerical Techniques for Classifying Individuals by P. Macnaughton-Smith. Biometrics, 22(3):638–639, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528199.

Raktoe:1966:BRM

[657] Bisnoedath L. Raktoe. Book review: A Modern Algebra for Biologists by Howard M. Nahikian. Biometrics, 22(3):639–640, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528200.

Hagans:1966:BRB

[658] J. A. Hagans. Book review: Biological Statistics — an Introduction by S. C. Pearce. Biometrics, 22(3):640–641, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528201.

Sprent:1966:BRM

[659] P. Sprent. Book review: Matrix Algebra for the Biological Sciences (Including Applications in Statistics) by S. R. Searle. Biometrics, 22(3):641– 642, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528202.

Balaam:1966:BRSb

[660] L. N. Balaam. Book review: Statistics for Scientists and Engineers by R. Lowell Wine. Biometrics, 22(3):642–643, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528203.

Anonymous:1966:BRb

[661] Anonymous. [Books received]. Biometrics, 22(3):643–644, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528204.

Anonymous:1966:Ac

[662] Anonymous. Abstracts. Biometrics, 22(3):645–651, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528205.

Anonymous:1966:BSc

[663] Anonymous. The Biometric Society. Biometrics, 22(3):652–659, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528206.

Anonymous:1966:NAc

[664] Anonymous. News and announcements. Biometrics, 22(3):660–663, September 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528207.

Anonymous:1966:FMd

[665] Anonymous. Front matter. Biometrics, 22(4):i–ii, December 1966. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528066.

Anonymous:1966:VI

[666] Anonymous. Volume information. Biometrics, 22(4):i–xxxi, December 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528065.

Zippin:1966:UCV

[667] Calvin Zippin and Peter Armitage. Use of concomitant variables and incomplete survival information in the estimation of an exponential survival parameter. Biometrics, 22(4):665–672, December 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528067.

Lellouch:1966:EPM

[668] J. Lellouch and A. Wambersie. Estimation par la méthode du maximum de vraisemblance des courbes de survie de microorganismes irradiés. (French) [Estimation by the maximum likelihood method of survival curves of irradiated microorganisms]. Biometrics, 22(4):673–683, December 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528068.

Tootill:1966:CSA

[669] J. P. R. Tootill, C. Nancy Hebert, and A. G. Adams. Comparative sampling acceptance schemes for testing the antigenicity of vaccines. Biometrics, 22(4):684–705, December 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528069.

Das:1966:IBD

[670] M. N. Das and G. A. Kulkarni. Incomplete block designs for bio-assays. Biometrics, 22(4):706–729, December 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2528070. See correction [1059].

Addelman:1966:SCC

[671] S. Addelman, D. W. Gaylor, and R. E. Bohrer. Sequences of combination chemotherapy experiments. Biometrics, 22(4):730–746, December 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528071.

Hopkins:1966:SCM

[672] J. W. Hopkins. Some considerations in multivariate allometry. Biometrics, 22(4):747–760, December 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528072.

Nissen-Meyer:1966:AEA

[673] S. Nissen-Meyer. Analysis of effects of antibiotics on bacteria by means of stochastic models. Biometrics, 22(4):761–780, December 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528073.

DeGroot:1966:CBS

[674] M. H. DeGroot and C. C. Li. Correlations between similar sets of measurements. Biometrics, 22(4):781–790, December 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528074.

Lu:1966:PPA

[675] K. H. Lu. A path-probability approach to irreversible Markov chains with an application in studying the dental caries process. Biometrics, 22(4):791–809, December 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528075.

Cole:1966:AMA

[676] J. W. L. Cole and James E. Grizzle. Applications of multivariate analysis of variance to repeated measurements experiments. Biometrics, 22(4):810–828, December 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528076.

Cunningham:1966:ATI

[677] E. P. Cunningham and C. R. Henderson. Analytical techniques for incomplete block experiments. Biometrics, 22(4):829–842, December 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528077.

Schutz:1966:EFB

[678] W. M. Schutz and C. Clark Cockerham. The effect of field blocking on gain from selection. Biometrics, 22(4):843–863, December 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528078.

Eberhart:1966:GMG

[679] S. A. Eberhart and C. O. Gardner. A general model for genetic effects. Biometrics, 22(4):864–881, December 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528079.

Goodall:1966:NSI

[680] David W. Goodall. A new similarity index based on probability. Biometrics, 22(4):882–907, December 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528080.

Dunn:1966:PCC

[681] Olive Jean Dunn and Paul D. Varady. Probabilities of correct classification in discriminant analysis. Biometrics, 22(4):908–924, December 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528081.

Holgate:1966:CMA

[682] P. Holgate. Contributions to the mathematics of animal trapping. Biometrics, 22(4):925–936, December 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528082.

Eisen:1966:NQT

[683] E. J. Eisen. 225. Note: The quasi-F test for an unnested fixed factor in an unbalanced hierarchal design with a mixed model. Biometrics, 22(4):937–942, December 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528083.

Anonymous:1966:Ad

[684] Anonymous. Abstracts. Biometrics, 22(4):943–962, December 1966. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528084.

Anonymous:1966:CMC

[685] Anonymous. Correction: Multiple covariate analysis. Biometrics, 22(4): 962, December 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528085. See

[565].

Anonymous:1966:BSd

[686] Anonymous. The Biometric Society. Biometrics, 22(4):963–971, December 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528086.

Anonymous:1966:NAd

[687] Anonymous. News and announcements. Biometrics, 22(4):972, December 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528087.

Anonymous:1967:FMa

[688] Anonymous. Front matter. Biometrics, 23(1):i–ii, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528276.

Pasternack:1967:SES

[689] Bernard S. Pasternack and Anthony Liuzzi. Some experimental sampling results for regression analysis applied to gamma ray spectrometer data, I. Biometrics, 23(1):11–26, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528278.

Plackett:1967:CTA

[690] R. L. Plackett and P. S. Hewlett. A comparison of two approaches to the construction of models for quantal responses to mixtures of drugs. Biometrics, 23(1):27–44, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528279.

Gilbert:1967:ACA

[691] N. Gilbert. Additive combining abilities fitted to plant breeding data. Biometrics, 23(1):45–49, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528280.

Kuzma:1967:CTL

[692] Jan W. Kuzma. A comparison of two life table methods. Biometrics, 23(1):51–64, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528281.

Mantel:1967:RPA

[693] Nathan Mantel. Ranking procedures for arbitrarily restricted observation. Biometrics, 23(1):65–78, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528282.

Maritz:1967:UGE

[694] J. S. Maritz and A. H. Munro. On the use of the generalised extremevalue distribution in estimating extreme percentiles. Biometrics, 23(1): 79–103, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528283.

Hartley:1967:EVC

[695] H. O. Hartley. Expectations, variances and covariances of ANOVA mean squares by “synthesis”. Biometrics, 23(1):105–114, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528284. See correction [796].

Bhattacharya:1967:SMR

[696] C. G. Bhattacharya. A simple method of resolution of a distribution into Gaussian components. Biometrics, 23(1):115–135, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528285.

Urbakh:1967:STD

[697] V. Yu. Urbakh. Statistical testing of differences in causal behaviour of two morphologically indistinguishable objects. Biometrics, 23(1):137– 143, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528286.

Lellouch:1967:MLT

[698] D. Schwartz et J. Lellouch. Les manquants dans L’Essai therapeutique. Biometrics, 23(1):145–152, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528287.

Cramer:1967:NET

[699] Elliot M. Cramer. 226. Note: Equivalence of two methods of computing discriminant function coefficients. Biometrics, 23(1):153, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528288.

Mantel:1967:NEM

[700] Nathan Mantel and Samuel W. Greenhouse. 227. Note: Equivalence of maximum likelihood and the method of moments in probit analysis. Biometrics, 23(1):154–157, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528289.

Hedges:1967:NDE

[701] A. J. Hedges. 228. Note: On the dilution errors involved in estimating bacterial numbers by the plating method. Biometrics, 23(1):158–159, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528290.

Cornfield:1967:MBC

[702] Jerome Cornfield. The meaning of bioassay: a comment. Biometrics, 23(1):160–162, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2528291. See [552].

F:1967:BRE

[703] W. T. F. Book review: Equality of Educational Opportunity, by J. S. Coleman, E. Q. Campbell, C. J. Hobson, J. McPartland, A. M. Mood, F. D. Weinfeld, R. L. York. Biometrics, 23(1):163, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528292.

Wallace:1967:BRF

[704] Bruce Wallace. Book review: The Foundations of Genetics, by F. A. E. Crew. Biometrics, 23(1):163, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528293.

Balaam:1967:BRI

[705] L. N. Balaam. Book review: Introduction to Statistical Methods, by Sylvain Ehrenfeld, Sebastian B. Littauer. Biometrics, 23(1):164, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528295.

F:1967:BRA

[706] W. T. F. Book review: Annual Report, 1 July, 1965-30 June, 1966, by G. B. Darling, H. Maki, I. Nagai. Biometrics, 23(1):164, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528294.

Campbell:1967:BRE

[707] R. C. Campbell. Book review: The Estimation of Probabilities: an Essay on Modern Bayesian Methods, by I. J. Good. Biometrics, 23(1):164–165, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528296.

Raktoe:1967:BRG

[708] B. L. Raktoe. Book review: Green’s Function Methods in Probability Theory, by J. Keilson. Biometrics, 23(1):165–166, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528297.

F:1967:BRR

[709] W. T. F. Book review: Replication: an Approach to the Analysis of Data from Complex Surveys, by P. J. McCarthy. Biometrics, 23(1):166, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528298.

Silveira:1967:BRA

[710] María Isabel Silveira and Carlos A. Gonzalez. Book review: Applications de la Statistique à la Demographie et à la Biologie. (Traité du calcul des probabilités et ses applications, Tome 3, Fascicule III), by R. Risser, C. Traynard. Biometrics, 23(1):166–167, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528299.

F:1967:BRHa

[711] W. T. F. Book review: Historical Statistics of Canada, by M. C. Urquhart, K. A. H. Buckley. Biometrics, 23(1):168, March 1967. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528300.

Holgate:1967:BRM

[712] P. Holgate. Book review: Mathematische Statistik, by B. L. van der Waerden. Biometrics, 23(1):168–169, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528301.

F:1967:BRBa

[713] W. T. F. Book review: Bibliography on Time Series and Stochastic Processes, by H. Wold. Biometrics, 23(1):169, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528302.

Anonymous:1967:AEV

[714] Anonymous. Annual epidemiological and vital statistics, 1961, by world health organization. Biometrics, 23(1):169–170, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528303.

Anonymous:1967:WHS

[715] Anonymous. World health statistics annual, 1962, by world health organization. Biometrics, 23(1):170–171, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528304.

F:1967:BRSa

[716] W. T. F. Book review: Sampling Methods in Morbidity Surveys and Public Health Investigations, by World Health Organization. Biometrics, 23 (1):171, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528305.

Anonymous:1967:BRa

[717] Anonymous. [Books received]. Biometrics, 23(1):171–173, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528306.

Anonymous:1967:Aa

[718] Anonymous. Abstracts. Biometrics, 23(1):174–176, March 1967. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528307.

Anonymous:1967:BSa

[719] Anonymous. The Biometric Society. Biometrics, 23(1):177–186, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528308.

Anonymous:1967:NAa

[720] Anonymous. News and announcements. Biometrics, 23(1):187–188, March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528309.

Karas:1967:PFW

[721] James Karas and I. Richard Savage. Publications of Frank Wilcoxon (1892–1965). Biometrics, 23(1):iv + 1–10 (1 plate), March 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528277.

Anonymous:1967:FMb

[722] Anonymous. Front matter. Biometrics, 23(2):i–ii, June 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528154.

Mead:1967:MME

[723] R. Mead. A mathematical model for the estimation of inter-plant competition. Biometrics, 23(2):189–205, June 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2528155. See corrections [928].

Eberhardt:1967:SDD

[724] L. L. Eberhardt. Some developments in ‘distance sampling’. Biometrics, 23(2):207–216, June 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528156.

Krause:1967:PSG

[725] G. F. Krause, P. B. Siegel, and D. C. Hurst. A probability structure for growth curves. Biometrics, 23(2):217–225, June 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528157.

Kodlin:1967:NRT

[726] D. Kodlin. A new response time distribution. Biometrics, 23(2):227–239, June 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528158. See correction

[1062].

Kemp:1967:CDS

[727] C. D. Kemp. On a contagious distribution suggested for accident data. Biometrics, 23(2):241–255, June 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528159.

Boen:1967:UPI

[728] James R. Boen and Byron Wm. Brown, Jr. Use of prior information to design a routine parallel line assay. Biometrics, 23(2):257–267, June 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528160.

Gart:1967:GOT

[729] John J. Gart and George H. Weiss. Graphically oriented tests for host variability in dilution experiments. Biometrics, 23(2):269–284, June 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528161.

Gross:1967:EUB

[730] Alan J. Gross and Nathan Mantel. The effective use of both positive and negative controls in screening experiments. Biometrics, 23(2):285– 295, June 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528162.

Taylor:1967:VOP

[731] J. Taylor. The value of orthogonal polynomials in the analysis of changeover trials with diary cows. Biometrics, 23(2):297–311, June 1967. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528163.

Day:1967:GML

[732] N. E. Day and D. F. Kerridge. A general maximum likelihood discriminant. Biometrics, 23(2):313–323, June 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528164.

Nasoetion:1967:SSP

[733] A. H. Nasoetion, C. Clark Cockerham, and D. F. Matzinger. Simultaneous selfing and partial diallel test crossing II. an evaluation of two methods of estimation of genetic and environmental variance. Biometrics, 23(2):325–334, June 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528165.

Kale:1967:ISI

[734] B. K. Kale and T. A. Bancroft. Inference for some incompletely specified models involving normal approximations to discrete data. Biometrics, 23(2):335–348, June 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528166.

Naylor:1967:NSS

[735] Alfred F. Naylor. 229. Note: Small sample considerations in combining 2 × 2 tables. Biometrics, 23(2):349–356, June 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528167.

Berger:1967:NER

[736] Agnes Berger and Ruth Z. Gold. 230. Note: On estimating recessive frequencies from truncated samples. Biometrics, 23(2):356–360, June 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528168.

Soller:1967:NOE

[737] M. Soller and A. Genizi. 231. Note: Optimum experimental designs for realized heritability estimates. Biometrics, 23(2):361–365, June 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528169.

Sutter:1967:OSL

[738] J. Sutter. Obituary: Sully Ledermann: 1915–67. Biometrics, 23(2):366– 367, June 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528170.

Sampford:1967:EN

[739] M. R. Sampford. Editorial note. Biometrics, 23(2):368, June 1967. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528171.

Searle:1967:BRA

[740] S. R. Searle. Book review: Applied Regression Analysis, by N. R. Draper, H. Smith. Biometrics, 23(2):369–370, June 1967. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528172.

Gramont:1967:BRP

[741] María Isabel Silveira Gramont. Book review: Probability and Statistics, by H. Freudenthal. Biometrics, 23(2):370–371, June 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528173.

Geisser:1967:BRL

[742] Seymour Geisser. Book review: Logic of Statistical Inference, by Ian Hacking. Biometrics, 23(2):371–373, June 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528174.

Robson:1967:BRM

[743] D. S. Robson. Book review: Methods in Drug Evaluation, by P. Mantegazza, F. Piccinini. Biometrics, 23(2):373, June 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528175.

Anonymous:1967:Ab

[744] Anonymous. Abstracts. Biometrics, 23(2):374–386, June 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528176.

Anonymous:1967:BSb

[745] Anonymous. The Biometric Society. Biometrics, 23(2):387–394, June 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528177.

Anonymous:1967:NAb

[746] Anonymous. News and announcements. Biometrics, 23(2):395, June 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528178.

Anonymous:1967:FMc

[747] Anonymous. Front matter. Biometrics, 23(3):i–ii, September 1967. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528008.

Li:1967:BIP

[748] C. C. Li. A biometrics invited paper. Genetic equilibrium under selection. Biometrics, 23(3):397–484, September 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528009.

Keyfitz:1967:MMD

[749] Nathan Keyfitz and Edmund M. Murphy. Matrix and multiple decrement in population analysis. Biometrics, 23(3):485–503, September 1967. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528010.

Berry:1967:MMR

[750] G. Berry. A mathematical model relating plant yield with arrangement for regularly spaced crops. Biometrics, 23(3):505–515, September 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528011.

Bennett:1967:GCE

[751] J. H. Bennett. A general class of enumerations arising in genetics. Biometrics, 23(3):517–537, September 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528012.

Rhyne:1967:MCS

[752] A. L. Rhyne and R. G. D. Steel. A multiple comparisons sign test: All pairs of treatments. Biometrics, 23(3):539–549, September 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528013.

Brown:1967:UCV

[753] G. H. Brown. The use of correlated variables for preliminary culling. Biometrics, 23(3):551–562, September 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528014.

Moore:1967:UNL

[754] R. H. Moore and R. K. Zeigler. The use of non-linear regression methods for analyzing sensitivity and quantal response data. Biometrics, 23(3):563–566, September 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528015.

Mantel:1967:NAF

[755] Nathan Mantel. 232. Note: Assumption-free estimators using U statistics and a relationship to the jackknife method. Biometrics, 23(3):567–571, September 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528016.

Rayner:1967:NSS

[756] A. A. Rayner. 233. Note: The square-summing check on the main effects and interactions in a 2^{n} factorial experiment as calculated by Yates’s algorithm. Biometrics, 23(3):571–573, September 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528017.

Good:1967:NCY

[757] I. J. Good. 234. Note: Checks on Yates’s algorithm. Biometrics, 23(3): 573, September 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528018.

Preece:1967:NCG

[758] D. A. Preece. 235. Note: Cyclic generation of Robinson’s balanced incomplete block designs. Biometrics, 23(3):574–578, September 1967. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528019.

Berenblut:1967:NAC

[759] I. I. Berenblut. 236. Note: The analysis of changeover designs with complete balance for first residual effects. Biometrics, 23(3):578–580, September 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528020.

Anonymous:1967:FSP

[760] Anonymous. French summaries of papers in the June 1967 issue. Biometrics, 23(3):581–586, September 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528021.

Finney:1967:MBC

[761] D. J. Finney. [The meaning of bioassay: a comment]. Biometrics, 23(3):587–588, September 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2528022. See [702, 552].

F:1967:BRHb

[762] W. T. F. Book review: Handbook of Tables for Probability and Statistics by W. H. Beyer. Biometrics, 23(3):589–590, September 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528023.

Searle:1967:BRP

[763] S. R. Searle. Book review: Principles of Statistics by M. G. Bulman. Biometrics, 23(3):590–591, September 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528024.

F:1967:BRM

[764] W. T. F. Book review: The MAA Problem Book II. Annual High School Contests of the MAA 1961–1965 by C. T. Salkind. Biometrics, 23(3):591, September 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528025.

F:1967:BRSb

[765] W. T. F. Book review: Statistics in Action. Readings in Business and Economic Statistics by T. J. Sielaff. Biometrics, 23(3):591–592, September 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528026.

F:1967:BRSc

[766] W. T. F. Book review: Statistical Methods in Malaria Eradication by S. Swaroop. Biometrics, 23(3):592–593, September 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528027.

Anonymous:1967:BRb

[767] Anonymous. [Books received]. Biometrics, 23(3):593–595, September 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528028.

Anonymous:1967:Ac

[768] Anonymous. Abstracts. Biometrics, 23(3):596–607, September 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528029.

Anonymous:1967:BSc

[769] Anonymous. The Biometric Society. Biometrics, 23(3):608–619, September 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528030.

Anonymous:1967:NAc

[770] Anonymous. News and announcements. Biometrics, 23(3):620–621, September 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528031.

Anonymous:1967:FMd

[771] Anonymous. Front matter. Biometrics, 23(4):i–ii, December 1967. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528416.

Anonymous:1967:VI

[772] Anonymous. Volume information. Biometrics, 23(4):i–xxviii, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528415.

Gower:1967:CSM

[773] J. C. Gower. A comparison of some methods of cluster analysis. Biometrics, 23(4):623–637, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528417.

Lachenbruch:1967:AUM

[774] Peter A. Lachenbruch. An almost unbiased method of obtaining confidence intervals for the probability of misclassification in discriminant analysis. Biometrics, 23(4):639–645, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528418.

Goodall:1967:DMC

[775] D. W. Goodall. The distribution of the matching coefficient. Biometrics, 23(4):647–656, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528419.

Pielou:1967:TRM

[776] E. C. Pielou. A test for random mingling of the phases of a mosaic. Biometrics, 23(4):657–670, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528420.

Mazumdar:1967:SVD

[777] M. Mazumdar and K. L. Kussmaul. A study of the variability due to coincident passage in an electronic blood cell counter. Biometrics, 23(4):671– 684, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528421.

Pettigrew:1967:RTP

[778] Hugh M. Pettigrew and William C. Mohler. A rapid test for the Poisson distribution using the range. Biometrics, 23(4):685–692, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528422.

Burnett-Hall:1967:ISC

[779] D. G. Burnett-Hall and W. A. O’N. Waugh. Indices of synchrony in cellular cultures. Biometrics, 23(4):693–716, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528423.

Cornell:1967:ESE

[780] Richard G. Cornell and Janace A. Speckman. Estimation for a simple exponential model. Biometrics, 23(4):717–737, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528424.

Mantel:1967:AKM

[781] Nathan Mantel. Adaptation of Kärber’s method for estimating the exponential parameter from quantal data, and its relationship to birth, death, and branching processes. Biometrics, 23(4):739–746, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528425.

Patwary:1967:AQR

[782] Kamini M. Patwary and K. David C. Haley. Analysis of quantal response assays with dosage errors. Biometrics, 23(4):747–760, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528426.

Freeman:1967:UCB

[783] G. H. Freeman. The use of cyclic balanced incomplete block designs for directional seed orchards. Biometrics, 23(4):761–778, December 1967.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528427.

Rees:1967:SDU

[784] D. H. Rees. Some designs of use in serology. Biometrics, 23(4):779–791, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528428.

Robinson:1967:ISP

[785] J. Robinson. Incomplete split plot designs. Biometrics, 23(4):793–802, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528429.

Clarke:1967:FWB

[786] G. M. Clarke. Four-way balanced designs based on Youden squares with 5, 6, or 7 treatments. Biometrics, 23(4):803–812, December 1967. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528430.

Das:1967:CAS

[787] M. N. Das and P. S. Rao. Construction and analysis of some new series of confounded asymmetrical factorial designs. Biometrics, 23(4):813–822, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528431.

Quartermain:1967:STS

[788] A. R. Quartermain and A. E. Freeman. Some transformations of scale and the estimation of genetic parameters from daughter–dam regression. Biometrics, 23(4):823–834, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528432.

Epstein:1967:BET

[789] Benjamin Epstein. Bacterial extinction time as an extreme value phenomenon. Biometrics, 23(4):835–839, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528433.

Lancaster:1967:QCP

[790] H. O. Lancaster. 237. Query: The combination of probabilities. Biometrics, 23(4):840–842, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528434.

Allaire:1967:NES

[791] F. R. Allaire. 238. Note: Expected selection differential for positive directional selection on normal variables within sets of finite subpopulations. Biometrics, 23(4):842–845, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528435.

Likes:1967:NSS

[792] Jiří Likeš. 239. Note: Sample size for the estimation of means of normal populations. Biometrics, 23(4):846–849, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528436.

Mather:1967:OGM

[793] Kenneth Mather. Obituary: Gordon M. L. Haskell; 1920–1967. Biometrics, 23(4):850–851, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528437.

Anonymous:1967:CCP

[794] Anonymous. Correction: Computing procedures for estimating components of variance in the two-way classification, mixed model. Biometrics, 23(4):852–853, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2528438. See [150].

Anonymous:1967:CEM

[795] Anonymous. Correction: The estimation of mortality and recruitment from a single-tagging experiment. Biometrics, 23(4):853, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528439. See [527].

Anonymous:1967:CEV

[796] Anonymous. Correction: Expectations, variances and covariances of ANOVA means squares by ‘synthesis’. Biometrics, 23(4):853, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528440. See [695].

Anonymous:1967:CNR

[797] Anonymous. Corrections: 218. Note. F -ratio probabilities from binomial tables. Biometrics, 23(4):854, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528441. See [629].

F:1967:BRSd

[798] W. T. F. Book review: Supplement to Tables for Testing Significance in a 2 × 2 Contingency Table. Five and One Per Cent Significance Points for 41 ≤ A ≤ 50, B ≤ A by B. M. Bennett, C. Horst. Biometrics, 23(4):855, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528442.

F:1967:BRT

[799] W. T. F. Book review: Tables of Bivariate Random Normal Deviates by P. V. Krishna Iyer, P. S. Simha. Biometrics, 23(4):855, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528443.

F:1967:BRI

[800] W. T. F. Book review: International Advisory Group of Forest Statisticians, 2nd Conference, Stockholm, 27th September–1st October, 1965 by B. Matern. Biometrics, 23(4):855–857, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528444.

Weeks:1967:BRS

[801] D. L. Weeks. Book review: Simultaneous Statistical Inference by Rupert G. Miller, Jr. Biometrics, 23(4):857–858, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528445.

F:1967:BRC

[802] W. T. F. Book review: Classical and Contagious Discrete Distributions by G. P. Patil. Biometrics, 23(4):858–859, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528446.

F:1967:BRP

[803] W. T. F. Book review: Public Health and Population Change by M. C. Sheps, J. C. Ridley. Biometrics, 23(4):860, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528447.

F:1967:BRSe

[804] W. T. F. Book review: The Selected Papers of E. S. Pearson by E. S. Pearson, L. H. C. Tippett. Biometrics, 23(4):860, December 1967.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528448.

F:1967:BRBb

[805] W. T. F. Book review: Biometeorology — Volume 2, by S. W. Tromp, W. H. Weihe. Biometrics, 23(4):861, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528449.

Anonymous:1967:BRc

[806] Anonymous. [Books received]. Biometrics, 23(4):861–862, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528450.

Anonymous:1967:Ad

[807] Anonymous. Abstracts. Biometrics, 23(4):863–877, December 1967. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528451.

Anonymous:1967:BSd

[808] Anonymous. The Biometric Society. Biometrics, 23(4):878–887, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528452.

Anonymous:1967:NAd

[809] Anonymous. News and announcements. Biometrics, 23(4):888–893, December 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528453.

Anonymous:1968:FMa

[810] Anonymous. Front matter. Biometrics, 24(1):i–ii, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528455.

Finney:1968:TBU

[811] D. J. Finney. Teaching biometry in the university. Biometrics, 24(1):1– 12, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528456.

Cunningham:1968:IPE

[812] E. P. Cunningham and C. R. Henderson. An iterative procedure for estimating fixed effects and variance components in mixed model situations. Biometrics, 24(1):13–25, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2528457. See corrections [1043].

Koch:1968:SAS

[813] Gary G. Koch and Pranab Kumar Sen. Some aspects of the statistical analysis of the “mixed model”. Biometrics, 24(1):27–48, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528458.

Grossman:1968:CAU

[814] M. Grossman and G. A. E. Gall. Covariance analysis with unequal subclass numbers: Component-estimation in quantitative genetics. Biometrics, 24(1):49–59, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528459.

Balaam:1968:TPD

[815] L. N. Balaam. A two-period design with t^{2} experimental units. Biometrics, 24(1):61–73, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528460.

Berkson:1968:AML

[816] Joseph Berkson. Application of minimum logit χ^{2} estimate to a problem of Grizzle with a notation on the problem of ‘no interaction’. Biometrics, 24(1):75–95, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528461.

Bartko:1968:ESF

[817] John J. Bartko, Samuel W. Greenhouse, and Clifford S. Patlak. On expectations of some functions of Poisson variates. Biometrics, 24(1): 97–102, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528462.

Anderson:1968:CST

[818] J. A. Anderson and J. A. Boyle. A comparison of statistical techniques in the differential diagnosis of non-toxic goitre. Biometrics, 24(1):103– 116, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528463.

Rossing:1968:CCD

[819] Robert G. Rossing and M. Bryan Danford. A comparison of continuous distributions of parameters of exponential decay curves. Biometrics, 24(1):117–134, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528464.

Gates:1968:LTM

[820] Charles E. Gates, William H. Marshall, and David P. Olson. Line transect method of estimating grouse population densities. Biometrics, 24(1):135– 145, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528465.

Pollard:1968:MTG

[821] J. H. Pollard. The multi-type Galton–Watson process in a genetical context. Biometrics, 24(1):147–158, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528466.

Choi:1968:TSD

[822] S. C. Choi. Truncated sequential designs for clinical trials based on Markov chains. Biometrics, 24(1):159–168, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528467.

Tallis:1968:SOG

[823] G. M. Tallis. Selection for an optimum growth curve. Biometrics, 24(1): 169–177, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528468.

Wyshak:1968:NDA

[824] Grace Wyshak. 239. Note: Distribution among relatives of genotypes for twinning. Biometrics, 24(1):179–185, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528469.

Cohen:1968:NEE

[825] Joel E. Cohen. 240. Note: On estimating the equilibrium and transition probabilities of a finite-state Markov chain from the same data. Biometrics, 24(1):185–187, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528470.

Cannings:1968:NEU

[826] C. Cannings. 241. Note: Equilibrium under selection at a multiallelic sexlinked locus. Biometrics, 24(1):187–189, March 1968. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528471. See corrections [929].

Goodman:1968:NMO

[827] Major M. Goodman. 242. Note: a measure of ‘overall variability’ in populations. Biometrics, 24(1):189–192, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528472.

Hills:1968:NNA

[828] M. Hills. 243. Note: a note on the analysis of growth curves. Biometrics, 24(1):192–196, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528473.

F:1968:BRA

[829] W. T. F. Book review: An Abstract of Statistical Methods in Grassland Research, by M. J. Morris. Biometrics, 24(1):197, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528474.

Michalos:1968:BRMa

[830] Alex C. Michalos. Book review: Mathematics of Choice, by Ivan Niven. Biometrics, 24(1):197, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528475.

Raktoe:1968:BRE

[831] B. L. Raktoe. Book review: Exercises in Probability and Statistics, by N. A. Rahman. Biometrics, 24(1):197–199, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528476.

Raktoe:1968:BRW

[832] B. L. Raktoe. Book review: Wahrscheinlichkeitstheorie, by Hans Richter. Biometrics, 24(1):199–200, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528477.

Michalos:1968:BRMb

[833] A. C. Michalos. Book review: Mathematics and Social Sciences I, by S. Sternberg, V. Capecchi, T. Kloek, C. T. Leenders. Biometrics, 24(1): 200–201, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528478.

F:1968:BRI

[834] W. T. F. Book review: Infectious Diseases: Cases, Deaths and Vaccinations, by WHO. Biometrics, 24(1):201, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528479.

Anonymous:1968:BRa

[835] Anonymous. [Books received]. Biometrics, 24(1):201–202, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528480.

Anonymous:1968:Aa

[836] Anonymous. Abstracts. Biometrics, 24(1):203–238, March 1968. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528481.

Anonymous:1968:BSa

[837] Anonymous. The Biometric Society. Biometrics, 24(1):239–243, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528482.

Anonymous:1968:NAa

[838] Anonymous. News and announcements. Biometrics, 24(1):244–245, March 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528483.

Anonymous:1968:FMb

[839] Anonymous. Front matter. Biometrics, 24(2):i–ii, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528033.

Mertz:1968:CPS

[840] David B. Mertz and Robert B. Davies. Cannibalism of the pupal stage by adult flour beetles: an experiment and a stochastic model. Biometrics, 24(2):247–275, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528034.

Merat:1968:DFI

[841] P. Mérat. Distributions de fréquences, interpretation du déterminisme génétique des caractères quantitatifs et recherche de ‘gènes majeurs’. (French) [Frequency distributions, interpretation of genetic determinism of quantitative traits and search for ‘major genes’]. Biometrics, 24(2): 277–293, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528035.

Cochran:1968:EAS

[842] W. G. Cochran. The effectiveness of adjustment by subclassification in removing bias in observational studies. Biometrics, 24(2):295–313, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528036.

Chapman:1968:APC

[843] D. G. Chapman and Jun mo Nam. Asymptotic power of chi square tests for linear trends in proportions. Biometrics, 24(2):315–327, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528037. See corrections [1042].

Bhapkar:1968:ACT

[844] V. P. Bhapkar. On the analysis of contingency tables with a quantitative response. Biometrics, 24(2):329–338, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528038.

Miettinen:1968:MPD

[845] Olli S. Miettinen. The matched pairs design in the case of all-or-none responses. Biometrics, 24(2):339–352, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528039.

Pasternack:1968:SES

[846] Bernard S. Pasternack and Anthony Liuzzi. Some experimental sampling results for regression analysis applied to gamma ray spectrometer data, II. Biometrics, 24(2):353–362, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528040.

Danford:1968:NEC

[847] M. Bryan Danford and Harry M. Hughes. The negative exponential with cumulative error. Biometrics, 24(2):363–376, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528041.

Wyshak:1968:EPT

[848] Grace Wyshak and Colin White. Estimation of parameters of a truncated Poissonian binomial. Biometrics, 24(2):377–388, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528042.

Afonja:1968:AGB

[849] Biyi Afonja. Analysis of a group of balanced block experiments having error variance and some treatments in common. Biometrics, 24(2):389– 400, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528043.

Causton:1968:AMP

[850] D. R. Causton. Adjoint matrices for polynomial regression. Biometrics, 24(2):401–411, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2528044. See corrections [927].

Bliss:1968:TCW

[851] C. I. Bliss and S. S. Whitman. A table for computing working angles. Biometrics, 24(2):413–422, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528045.

Tietjen:1968:NTS

[852] Gary L. Tietjen and Roger H. Moore. 244. Note: On testing significance of components of variance in the unbalanced nested analysis of variance. Biometrics, 24(2):423–429, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/ stable/2528046. See corrections [930].

Misra:1968:NST

[853] R. K. Misra. 245. Note: Statistical tests of hypotheses concerning the degree of dominance in monofactorial inheritance. Biometrics, 24(2):429– 434, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528047.

Mantel:1968:NSC

[854] Nathan Mantel. 246. Note: Simultaneous confidence intervals and experimental design with normal correlation. Biometrics, 24(2):434–437, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528048.

Reinhardt:1968:NCE

[855] H. E. Reinhardt. 247. Note: Characterizing the exponential distribution. Biometrics, 24(2):437–439, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528049.

Namkoong:1968:NEN

[856] G. Namkoong and D. L. Miller. 248. Note: Estimation of non-linear parameters for a non-asymptotic function. Biometrics, 24(2):439–440, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528050.

VanVleck:1968:BRM

[857] L. D. Van Vleck. Book review: Manual of Procedures in Quantitative Genetics, by Walter A. Becker. Biometrics, 24(2):441, June 1968. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528051.

Ashton:1968:BRPa

[858] G. C. Ashton. Book review: Psychological Tests and Personnel Decisions, by L. J. Cronbach, C. Gleser. Biometrics, 24(2):442, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528052.

Rodine:1968:BRI

[859] R. H. Rodine. Book review: Introduction to Probability and Statistics, by William Mendenhall. Biometrics, 24(2):442–443, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528053.

Campbell:1968:BRL

[860] R. C. Campbell. Book review: The Linear Hypothesis: a General Theory, by G. A. F. Seber. Biometrics, 24(2):443, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528054.

Mode:1968:BRB

[861] C. J. Mode. Book review: Biomathematics, Vol. I. – Algebra, Geometry, and Calculus, by C. A. B. Smith. Biometrics, 24(2):443–444, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528055.

Schmidt:1968:BRB

[862] R. N. Schmidt. Book review: The Bases of Fortran, by R. E. Smith. Biometrics, 24(2):444–445, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528056.

Michalos:1968:BRMc

[863] Alex C. Michalos. Book review: Mathematics in the Social Sciences and Other Essays, by Richard Stone. Biometrics, 24(2):445, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528057.

Searle:1968:BRH

[864] S. R. Searle. Book review: Handbook of Numerical Matrix Inversion and Solution of Linear Equations, by Joan R. Westlake. Biometrics, 24(2): 445–446, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528058.

Ashton:1968:BRPb

[865] Gordon C. Ashton. Book review: Prediction Analysis, by John R. Wolberg. Biometrics, 24(2):446–447, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528059.

Anonymous:1968:BRb

[866] Anonymous. [Books received]. Biometrics, 24(2):447–448, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528060.

Anonymous:1968:Ab

[867] Anonymous. Abstracts. Biometrics, 24(2):449–458, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528061.

Webb:1968:BS

[868] H. R. Webb. The Biometric Society. Biometrics, 24(2):459–468, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528062.

Anonymous:1968:NAb

[869] Anonymous. News and announcements. Biometrics, 24(2):469–470, June 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528063.

Anonymous:1968:FMc

[870] Anonymous. Front matter. Biometrics, 24(3):i–ii, September 1968. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528311.

Li:1968:FWP

[871] C. C. Li. Fisher, Wright, and path coefficients. Biometrics, 24(3):471– 483, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528312.

Bennett:1968:MSC

[872] J. H. Bennett. Mixed self- and cross-fertilization in a tetrasomic species. Biometrics, 24(3):485–500, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528313.

Ohta:1968:SPN

[873] Tomoko Ohta and Ken-Ichi Kojima. Survival probabilities of new inversions in large populations. Biometrics, 24(3):501–516, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528314.

Gill:1968:PON

[874] J. L. Gill and E. L. Jensen. Probability of obtaining negative estimates of heritability. Biometrics, 24(3):517–526, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528315.

Blischke:1968:VME

[875] W. R. Blischke. Variances of moment estimators of variance components in the unbalanced r-way classification. Biometrics, 24(3):527–539, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528316.

Pike:1968:DCG

[876] M. C. Pike and P. G. Smith. Disease clustering: a generalization of Knox’s approach to the detection of space-time interactions. Biometrics, 24(3):541–556, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528317.

Gart:1968:MAE

[877] John J. Gart. The mathematical analysis of an epidemic with two kinds of susceptibles. Biometrics, 24(3):557–566, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528318.

Bhapkar:1968:HNI

[878] V. P. Bhapkar and Gary G. Koch. On the hypotheses of ‘no interaction’ in contingency tables. Biometrics, 24(3):567–594, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528319.

Frome:1968:MLE

[879] Edward L. Frome and John J. Beauchamp. Maximum likelihood estimation of survival curve parameters. Biometrics, 24(3):595–605, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528320.

Jacquez:1968:LRN

[880] John A. Jacquez, Frances J. Mather, and Charles R. Crawford. Linear regression with non-constant, unknown error variances: Sampling experiments with least squares, weighted least squares and maximum likelihood estimators. Biometrics, 24(3):607–626, September 1968. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528321.

Llewelyn:1968:LLT

[881] F. W. M. Llewelyn. The log(-log) transformation in the analysis of fruit retention records. Biometrics, 24(3):627–638, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528322.

Sprent:1968:LRG

[882] P. Sprent. Linear relationships in growth and size studies. Biometrics, 24(3):639–656, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528323.

Pearce:1968:ALS

[883] S. C. Pearce. Alternatives to a Latin square. Biometrics, 24(3):657–662, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528324.

Davies:1968:SMT

[884] P. Davies. A sequential method of testing the linear trends of responses in dose trials. Biometrics, 24(3):663–677, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528325.

Jolicoeur:1968:IES

[885] Pierre Jolicoeur. Interval estimation of the slope of the major axis of a bivariate normal distribution in the case of a small sample. Biometrics, 24(3):679–682, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528326.

Chaddha:1968:ECD

[886] R. L. Chaddha and L. F. Marcus. An empirical comparison of distance statistics for populations with unequal covariance matrices. Biometrics, 24(3):683–694, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528327.

Emerson:1968:NCO

[887] Phillip L. Emerson. Numerical construction of orthogonal polynomials from a general recurrence formula. Biometrics, 24(3):695–701, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528328. See corrections [1044].

Trappl:1968:NEV

[888] Robert Trappl. 249. Note: Ein einfaches Verfahren zur Erzeugung von Symbolfolgen mit vorgegebener relativer Dyadenkontextredundanz. (German) [249. Note: A simple method for generating symbol sequences with given relative dyad context redundancy]. Biometrics, 24(3):703–707, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528329.

Ireland:1968:NMD

[889] C. T. Ireland and S. Kullback. 250. Note: Minimum discrimination information estimation. Biometrics, 24(3):707–713, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528330.

Leyton:1968:NRC

[890] Morley K. Leyton. 251. Note: Rapid calculation of exact probabilities for 2×3 contingency tables. Biometrics, 24(3):714–717, September 1968.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528331.

Clarke:1968:NSA

[891] R. T. Clarke. 252. Note: The statistical analysis of a radio-active tracer experiment to determine root activity in potato plants. Biometrics, 24(3):717–722, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528332.

Jowett:1968:TBU

[892] D. Jowett and R. L. Chamberlain. [Teaching biometry in the university]. Biometrics, 24(3):723–725, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528333.

Linder:1968:OAF

[893] A. Linder. Obituary: Adolphe Franceschetti, 1896–1968. Biometrics, 24(3):726, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528334.

Balaam:1968:BRS

[894] L. N. Balaam. Book review: Statistical Methods for Research in the Natural Sciences by C. I. Bliss. Biometrics, 24(3):727–728, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528335.

Campbell:1968:BRE

[895] R. C. Campbell. Book review: Estimation Theory by R. Deutsch. Biometrics, 24(3):728–729, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528336.

Gomes:1968:BRS

[896] F. P. Gomes. Book review: Statistics of Crop Response to Fertilizers by FAO. Biometrics, 24(3):729, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528337.

Price:1968:BRB

[897] K. Price. Book review: Basic Statistics for Students of Economics, Commerce, Accountancy and the Biological Sciences by A. M. Goon, M. K.

Gupta, B. Dasgupta. Biometrics, 24(3):729–730, September 1968. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528338.

Glodek:1968:BRB

[898] P. Glodek. Book review: Biometrische Methoden in der Tierproduktion by A. Haiger. Biometrics, 24(3):730–731, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528339.

Glodek:1968:BRG

[899] Peter Glodek. Book review: Gleichzeitige Selektion nach mehreren Merkmalen und deren Anwendung in der Pflanzen- und Tierzüchtung by D. Herrendorfer. Biometrics, 24(3):731, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528340.

Glodek:1968:BRE

[900] P. Glodek. Book review: Elemente der Tierzucht by H. L. Le Roy. Biometrics, 24(3):731–732, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528341.

Loewen:1968:BRE

[901] D. J. Loewen. Book review: Econometric Techniques and Problems by C. E. V. Leser. Biometrics, 24(3):732–733, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528342.

Anonymous:1968:BRc

[902] Anonymous. [Books received]. Biometrics, 24(3):733–735, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528343.

Cavalli-Sforza:1968:BS

[903] Luigi Cavalli-Sforza. The Biometric Society. Biometrics, 24(3):736–743, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528344.

Anonymous:1968:NAc

[904] Anonymous. News and announcements. Biometrics, 24(3):744–747, September 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528345.

Anonymous:1968:FMd

[905] Anonymous. Front matter. Biometrics, 24(4):i–ii, December 1968. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528869.

Anonymous:1968:VI

[906] Anonymous. Volume information. Biometrics, 24(4):i–xxxviii, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528868.

Searle:1968:ALH

[907] S. R. Searle. Another look at Henderson’s methods of estimating variance components. Biometrics, 24(4):749–787, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528870.

Cox:1968:SOT

[908] C. Philip Cox. Some observations on the teaching of statistical consulting. Biometrics, 24(4):789–801, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528871.

Parker:1968:SAE

[909] Richard A. Parker. Simulation of an aquatic ecosystem. Biometrics, 24(4):803–821, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528872.

Lachenbruch:1968:EPM

[910] Peter A. Lachenbruch. On expected probabilities of misclassification in discriminant analysis, necessary sample size, and a relation with the multiple correlation coefficient. Biometrics, 24(4):823–834, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528873.

Jackson:1968:MVL

[911] Esther C. Jackson. Missing values in linear multiple discriminant analysis. Biometrics, 24(4):835–844, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528874.

Horton:1968:MCC

[912] I. F. Horton, J. S. Russell, and A. W. Moore. Multivariate-covariance and canonical analysis: a method for selecting the most effective discriminators in a multivariate situation. Biometrics, 24(4):845–858, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528875.

Balakrishnan:1968:DBP

[913] V. Balakrishnan and L. D. Sanghvi. Distance between populations on the basis of attribute data. Biometrics, 24(4):859–865, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528876.

Hanson:1968:SRA

[914] W. D. Hanson and Eduardo Casas. Spatial relationship among eight populations of zea mays l. utilizing information from a diallel mating design. Biometrics, 24(4):867–880, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528877.

Kuehl:1968:RPD

[915] R. O. Kuehl, J. O. Rawlings, and C. Clark Cockerham. Reference populations for diallel experiments. Biometrics, 24(4):881–901, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528878.

Hinkelmann:1968:MVP

[916] Klaus Hinkelmann. Missing values in partial diallel cross experiments. Biometrics, 24(4):903–913, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2528879. See correction [1061].

Yasuda:1968:EIC

[917] Norikazu Yasuda. Estimation of the inbreeding coefficient from phenotype frequencies by a method of maximum likelihood scoring. Biometrics, 24(4):915–935, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528880.

Wilton:1968:SIQ

[918] J. W. Wilton, D. Anthony Evans, and L. D. Van Vleck. Selection indices for quadratic models of total merit. Biometrics, 24(4):937–949, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528881.

VanVleck:1968:SBE

[919] L. D. Van Vleck. Selection bias in estimation of the genetic correlation. Biometrics, 24(4):951–962, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528882.

Rao:1968:EVC

[920] J. N. K. Rao. On expectations, variances, and covariances of ANOVA mean squares by ‘synthesis’. Biometrics, 24(4):963–978, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528883.

Nelder:1968:WRQ

[921] J. A. Nelder. Weighted regression, quantal response data, and inverse polynomials. Biometrics, 24(4):979–985, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528884.

Davis:1968:CCW

[922] C. E. Davis and Dana Quade. On comparing the correlations within two pairs of variables. Biometrics, 24(4):987–995, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528885.

Mantel:1968:CIA

[923] Nathan Mantel and Charles R. Stark. Computation of indirect-adjusted rates in the presence of confounding. Biometrics, 24(4):997–1005, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528886.

Becker:1968:SEF

[924] N. G. Becker. The spread of an epidemic to fixed groups within the population. Biometrics, 24(4):1007–1014, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528887.

James:1968:NIS

[925] J. W. James. 253. Note: Index selection with restrictions. Biometrics, 24(4):1015–1018, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528888.

Mantel:1968:NKS

[926] Nathan Mantel. 254. Note: Kolmogorov–Smirnov tests and Rényi’s modification. Biometrics, 24(4):1018–1023, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528889.

Anonymous:1968:CAM

[927] Anonymous. Corrections: Adjoint matrices for polynomial regression. Biometrics, 24(4):1025, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528891. See [850].

Anonymous:1968:CMM

[928] Anonymous. Corrections: A mathematical model for the estimation of inter-plant competition. Biometrics, 24(4):1025, December 1968. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528892. See [723].

Anonymous:1968:CNE

[929] Anonymous. Corrections: 241. Note: Equilibrium under selection at a multi-allelic sex-linked locus. Biometrics, 24(4):1025, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528890. See [826].

Anonymous:1968:CNT

[930] Anonymous. Corrections: 244. Note: On testing significance of components of variance in the unbalanced nested analysis of variance. Biometrics, 24(4):1025–1026, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528893. See [852].

Anonymous:1968:Ac

[931] Anonymous. Abstracts. Biometrics, 24(4):1027–1043, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528894.

Tukey:1968:TPT

[932] John W. Tukey. The true purpose of transformation (abstract 1545). Biometrics, 24(4):1041, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:1968:BSb

[933] Anonymous. The Biometric Society. Biometrics, 24(4):1045–1050, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528895.

Anonymous:1968:NAd

[934] Anonymous. News and announcements. Biometrics, 24(4):1051–1055, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528896.

Anonymous:1969:FMa

[935] Anonymous. Front matter. Biometrics, 25(1):i–ii, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528671.

Reyment:1969:MPG

[936] R. A. Reyment. A multivariate paleontological growth problem. Biometrics, 25(1):1–8, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528672.

Hashiguchi:1969:EGC

[937] Shoko Hashiguchi and Hiroko Morishima. Estimation of genetic contribution of principal components to individual variates concerned. Biometrics, 25(1):9–15, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528673.

Bosso:1969:AFA

[938] José A. Bosso, Oscar M. Sorarrain, and Ewald E. A. Favret. Application of finite absorbent Markov chains to sib mating populations with selection. Biometrics, 25(1):17–26, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528674.

Falk:1969:SSC

[939] H. Falk and C. T. Falk. Stability of solutions to certain nonlinear difference equations of population genetics. Biometrics, 25(1):27–37, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528675.

Haigh:1969:EPS

[940] John Haigh. An enumeration problem in self-sterility. Biometrics, 25(1): 39–47, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528676.

Patel:1969:SAD

[941] R. M. Patel, C. Clark Cockerham, and J. O. Rawlings. Selection among diallel classified variables. Biometrics, 25(1):49–61, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528677.

Brown:1969:ESD

[942] G. H. Brown. An empirical study of the distribution of the sample genetic correlation coefficient. Biometrics, 25(1):63–72, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528678.

TenHoopen:1969:MCT

[943] M. Ten Hoopen and H. A. Reuver. On measures of correlation in time series of events. Biometrics, 25(1):73–78, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528679.

Gregor:1969:ADD

[944] J. Gregor. An algorithm for the decomposition of a distribution into Gaussian components. Biometrics, 25(1):79–93, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528680.

Blumenson:1969:MAG

[945] Leslie E. Blumenson and Irwin D. J. Bross. A mathematical analysis of the growth and spread of breast cancer. Biometrics, 25(1):95–109, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528681.

Day:1969:TSD

[946] N. E. Day. Two-stage designs for clinical trials. Biometrics, 25(1):111– 118, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528682.

Bishop:1969:ITD

[947] Yvonne M. M. Bishop and Stephen E. Fienberg. Incomplete twodimensional contingency tables. Biometrics, 25(1):119–128, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528683.

Barnett:1969:SPL

[948] V. D. Barnett. Simultaneous pairwise linear structural relationships. Biometrics, 25(1):129–142, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528684.

Pahl:1969:TGF

[949] P. J. Pahl. On testing for goodness-of-fit of the negative binomial distribution when expectations are small. Biometrics, 25(1):143–151, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528685.

Han:1969:THV

[950] Chen-Pai Han. Testing the homogeneity of variances in a two-way classification. Biometrics, 25(1):153–158, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528686.

Patterson:1969:QBE

[951] H. D. Patterson. 255. Query: Baule’s equation. Biometrics, 25(1):159– 164, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528687.

Wishart:1969:NAH

[952] David Wishart. 256. Note: an algorithm for hierarchical classifications. Biometrics, 25(1):165–170, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528688.

Dahiya:1969:NFS

[953] R. C. Dahiya and John Gurland. 257. Note: Functions of the sample mean and sample variance of a Poisson variate. Biometrics, 25(1):171– 173, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528689.

Dudzinski:1969:NUI

[954] M. L. Dudziński and D. Bennett. 258. Note: The use of λ as an index of precision. Biometrics, 25(1):174–176, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528690.

Lever:1969:NCL

[955] W. E. Lever. 259. Note: Confidence limits for quantiles of mortality distributions. Biometrics, 25(1):176–178, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528691.

F:1969:BRC

[956] W. T. F. Book review: The Carcinogenic Action of Mineral Oils: a Chemical and Biological Study, by Carcinogenic Action of Mineral Oils Committee. Biometrics, 25(1):179, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528692.

Michalos:1969:BRF

[957] Alex C. Michalos. Book review: Fundamentals of Statistics. Vol. I, by A. M. Goon, M. K. Gupta, B. Das Gupta. Biometrics, 25(1):179–180, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528693.

F:1969:BRB

[958] W. T. F. Book review: Bibliography of Statistical Literature Pre-1940, by M. G. Kendall, A. G. Doig. Biometrics, 25(1):180, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528695.

Settle:1969:BRB

[959] T. W. Settle. Book review: Basic Statistics, by William L. Hays. Biometrics, 25(1):180, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528694.

Finucan:1969:BRP

[960] H. M. Finucan. Book review: Probability and Statistics for Engineers, by I. Miller, J. E. Freund. Biometrics, 25(1):180–181, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528696.

F:1969:BRE

[961] W. T. F. Book review: Energy Flow in Biology, by H. J. Morowitz. Biometrics, 25(1):181, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528697.

Stern:1969:BRF

[962] K. Stern. Book review: Forstliche Biometrie, by Michail Prodan. Biometrics, 25(1):181–182, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528698.

Warren:1969:BRF

[963] J. A. Warren. Book review: Forest Biometrics, by M. Prodan. Biometrics, 25(1):182, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528699.

Settle:1969:BRL

[964] T. W. Settle. Book review: Looking at History Through Mathematics, by Nicolas Rashevsky. Biometrics, 25(1):183, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528700.

F:1969:BRV

[965] W. T. F. Book review: Veb Deutscher Landwirtschaftsverlag Berlin und Verlag für Landwirtschaftliche Bücher und Zeitschriften Budapest. Biometrics, 25(1):183–184, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528701.

Cox:1969:BRE

[966] C. Philip Cox. Book review: Elementary Statistical Methods, by G. B. Wetherill. Biometrics, 25(1):184–185, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528702.

Anonymous:1969:BRb

[967] Anonymous. [Books received]. Biometrics, 25(1):185–187, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528703.

Anonymous:1969:Aa

[968] Anonymous. Abstracts. Biometrics, 25(1):189–196, March 1969. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528704.

Anonymous:1969:BSa

[969] Anonymous. The Biometric Society. Biometrics, 25(1):197–202, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528705.

Anonymous:1969:NAa

[970] Anonymous. News and announcements. Biometrics, 25(1):203–205, March 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528706.

Anonymous:1969:FMb

[971] Anonymous. Front matter. Biometrics, 25(2):i–ii, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528787.

Elandt-Johnson:1969:BIP

[972] Regina C. Elandt-Johnson. A. Biometrics invited paper. Survey of histocompatibility testing: Biological backgrounds, probabilistic and statistical models and problems. Biometrics, 25(2):207–283, June 1969. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528788.

Sykes:1969:DSP

[973] Z. M. Sykes. On discrete stable population theory. Biometrics, 25(2): 285–293, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528789.

Rouvier:1969:PVG

[974] R. Rouvier. Ponderation des valeurs génotypiques dans la selection par index sur plusieurs caractères. Biometrics, 25(2):295–307, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528790.

Usher:1969:MMF

[975] M. B. Usher. A matrix model for forest management. Biometrics, 25(2): 309–315, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528791.

Gates:1969:SSE

[976] Charles E. Gates. Simulation study of estimators for the line transect sampling method. Biometrics, 25(2):317–328, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528792.

Kimball:1969:MEC

[977] A. W. Kimball. Models for the estimation of competing risks from grouped data. Biometrics, 25(2):329–337, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528793.

Miettinen:1969:IMM

[978] Olli S. Miettinen. Individual matching with multiple controls in the case of all-or-none responses. Biometrics, 25(2):339–355, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528794.

Grizzle:1969:AGD

[979] James E. Grizzle and David M. Allen. Analysis of growth and dose response curves. Biometrics, 25(2):357–381, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528795. See corrections [1057].

Bishop:1969:FCT

[980] Yvonne M. M. Bishop. Full contingency tables, logits, and split contingency tables. Biometrics, 25(2):383–399, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528796.

Causton:1969:CPF

[981] D. R. Causton. A computer program for fitting the Richards function. Biometrics, 25(2):401–409, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528797.

Healy:1969:NRP

[982] M. J. R. Healy. 259 note: Rao’s paradox concerning multivariate tests of significance. Biometrics, 25(2):411–413, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528798.

Rees:1969:NSO

[983] David H. Rees. 260 note: Some observations on change-over trials. Biometrics, 25(2):413–417, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528799.

Kerr:1969:NES

[984] J. D. Kerr. 261 note: The estimation of the ‘short’ distribution. Biometrics, 25(2):417–420, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528800.

McCrea:1969:NHT

[985] J. M. McCrea. 262 note: The half-table ratio estimator for a simple exponential model. Biometrics, 25(2):420–423, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528801.

Broemeling:1969:NCI

[986] L. D. Broemeling. 263 note: Confidence intervals for measures of heritability. Biometrics, 25(2):424–427, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528802.

Gaylor:1969:NEM

[987] D. W. Gaylor and T. D. Hartwell. 264 note: Expected mean squares for nested classifications. Biometrics, 25(2):427–430, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528803.

Hyams:1969:SOT

[988] Lyon Hyams and T. J. Boardman. [ “Some observations on the teaching of statistical consulting”]. Biometrics, 25(2):431–434, June 1969. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528804.

F:1969:BRU

[989] W. T. F. Book review: United States Life Tables by Causes of Death: 1959-61, by F. Bayo. Biometrics, 25(2):435, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528805.

Anonymous:1969:BRA

[990] Anonymous. Book review: The Advanced Theory of Statistics. Vol. 3, by M. G. Kendall, A. Stuart. Biometrics, 25(2):435–436, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528806.

Sykes:1969:BRW

[991] Z. M. Sykes. Book review: World Population: an Analysis of Vital Data, by N. Keyfitz, W. Flieger. Biometrics, 25(2):436–437, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528807.

F:1969:BRSa

[992] W. T. F. Book review: Statistical Dictionary of Terms and Symbols, by A. K. Kurtz, H. A. Edgerton. Biometrics, 25(2):438, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528808.

F:1969:BRSb

[993] W. T. F. Book review: Studies of Inbreeding and Ageing in Female Mice, by L. E. Liljedahl. Biometrics, 25(2):438, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528809.

Balaam:1969:BRD

[994] L. N. Balaam. Book review: The Design and Analysis of Scientific Experiments, by K. C. Peng. Biometrics, 25(2):438–439, June 1969. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528810.

F:1969:BRT

[995] W. T. F. Book review: Transformations of Variables Used in the Analysis of Experimental and Observational Data. A Review, by H. Thöni. Biometrics, 25(2):439, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528812.

Solomon:1969:BRA

[996] D. L. Solomon. Book review: Applied Statistical Decision Theory, by H. Raiffa, R. Schlaifer. Biometrics, 25(2):439, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528811.

Anonymous:1969:BRc

[997] Anonymous. [Book received]. Biometrics, 25(2):439–440, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528813.

Anonymous:1969:Ab

[998] Anonymous. Abstracts. Biometrics, 25(2):441–448, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528814.

Anonymous:1969:BSb

[999] Anonymous. The Biometric Society. Biometrics, 25(2):449–453, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528815.

Anonymous:1969:NAb

[1000] Anonymous. News and announcements. Biometrics, 25(2):454–455, June 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528816.

Anonymous:1969:FMc

[1001] Anonymous. Front matter. Biometrics, 25(3):i–ii, September 1969. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528898.

Skellam:1969:MIS

[1002] J. G. Skellam. Models, inference, and strategy. Biometrics, 25(3):457– 475, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528899.

Hewlett:1969:MPD

[1003] P. S. Hewlett. Measurement of the potencies of drug mixtures. Biometrics, 25(3):477–487, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528900.

Grizzle:1969:ACD

[1004] James E. Grizzle, C. Frank Starmer, and Gary G. Koch. Analysis of categorical data by linear models. Biometrics, 25(3):489–504, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528901. See corrections [1056].

Gilbert:1969:EUV

[1005] Ethel S. Gilbert. The effect of unequal variance-covariance matrices on Fisher’s linear discriminant function. Biometrics, 25(3):505–515, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528902.

Samuel:1969:CSR

[1006] Ester Samuel. Comparison of sequential rules for estimation of the size of a population. Biometrics, 25(3):517–527, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528903.

Sheps:1969:MRD

[1007] Mindel C. Sheps, Donald P. Doolittle, and Mary L. New. Mammalian reproductive data fitted to a mathematical model. Biometrics, 25(3):529– 535, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528904.

Gibbons:1969:ENC

[1008] Jean D. Gibbons and Samuel Litwin. Estimation of the number of critical sites in limited genome expression during viral infection of bacteria. Biometrics, 25(3):537–544, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528905.

Watkins:1969:GCN

[1009] R. Watkins. Genetic components for non-inbred diploid species when the expectations for the epistatic variance are restricted. Biometrics, 25(3):545–551, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528906.

Machin:1969:PRS

[1010] D. Machin and R. M. MacIver. The prediction of response to selection in breeding programmes when all daughters of selected parents are retained. Biometrics, 25(3):553–560, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528907.

Freeman:1969:UCB

[1011] G. H. Freeman. The use of cyclic balanced incomplete block designs for non-directional seed orchards. Biometrics, 25(3):561–571, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528908.

Hartley:1969:DMM

[1012] H. O. Hartley and S. R. Searle. A discontinuity in mixed model analyses. Biometrics, 25(3):573–576, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528909.

Turner:1969:NGL

[1013] Malcolm E. Turner, Jr., Brent A. Blumenstein, and Jeanne L. Sebaugh. 265 note: a generalization of the logistic law of growth. Biometrics, 25(3):577–580, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528910.

Foss:1969:NMO

[1014] S. D. Foss. 266 note: a method for obtaining initial estimates of the parameters in exponential curve fitting. Biometrics, 25(3):580–584, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528911.

Watson:1969:NLR

[1015] G. S. Watson. 267 note: Linear regression on proportions. Biometrics, 25(3):585–588, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528912.

Arnold:1969:NMT

[1016] J. C. Arnold. 268 note: a modified technique for improving an estimate of the mean. Biometrics, 25(3):588–591, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528913.

Kulshreshtha:1969:NEM

[1017] A. C. Kulshreshtha. 269 note: On the efficiency of modified BIB designs for bio-assays. Biometrics, 25(3):591–593, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528914.

Gabriel:1969:NNP

[1018] K. R. Gabriel and P. A. Lachenbruch. 270 note: Non-parametric ANOVA in small samples: a Monte Carlo study of the adequacy of the asymptotic approximation. Biometrics, 25(3):593–596, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528915.

Davidson:1969:NRB

[1019] Roger R. Davidson. 271 note: On a relationship between two representations of a model for paired comparisons. Biometrics, 25(3):597–599, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528916.

Blumenthal:1969:BRS

[1020] S. Blumenthal. Book review: Sequential Identification and Ranking Procedures, by R. E. Bechhofer, J. Kiefer, M. Sobel. Biometrics, 25(3):601– 602, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528917.

McPherson:1969:BRP

[1021] D. G. McPherson. Book review: Principles of Statistics, by M. G. Bulmer. Biometrics, 25(3):602–603, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528918.

Peritz:1969:BRP

[1022] E. Peritz. Book review: Procese Stohastice si Aplicatii in Biologie si Medicină, by M. Iosifescu, P. Tăutu. Biometrics, 25(3):603, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528919.

Rao:1969:BRS

[1023] J. N. K. Rao. Book review: Survey Sampling by Leslie Kish. Biometrics, 25(3):603–604, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528920.

Schonemann:1969:BRF

[1024] Peter H. Schönemann. Book reviews: Faktorenanalyse, by K. Überla; Dimensionen des Verhaltens, by Kurt Pawlik. Biometrics, 25(3):604–607, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528921.

VanVleck:1969:BRE

[1025] L. D. Van Vleck. Book review: Evolution and the Genetics of Populations. Vol. 1. Genetic and Biometric Foundations, by Sewell Wright. Biometrics, 25(3):607, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528922.

Anonymous:1969:Ac

[1026] Anonymous. Abstracts. Biometrics, 25(3):609–611, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528923.

Bancroft:1969:BS

[1027] T. A. Bancroft. The Biometric Society. Biometrics, 25(3):613–614, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528924.

Anonymous:1969:NAc

[1028] Anonymous. News and announcements. Biometrics, 25(3):615–616, September 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528925.

Anonymous:1969:FMd

[1029] Anonymous. Front matter. Biometrics, 25(4):i–ii, December 1969. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528564.

Anonymous:1969:VI

[1030] Anonymous. Volume information. Biometrics, 25(4):ix–xxvi, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528563.

Cornfield:1969:BOA

[1031] Jerome Cornfield. The Bayesian outlook and its application. Biometrics, 25(4):617–657, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528565.

Goodman:1969:APG

[1032] Leo A. Goodman. The analysis of population growth when the birth and death rates depend upon several factors. Biometrics, 25(4):659–681, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528566.

Gould:1969:RTA

[1033] A. Lawrence Gould. A regression technique for angular variates. Biometrics, 25(4):683–700, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528567.

Reisch:1969:PTC

[1034] Joan S. Reisch and J. T. Webster. The power of a test in covariance analysis. Biometrics, 25(4):701–714, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528568.

Ashford:1969:GCA

[1035] J. R. Ashford and Susannah Brown. Generalised covariance analysis with unequal error variances. Biometrics, 25(4):715–724, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528569.

Feldman:1969:CSS

[1036] Sydney Feldman, Donald F. Klein, and Gilbert Honigfeld. A comparison of successive screening and discriminant function techniques in medical taxonomy. Biometrics, 25(4):725–734, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528570.

Bliss:1969:RCB

[1037] C. I. Bliss and Paul Griminger. Response criteria for the bioassay of vitamin K. Biometrics, 25(4):735–745, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528571.

Cannings:1969:GMS

[1038] C. Cannings. A graphical method for the study of complex genetical systems with special reference to equilibria. Biometrics, 25(4):747–754, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528572.

Hinkelmann:1969:EHE

[1039] Klaus Hinkelmann. Estimation of heritability from experiments with related dams. Biometrics, 25(4):755–766, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528573.

Thompson:1969:IEV

[1040] R. Thompson. Iterative estimation of variance components for nonorthogonal data. Biometrics, 25(4):767–773, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528574.

Marriott:1969:NAD

[1041] F. H. C. Marriott. 272. Note: Associated directions. Biometrics, 25(4):775–776, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528575.

Anonymous:1969:CAP

[1042] Anonymous. Corrections: Asymptotic power of chi square tests for linear trends in proportions. Biometrics, 25(4):777, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528576. See [843].

Anonymous:1969:CIP

[1043] Anonymous. Corrections: An iterative procedure for estimating fixed effects and variance components in mixed model situations. Biometrics, 25(4):777–778, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2528577. See [812].

Anonymous:1969:CNC

[1044] Anonymous. Corrections: Numerical construction of orthogonal polynomials from a general recurrence formula. Biometrics, 25(4):778, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528578. See [887].

Anonymous:1969:CSR

[1045] Anonymous. Corrections: Some row and column designs for two sets of treatments. Biometrics, 25(4):778, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528579. See [584].

Anonymous:1969:CTS

[1046] Anonymous. Corrections: Two sequential two-sample grouped rank tests with applications to screening experiments. Biometrics, 25(4):778, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528580. See [250].

Howell:1969:CPF

[1047] John R. Howell. [’A computer program for fitting the Richards function’]. Biometrics, 25(4):779, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528581.

Michalos:1969:BRMa

[1048] Alex C. Michalos. Book review: The Mathematical Sciences – A Collection of Essays by G. A. W. Boehm. Biometrics, 25(4):781, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528582.

F:1969:BRF

[1049] W. T. F. Book review: Functions and Graphs by I. M. Gelfand, E. G. Glagoleva, E. E. Shnol. Biometrics, 25(4):781–782, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528583.

Michalos:1969:BRR

[1050] Alex C. Michalos. Book review: Readings in Mathematical Social Science by Paul F. Lazarfeld, Neil W. Henry. Biometrics, 25(4):782, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528584.

Michalos:1969:BRMb

[1051] Alex C. Michalos. Book review: Mathematical Thinking in Behavioral Sciences by David M. Messick. Biometrics, 25(4):782–783, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528585.

Anonymous:1969:BRd

[1052] Anonymous. [Books received]. Biometrics, 25(4):783–785, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528586.

Anonymous:1969:Ad

[1053] Anonymous. Abstracts. Biometrics, 25(4):787–805, December 1969. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528587.

Anonymous:1969:BSc

[1054] Anonymous. The Biometric Society. Biometrics, 25(4):806–815, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528588.

Anonymous:1969:NAd

[1055] Anonymous. News and announcements. Biometrics, 25(4):816–818, December 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528589.

Anonymous:1970:CAC

[1056] Anonymous. Corrections: Analysis of categorical data by linear models. Biometrics, 26(4):860, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528736. See [1004].

Anonymous:1970:CAG

[1057] Anonymous. Corrections: Analysis of growth and dose response curves. Biometrics, 26(4):860, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528735. See [979].

Anonymous:1970:CSV

[1058] Anonymous. Corrections: The sampling variance of the correlation coefficients estimated in genetic experiments. Biometrics, 26(4):860, December 1970. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2528737. See [599].

Anonymous:1971:CIB

[1059] Anonymous. Correction: Incomplete block designs for bio-assays. Biometrics, 27(4):1103, December 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2528848. See [670].

Anonymous:1974:CTP

[1060] Anonymous. Corrections: The two-period change-over design and its use in clinical trials. Biometrics, 30(4):727, December 1974. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2529243. See [511].

Anonymous:1979:CMV

[1061] Anonymous. Correction: Missing values in partial diallel cross experiments. Biometrics, 35(4):905, December 1979. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2530135. See [916].

Anonymous:2005:CK

[1062] Anonymous. Correction of Kodlin (1967). Biometrics, 61(1):311, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [726].